A Complete Bibliography of Publications in Biometrika for the decade 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.06

Title word cross-reference

#15763 [BR81]. #15917 [HP80]. #4143 [AR81]. #4315 [Har79].

(p, q) [And75c]. 25 [Mul77]. 2^{k} [Fie72]. 2 × 2 [Alt71b, Cop73, Gar70, Ham70, Mee78, Pla77, Whi78, Zel71]. 2 × 2 × 2 [VB78]. 2 × s × t [Gar72b, Gar73]. b_{2} [Bow73, BS75, DT71, DP73, DP74]. C_{P} [Spj77a, Spj78]. χ [JP69, JP70]. D [D’A72, Hil78, Pes75, Pes76, Tit75]. E(X) = E{E(X|Y)} [Eni73]. F [AP70, And71b, BT73, DR73, Dav75, Dav76a, Dav70b, Hir79, McK74, MH71, PT70, Pre74a]. G_{p, p}^{p,0} [Dav79]. G_{1} [DS75, D’A70a]. I × J × K [HWB^{+}77]. K [Bre70, Pol75, GP70, HH71a, Lew72, Ode72]. M [BN73, Kor72, AR72, AR81, Pre79a, Saw78]. G_{∈} [JOZD70]. N [Mor79, SF79, She71a]. R [Kor72, MH78]. R^{2} [Hos79]. ρ [Ham70]. R × C [LB77]. S_{B} [Buk72, JK71a, JK71b]. S_{U} [Joh74b]. σ [Pre71b]. \sqrt{b}_{1} [BS73b, DT73, BS75, Mul77]. √b_{1} [DP73, DP74]. t [BDJS78, Boh73, Dut75, FK72, HH71a, Hin77b, Hin78a, Mic75, Pre74c, SS70, YD73, Yue74]. T^{2} [Kor72]. T_{0}^{2} [Dav70a, Dav72, McK74, HS72b, Tik71b]. U [Blo76, LS76b, LS77a, LS77b]. V [CR72]. V_{n} [BS73a]. W′ [Wei74, Kor72, Wei74]. X < t^{2}/X [Juo70].

-Ancillarity [BN73]. -Ary [SF79]. -estimates [MH78]. -Independence [DR73]. -Minimax [JOZD70, DS75]. -Optimal [Hil78]. -Optimality [Tit75]. -optimum [Pes76, Pes75]. -Sample [Pol75, Lew72, Ode72]. -sphere [Saw78]. -Statistics [GP70, LS77a, LS76b, LS77b, Blo76].

1 [AR72]. 1977 [Efr79].

30 [Har79].

49 [AR81].

51 [Har79]. 52 [BR81, HP80, Joh74b]. 54 [MT72]. 55 [DS76a]. 57 [Dav72]. 58 [Ken72]. 59 [AR81, Joa75].

60 [DP74, Rao73a]. 62 [BR81, Bin76, CS76, Dav76a, Hin76a, HP76, HP80, Lou76, Pes76, Sch76, Smi76, The76]. 63 [DC77, LS77a, Sha78b, Wil78]. 64 [Ali78, Hin78c]. 65 [Pre83]. 66 [CS82, JW81].

80a [Pre83]. 80i [JW81]. 81f [CS82].

A. [Goo79a]. Abbe [Ken72, Ken71]. Abraham [DP72]. Absolute [Lac75, HH71a]. Absorbing [Mat70]. abstract [Tuk58]. Acceptability [Joh74a]. Acceptance [BS76b]. Accuracy [EH78b, RF70, Sow71]. Adding [Cha76]. Additions [AP70]. Additive [Dar74]. additivity [Yat72]. Adequacy [God79]. Adjusted [AL76, ER77, Fea77]. Admissible [FV71, Van73b]. Adult [Seb70]. Affinity [CD78]. Against [Fuj79, Kno75a, Kno77, LS76b, LS77b, Mar76a, MP78, Pet72, Spi77, Sto77, AS72, LS77a, Ode72]. Age [Seb71, CH72, Sti70]. Age-Specific [Seb71]. Aggregates [Tia72]. Aggregation [TW76]. Agnesi [Sti74]. Agreement [Gor79a, Gor79b, HS78c]. AIC [Aka79]. Aided [BP74]. Akaike [BD77, Koh77, Shi76]. Algebraic [KP76]. Algorithm [Ans79, BS00, FK74, Geb70, Gor73, Hit03, deC72]. Algorithms [Atk79]. Allocation [FL71, GJ79b, Gla78, HHvdB74, Lou75, Mor75a, New71, SY74, ST71, Wei78, Lou76]. Allowing [Pol75, RK75]. Alternative [Dav77, Hay75, Kno77, Leo76, Mar76a, MH71, Sow72]. Alternatives [AS72, Bro74b, Cha74b, Fuj79, Hir78, Kno75a, Koe79, LS76b, LS77b, Nel77, Pet72, Spi77, Spj77a, You73a, LS77a, Ode72, Spj78]. Amalgamation [Ait79]. Amendments [Ali78, Ano81, AR81, Bel79, BR81, Bin76, BMW79, Cha73a, CS76, Cox77, CS82, DP74, Dav76a, DPF74a, DH77, De 77, DS76a, DC77, Gar73, Har79, HPF71, Hil77, HH71b, Hin78b, Hin78a, HP76, HP80, Hol72a, JW81, Kei74a, LSG83, LMH71, Ofo75, Pea72, Pes76, Pfa74a, Pre83, Rob77, Rog78, Saw79, Sch76, Sha78b, Som76, Spj78, Tar90, The76, Wis90, Yan80, You72a, Dav72, Hin76a, Joa75, Ken72, LS77a, Lou76, MT72, Smi76]. Among [Ano81, BO78, Pre75]. Amongst [GL70]. Analogy [Fin79a]. Analyses [HP74, Hol78, Ran74]. Analysis [AB79, Ali77, Ali78, Ali79, Alt71a, Alt71b, Alt76, AC77, BM79, Bin78, BT77, Bro74b, CC77, Cha77, Che71, Cla74, CS79a, Cor77, Cox79, Daw77, Dun76, Feu79, FV79, Gab71, Gol76, Goo73, Goo74, JW71, Jör70, KHB70a, Ket71, Kor72, Krz79, LW78, LM73, Lee79, MPG76, McC78, McK77, McL74, McL76, ML74b, Mor78, MDS74, Nea70, Nel77, Olk72, PCdCM74, Poc77, RS78c, See75, SM72, SWH75, Tar75, The75, TP74, TA71, Tit77, UL76, Wal67, Wal70a, Whi78, WJ79, Wil70a, WL77, Wor75, DS75, MH78, The76, Zel71]. Analytical [Rao73b, Rao73a]. Analyzing [CL76]. Ancillarity [BN73, FR79]. Anderson [Pet76b]. Anderson-Darling [Pet76b]. Angular [Upt74]. Animal [Hol73]. Animals [Ano81, BO78]. Anomalies [CC77]. Ansari [Lep71, Lep73]. Ansari-Bradley [Lep71, Lep73]. Apparent [McL76]. Application [BEH76, Cla78, Gab71, GK71, Hir79, The75, Web70, Whi78, The76]. Applications [Dur75, Has70, Mar70a, Olk72, HH71a]. Applied [Bro78, Mar79b]. Approach [BO75, BE72, Bro72, DS73, DS68, Fea75, Gei74, Goo73, HG76, KS72, KM73, MW77, Rob78c, Shi78, ST73, ST74, Smi75, Sri70, You77, DS76a, Smi76]. Approaches [DT73, Leo76]. Approximate [BS79, Dav75, Gol74, Mic75, MH71, Que49, Ram72, Sis76, Dav76a, YD73]. Approximately [MG72]. Approximating [HS72b, Ped79]. Approximation [Hir79, Kon78, SS70, She71a]. Approximations [Dav70b, Dav71, Hil76, Hil77, Hin77b, Law74, McC73, Phi78, SH76, Upt74, DT72, McK74, Hin78a]. Arising [CC77, Atk74]. Armed [Jon75, WAK77]. Arthur [Mac77a]. Ary [SF79]. Aspects [Bro73, Gil79, Tit75]. Assay [Arm70]. Assessing [EH78b, Gea70]. Assigned [Sam70]. Assignment [HS75]. Associated [DB69, DB70, Ked76, Mor75a, Pee71, Sti71]. Association [Cla78, Tjø78]. Associations [Mes73]. Assuming [GLGH79]. Assumptions [HM75b]. Asymmetrical [LJ76]. Asymmetrically [Pre79b]. Asymptotic [Bai79, Bri74, BW78, Cha74a, Dun76, FR79, Hai71, Hol72b, Hol72a, Hos79, Kay79, Lee71, McL74, Mor73a, Oak77, Roy77, Ste75a, Tho76, Tia72, Wat76]. Asymptotically [Mor70b, DL73]. Asymptotics [Sto77]. Attenuation [BP75]. Attributable [Wal75]. Attributes [HG76]. Autocorrelated [BW74, Bli73, Kno75a]. Autocorrelation [AL71, And76b, Bas72, DC78, HW78, HM75a, Kno77, Wic73]. Autocorrelations [And75b, AN79, DPF74a, DPF74b]. Autocovariance [And76a, Roy77]. Autocovariances [Tua79]. Autoregression [Gea70, Phi78, Tro79]. Autoregressions [dJ76]. Autoregressive [Aka73, Aka79, Ali77, And75d, Ans79, AN79, Bai79, BD77, BS76b, CT77, God75, Han79, HP79, Hos79, Koh77, Lev72, LB79, McC73, New74, NH79, PDT78, Pie71, Ree72, Rud74, Sha73, Shi76, Tua79, Ali78, And75c]. Autoregressive- [HP79]. Autoregressive-Moving [Ali77, And75d, Ans79, AN79, CT77, Han79, Hos79, Koh77, LB79, New74, NH79, PDT78, Pie71, Sha73, Tua79, Ali78]. Average [Aka73, Ali77, And75b, And75d, And76a, And76b, Ans79, AN79, Bha76, CT77, DPF74a, DPF74b, God75, God77, Han79, HP79, Hos79, Koh77, LB79, New74, NH79, PK78, PDT78, Pie71, Roy77, Sha73, Tua79, Wic73, Ali78, And75c]. Averages [God78, McC73].

B. [BN78, Jam78, Rob78a, Spr78]. Back [Ano70a, Ano70b, Ano70c, Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, Ano72c, Ano73a, Ano73b, Ano73c, Ano74a, Ano74b, Ano74c, Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano79b, Ano79c]. Bacon [Wil75a]. Bad [And71b]. Balanced [Efr71, SF79, Sin78b, Wil75b]. Band [Seb70, Seb71]. Bandit [GJ79b, Jon75, WAK77]. Bands [KS72, SW73]. Bartholomew [Sis76]. Bartlett [SHA78a]. Based [And71c, BS76b, BS75, Cre78, Cre79, Far79, GM78, HBA70, HS79, KP73, Pre74a, Sco79, Sim77, Ste79a, Tam78, Ton71, You70a, You72a, HR77, Lew72, MH78]. Basic [AB76, BEH76, God75, PCdCM74]. Batch [BS76b]. Bayes [CK70, Aka78, BM72, CK68, Cop72, CD78, Cox75b, Cox75c, EM72, EK72, GK70, Geo71, GK71, GD74, Lai78, Leo74, ML74a, O’H73, OL76, Pre71c, RK75, Smi73, SD72]. Bayesian [Smi76, AB79, Aka79, Alt71b, Bin78, BO75, Bre79, Bro74b, CC77, DS73, ESF73, Fea75, Fre70, GJ79a, Gol74, Gol76, Har74, Jon74b, Leo72, Leo73, MEA76, Rob78b, Sha77, Smi75, SD72, WAK77, Wat74, Woo76, You77]. be [DGS74, Efr71, Kha78]. Behaviour [BW78, Roy77, Tia72, TW79, Wic73, Yat72, YD73]. Behrens [Pfa74a, Dal78, MS70, Nai74, Pfa74b, SS73]. Behrens-Fisher [Dal78, Nai74, Pfa74b, SS73]. Bernoulli [She70]. Best [KMN75, LC76, TW79]. Beta [Atk79]. Better [HSW72, SW70, TY74, Fus73]. Between [AG74, AF75a, Els75, Fre73b, Han75, HS78c, Jar79, Mes73, PT70, PP75, Yan78, Yan80, AF75b, BW71, Gor79a, Gor79b]. Bhattacharyya [BR74, Whi73b]. Bias [McC73, McL76, Que56, RF70, Sow71, Sow72, Tuk58, Wei78]. bibliographical [DP72]. Bibliography [Cox70a]. Bienaymé [HS72a]. Bimodal [AS72]. Binary [AA76, Bro73, Far77, GP75, Krz79, San71, Tit77]. Binomial [Bes74, Bin75, Bis72a, Bis72b, GK71, HH71b, HSW72, Jon74b, Leo72, ML74a, NS72, RS78c, SWH75, SW70, TY74, Tar79b, Tar90, UL76, Wal70b, Wis68, Wis90, You70b, You73b, Bin76, Fus73, HH70]. Bioassay [Chm76, Mil73]. Biometrika [HP80, Ali78, AR81, BR81, Bin76, CS76, CS82, DP74, Dav76a, Dav72, DS76a, DC77, Har79, Hin76a, Hin78c, HP76, Joa75, Joh74b, JW81, Ken72, LS77a, Lou76, MT72, Pes76, Pre83, Rao73a, Sch76, Sha78b, Smi76, The76, Wil78]. Biplot [Gab71]. Bird [Seb71]. Bird-Band [Seb71]. Birds [Seb70]. Birth [Kei74a, Kei74b, Ren72, vK73]. Bivariate [Ahs71, Cla78, HS79, Hol71a, Pal72, Sam70, Shi78, TN79, Wyn77]. Black [Mac77a]. Block [BT74, CS74, CS75, Dow76, Fre76, Jar77, JH78, KHB70b, PT71, PW76, Pea70, Pea71, Pea72, SF79, SD79, Sin78b, Wor75, CS76, Woo78]. Blocking [Rob70]. Book [Cox70a, Cox72a, Cox72b, HW72, Sto72]. Bootstrap [Efr79]. Borne [Aba74, Dan72, Den72]. Both [Bho78, Bho79, Ekb76b, Goo74, JK71b, LS77b]. Bound [Loy79, MS76, deC72]. Boundary [MV76]. Bounded [Dar71]. Bounded-Sum [Dar71]. Bounding [And75b]. Bounds [AG74, BR74, BF72, CS79a, Coo79, Jar77, MG72, Mui74, Sat77, Wyn75]. Box [DTN77, McL77]. Box-Jenkins [McL77]. Box-Pierce [DTN77]. Bradley [Atk72b, Lep71, Lep73]. Bradley-Terry [Atk72b]. Branch [FK74, deC72]. Branching [BS76a, Bec74, Hey75, San71, Sti71, Swe78, CH72, HS72a, Sti70]. Burr [Rod77].

C [SBL79]. Calculated [CS71]. Calculations [Bar65, FH70a, HMM77, Kal74, Kay79, MRR^{+}53]. Calibration [Ait77, Ait79]. Cambridge [Boo75]. Cancer [HJ79]. Canonical [BT77, Che71, FV79, JW71, Ket71]. Capture [Ano81, BO78]. Carcinogenesis [Ber75]. Carlo [Bar65, DL73, Has70, Hut71, Lay74, LDG75, Mor75b, Pes73, Tik70, Yat72]. Carrier [Aba74, Dan72, Den72]. Carrier-Borne [Aba74, Dan72, Den72]. Case [AAK72, Ekb76a, Ekb76b, Kno75a, MN74, PP79, Saw70, You73a, DS75]. Case-Control [PP79]. Cases [AB76, HHvdB74]. Catchability [Pol75]. Categorical [Cla74, Cla76, McC77]. Categories [Goo79b, McC78]. Categorized [HG76]. Cauchy [Cop75, HBA70, Law72, Sti74]. Causes [BD76]. Cautionary [Dye74]. Cell [CH74]. Cells [Str77]. Censored [Bli70, BJ79, Cla76, DF79, Far77, HS79, Hea78, JW79, KM79, LW78, LSZ78, Lat77, Mac72, MN74, Mil76, Oak77, PR77, Pet76a, Pre78, Pre70, Pre79b, Saw58a, Saw58b, Saw79, Tik70, Tik71a, WL77, You73a, JW81, KG76, PS76, Pet77, Pre83]. Censoring [Hyd77, LGGB79, Pre73, You70a, You72a]. Censorship [Bre70]. census [Fie72]. Central [JP69, Mat70, PT70]. Centre [Luc77]. Century [Har10]. Certain [Cra72, DGK71, Hol72b, Hol72a, KLC76, Lee71, Man70b, Rea73, Roy70, Upt74, Wed76, Wyn77]. Certainty [AB76]. Chain [Dev73, Han72, Ked76]. Chains [Bas72, Has70, Mat70, Pes73, Sha76a, Sha75, Sha78b]. Change [HH71b, Hin71, Kie75, Pat70, Smi75, Hin70b, HH70, Smi76]. Change-Over [Pat70]. Change-Point [HH71b, Hin71, Smi75, Hin70b, HH70, Smi76].

Changeover [Pat73]. Changepoint [Cob78]. Changes [Gea70]. Characteristic [Dav73]. Characterization [Her76]. Charlier [DT72]. Chi [Bri74, Dah71, DG72, Goo69, Goo70, Han75, Har64, Kha78, Mol77, Sie79, Sla73, Har65, Har79]. Chi-Square [Har64, Har65, Har79]. Chi-Squared [Bri74, Dah71, DG72, Goo69, Goo70, Han75, Kha78, Mol77, Sie79, Sla73]. Choice [Bro74c, Han72, Man70a, Rob78c]. Choosing [Atk78, HSW72, Sil78a]. Chronic [Cla78]. Circle [BM78a]. Circular [LO78, Mar76b]. Class [Afo72, Ber74, CS74, Dav71, DTL76, LW78, McC78, PW76, Sub69, Sub70]. Classical [GO77, Saw70]. Classification [BW78, Har69, Har70, Hin72]. Classifications [MT67, MT72]. Classified [AB76]. Clinical [Ait79, BM79, FL71, Gla78, MI77, Poc77, SWH75, UL76, WJ79]. Closed [Ano81, Bar75a, BO78, MPG76, Fie72]. closer [Har10]. Closest [Mor79]. Cluster [Bin78]. Clustering [Dom72, FV71, HIR72, KR76, Str75, Van73b]. Coal [Jar79]. Coal-Mining [Jar79]. Coefficient [Bas72, DC78, Hin77a, Hin78b, Joh79, Kon78, Lee72b, MP78, Ram72, Thi74, Boo75, Hin78c]. Coefficients [Cho77, DGK75, LO78, Mar76b, Mor70a, Wol76, Woo70, DL73]. Coherence [HT71]. Columns [SD79]. Combination [Arm70, Gar70, RW75, Lep71]. Combinations [Rab70]. Combined [McK78, Lep73]. Comments [Bar75b, BN75, BN78, Bir75, EH78a, Fin79b, Hea79, HS78b, Jam78, Kal75a, Kem79, Kor72, McL75, Rob78a, Sch78b, SBL79, Spr78]. Common [Cha76, Gar70, Mac77b]. Comparative [BM79]. Comparing [Bre70, Ekb76a, Ekb76b, Gai74, GW75, Lou75, Wil70b, Lou76]. Comparison [Cla76, DS73, DH75b, FK74, Fre73b, FR72, Gor73, GP75, Gue72, Hei78, HS75, Kap70, Koc79, Lay74, LS75b, LGGB79, LB77, MT77, Nic77, PDB77, Pre76, Sli70, ST71, TY74, Wei74, dC71, deC72, Hut71]. Comparisons [BEH76, BP76, DB69, DB70, DB71, De 77, DS76b, Dye74, FL76, GL70, Lee71, McC77, MT67, dBP77, SHA78a, MT72]. Competing [BD76, Hol78, Sea77]. Complete [Wil76, Wil78]. Complex [KLC76, SD76]. Component [Gab71, See75, Wal71]. Components [Afo72, Har69, Har70, Har74, Mac74b, MH71, TA71]. Composite [Cha74a, Hay75, Joa72, Mor70b, Joa75]. Compound [Hol70, Ray71]. Compounds [And79]. Computation [PK78]. Computer [Atk79, BP74]. Computer-Aided [BP74]. Computing [MRR^{+}53, Dut75]. Concentration [Sch78a]. concept [Sti73]. Concerning [Hem77, Koh77, Sha75]. Concomitant [BP76]. Concomitants [BGR76]. Concordance [LS75a, Woo70]. Concurrences [Fre73a]. Conditional [Cob78, GP70, God76a, Hin77a, Hin78b, Hin78c, Law72, McL74, Mee79, Ped79, Spr75]. Conditionality [Kal75b, McL75]. Conditioned [CL76, Cox76c]. Conditions [Cha74a, Goo69, Goo70, Kha78]. Cone [LW78]. Confidence [AKH70, AR72, BF72, BMW78, BMW79, Buh77, DS76b, FH70a, Hin77b, Hin78a, KS72, Law72, Law74, MG72, Mar79a, Moh79, Odé73, Rob75, Rob77, Spj72, SW73, Swi76, Tar79a, Van70, Van73a, Wal70b, Wyn75, AR81, Dal78, Tuk58]. Configuration [You73b]. Conflict [CS73]. Confounding [BGP77, Bai77]. Connected [Wyn76]. Connectedness [ER75, Her76]. Consideration [TH78, You70a, You72a]. Considerations [Sha77]. Consistency [DS72, De 77, dBP77, Sar75]. Constant [KM79, Seb71, WL77]. Constant-Sum [KM79, WL77]. Constantine [BBLS00]. Constrained [Lee79]. Constraints [OL76]. Constructing [AH73, Fre76]. Construction [Bro77b, Dey77, HPF70, HPF71, KS72, Rob78b, SW73]. Contagious [BAS70]. Contamination [FH70b]. Contingency [Dar74, Gar73, Goo73, Goo79b, GD74, HWB^{+}77, Lai78, McC78, Mee78, MT77, SD76, VB78, Whi78, Wil76, Fie72, Gar72b, Wil78, Zel71]. Continuity [Cox70b]. Continuous [BM72, Krz79, PW75, PDT77]. Contours [BS75]. Contrasts [BEH76, Har74, PCdCM74]. Control [Bro77a, Cha74b, GK70, PP79, SY74, Sli70, ST71, Wei78]. Controllable [Fed74]. Controlled [Gla78]. Convergence [BR74, God78]. Convex [And76b, O’H73]. Correct [Som71b]. Correction [Ait71, BGKO71, BP75, CK70, Cox70b, Dan70, HPF71, Hin70a, HH71b, Hin78c, JP70, Wil78]. Corrections [Ali78, AP70, Ano81, Bel79, Bin76, BMW79, Cha73a, CS76, Cox77, CS82, DP74, Dav76a, DB70, DPF74a, DH77, Dav72, De 77, DS76a, DC77, Gar73, Goo70, Har70, Hil77, Hin76a, Hin78b, Hin78a, HP76, Hol72a, Joa75, Kei74a, Ken72, LSG83, LMH71, LS77a, Lou76, MT72, Ofo75, Pea72, Pes76, Pfa74a, PG70, Pre83, Rao73a, Rob77, Rog78, Saw79, Sch76, Sha75, Smi76, Som76, Spj78, Sub70, Tar90, The76, Wal70a, Wis90, Yan80, You72a, AR81, BR81, Har79, HP80, Joh74b, JW81, Sha78b]. Correctly [LC76]. Correlated [Bho78, Bho79, Bro74a, DL72, Gas75, GNP73, Hin69, Hin70a, Mor73b, Nai75, Wil70b, TA71]. Correlation [Ait69, Ait71, BGR76, Bes72, Che71, Cho77, DGK75, DW71, Els75, FV79, HM76, Hin78d, Joh79, Kon78, Lee72b, LO78, Mar76b, MP78, MG73, Mor70a, Ram72, Rot79, Sam70, SPB78, Shi77, Ste79b, Wol76, Boo75, DL73, Que49]. Correlations [Els75, Ros79]. Corrigenda [Ano48a, Ano48b, Har65, Saw58a]. Count [Gea70]. Counterexamples [Rob75, Rob77]. Counts [Mac72]. Covariables [Car77, KM77]. Covariance [BT74, Cha76, Fea77, Jör70, KLC76, Lay74, Lee79, Mui74, NP73, Sha73, Ste75a, TP74]. Covariances [Pre74d]. Coverage [LS77c]. Covey [Gut71]. Covey-Crump [Gut71]. Cox [Cor77, DC77, KP73]. Cramér [Pet76a, God76b, HR77, KG76, PS76, Pet77, Pet78, SG74, ZP77]. Cramér-von [HR77, KG76, PS76, Pet77, Pet78, SG74, ZP77]. Cramér-Wold [God76b]. Criteria [Dav71, GP75, HP72, JM78, OW78, Pee71]. Criterion [BD77, Box71, God78, Hay75, Kie75, Kor72, Lee72a, NP73, PG69, PG70, Shi76, Dav79]. Critical [BY77, Sis76]. Cross [SS77, Sto74, Sto77]. Cross-Product [SS77]. Cross-Validation [Sto74, Sto77]. Crossings [BL76, Ked78, Lin74]. Crump [Gut71]. Cube [Pes75, SHA78a, Pes76]. Cuboidal [Atk73]. Cumulative [Hin71, KS72]. Cure [O’N79]. Curtailment [PS73]. Curve [BCK78, Mac70, DP72]. Curves [Buk72, Fea75, Fea77, JK71a, JK71b, LGGB79, MV76, Joh74b]. Cusum [BE72]. Cut [Bro75]. Cyclic [Fre76, HW73, JH78, Joh73, Rog77, Rog78].

D [Bar75b, BN75, BN78, Bir75, Cox72a, Hea79, Jam78, Kem79, McL75, Rob78a, Spr78, Sto72]. Daniel [She70]. Darling [Pet76b]. Data [AL79, And71a, Bec79, Bho79, BJ79, Cla74, Cla76, Cox75a, Cox76a, Cox77, Cox79, DR73, DF79, Ekb76a, Ekb76b, Far78, Far79, Gas75, JW79, KM79, Kay79, LSZ78, Lat77, Leo72, LS74, Lit76, MG72, MM76, McC77, Mil76, ML74b, Mor73b, Mur79, Nai75, Nea70, Oak77, PCdCM74, Pre78, Rog77, Rog78, RS78c, Rub76, Sco79, Tit77, DC77, JW81, KG76, PS76, Pet77, Pre83, Yat72]. Data-Based [Sco79]. Data-Splitting [Cox75a]. David [Boo75]. Death [Ren72, vK73]. Decision [Bro72, Ray71]. Decisive [Man70b]. Decomposition [Gil79, JW71, TH78]. decrease [Mar70b]. Decreasing [LS79]. decrements [Sea77]. Definition [Dar71, Zen75]. Degree [GL70]. Degrees [Dav75, Pre74a, Sie79, Dav76a]. Degrees-of-Freedom [Pre74a]. Delay [HT71, HT73]. Dempster [CC77]. Densities [BP74, Fel74, GG71]. Density [Bru78, Cox76c, Dig75, Dig77, Gra79, Lew75, LS77c, Mur77, Mur79, PDT77, Sil78a, Sin79]. Departure [PDB77, DP73, DP74]. Departures [Bow73, BS75]. Dependence [Gre70, RT73]. Dependent [Cho77, KM77, Thi74]. derivation [DP72]. Derivatives [Sin79]. Description [Fin79a]. Design [AF75a, AF75b, BW74, Ber75, Box71, Bro77a, Bro73, Cot79, Her76, Hil78, JM78, Kie75, KP76, Lay75, Luc77, PCdCM74, Poc77, ST73, ST74, Sil78b, Sin78b, Tit75, Vuc77, Wei78, Yan76, Atk70, Woo78]. Designed [And71b, AH73]. Designs [Afo72, Atk73, BGP77, Bai77, Bec70, Bel77, Bel79, BD75, Bro72, BP76, CH75, CS74, CS75, Cor72, CCS70, Dey77, DGP77, ER75, ER77, Fre73a, Fre76, Gut71, HW73, HPF70, HPF71, HC72, HJ79, Jar77, JH78, JMM75, Joh73, KHB70b, LB77, Pat70, Pat73, PW76, Pea70, Pea72, Pes75, PPK73, Rob70, SS75, SF79, SD79, Wil75b, CS76, Pes76]. Detect [Atk72a]. Detectible [San71]. Detecting [JM78]. Detection [DGK75, HH79]. Determination [Cox76b, Yan78, Yan80]. Determining [Yan76]. Deviation [Ano48a, Ano48b, Gea36, Hea78, Pre70, Pre71a, Pre79b, SBL79, Tik71a]. Deviations [AG74]. Diagnosis [AB76, Ait77, Ait79, AL79, BP74]. Diagnostic [Bro77b, GP75]. Diagonal [MT77]. Dichotomous [Gil79]. Dictionary [Cox70a]. Did [ET76]. Difference [Buh77, CSW76, Hea78, Lac75, LS74, Pre79b, Sha70a]. differencing [And75c]. Different [BCK78, DGP77, GP75, Gre79, Kap70]. Differential [BP74, Dav79]. Differentials [GSW75]. Dilution [Hwa76]. Dilutions [CH75]. Dimensional [Bes72, DG76, Lew75, PPK73]. Dimensionality [FV79]. Dimensions [Adl77, AH76, Bel77, Bel79, FK72]. Direct [Fre73a, God75, God77]. Direction [Yan76]. Directional [ML74b]. Directions [AS72, Lay75]. Dirichlet [Olk72]. Disasters [Jar79]. Discounted [GJ79b]. discovery [Pla72, Sti73]. Discrete [Alt76, BP74, Cox70a, WA78]. Discriminant [AC77, GM78, Krz79, Lac75, McK77, McL74, McL76, Mor75a, ML74b, DS75]. Discriminating [AF75a, AF75b]. Discrimination [AA76, And72, And75a, Lac75, LMH67, LMH71, Pre75]. Disease [BAS70, Bec79, BP74, Cla78, PP79]. Diseases [Bec72]. Dispersed [Ros73]. Display [Gab71]. Dissecting [Cha76]. Distance [CL76, Cox76c, Dig75]. Distributed [Car77, Kha78]. Distribution [Ahs71, AN79, AKH70, AR72, Bar65, Ber75, BM78a, Bin75, BGKO66, BGKO71, BS73b, BMW78, BMW79, Bri74, BE72, CR72, Cop75, D’A70a, DT73, DGK71, GP70, Goo69, Goo70, GK71, HBA70, Har64, Has71, Hil76, Hil77, Hin76b, Hos79, Joh72, KS72, Koc79, Kon78, LS79, Lee71, Lee72a, Leo74, Lwi75, Man70b, MEA76, Mar77, Mic75, Moh79, Mol77, MB73, MG78, Mui74, Mul77, NP73, O’N79, Olk72, Pal72, PP75, PK79, PG69, PG70, Pre71a, Pre74d, RS70, Ram72, Ree72, Sam70, SFC72, Sch75, Sha70a, Sha70b, She71b, Shi78, Sie79, Sin70, SS77, SW73, Ste75b, Ste77, Ste79a, Sty70, TW77, Tar79b, Tar90, Tik71b, Upt73, Upt74, Van70, Van73a, Vit74, Wal75, Wat76, Wil75a, You70b, You71, You73b, AR81, Bin76, Boo75, HH71a, Har65]. distribution [Har79, McK74, Pet77, Pet78, Pre74c, Sch76, Sti74]. Distribution-Free [BMW78, BMW79, CR72, Koc79, O’N79, TW77, Van70, Van73a]. Distributions [BD76, Beh70, BTG77, BGKO66, BGKO71, BS73b, BS79, Cox70a, Dav70a, Dav76b, DS72, Daw77, FR72, Gol74, Gre79, Han75, HS79, HMM77, HM75b, HH79, Hol70, Joh72, JK73, Kal74, KM77, Koc79, KP78, KLC76, Law72, Law74, MS75, McL74, Mul70, Nic70, PT70, Ped79, Rod77, Sch78a, Sri70, Sub69, Sub70, SH76, Tar79a, TW77, Wat71, Wat74, Whi73b, You73a, DP73, DP74, Dav72, DS75, Saw78, SPB78, WA78, Cox70a]. Disturbance [Pie72]. Disturbances [HP79]. Diversity [GJ79a]. Divisible [Dey77, Fre76]. Domain [LS75b, Nic73, dJS73]. Dose [Fre70, HJ79, Rob78c]. Double [Rao73b, Tam78, Rao73a]. Doubly [Pre70]. Doubt [HHvdB74]. Downton [Den72]. Duration [Bar75a, San71]. Durbin [HM75a]. Dynamic [DH75b, GJ79b, Pie72, TW76]. Dynamic-Disturbance [Pie72].

Easy [Gue72]. Eberhardt [HH79]. Edgeworth [Dav76b, DT72, Ped79, Phi78]. Edgeworth-Saddlepoint [Ped79]. Editorial [Pea45]. Edward [Har10]. Effect [DH75a, DH77, Gei74, Hwa76, Mar71, McK78, TW76]. Effective [Fre70]. Effects [BGP77, Cha76, CH75, Fre73a, Hir78, LJ76, TA71]. Efficiencies [Luc77, dBP78]. Efficiency [Bes77, CS74, CS75, FH70a, GM78, Gea70, HMM77, Hey75, Hol72b, Hol72a, Jar77, Kal74, KM77, Kay79, Kno75b, Pea70, Pea72, Wat72, Wil75b, CS76]. Efficiency-Balanced [Wil75b]. Efficient [BS76a, DA70b, Nic76, PR77, Swe78]. Efron [BN78, Jam78, Rob78a, Spr78]. Electron [Bar65]. Element [Bis72a, Bis72b]. elementary [Sär71]. Elements [Sis72]. Empirical [BM72, CK68, CK70, Cop72, Cox75b, DP74, EM72, EK72, GK70, Geo71, GK71, Lai78, ML74a, Pre71c, RK75, Ste79a, Wei74, DP73]. EPF [BD77]. Epidemic [Bar75a, Den72, Kry72, Kry74, McN72]. Epidemics [Aba74, Dan72]. Epidemiological [Cla78]. Equal [Ahs71, GL70, ZP77]. Equality [Bho78, Bho79, Cho77, Cor72, KP78, Lay74, LGGB79, Mar76a, MG73, Mor73b, Nai75, SA76, TW77, Wol76]. Equally [GNP73]. Equation [BM78b, MRR^{+}53, Atk74, Dav79]. Equations [God76a]. Equatorial [AS72]. Equiradial [Cor72]. Equivalence [Ham70, Hil78, Whi73a]. Ergodic [Han72]. Ernst [Ken72, Ken71]. erratum [Boo75]. Error [Bai79, Blo72, CH75, Fed74, GW78, Har74, Hea77, Jan75, KM73, McL74, McL76, RK75, She71a, Boo75, Woo78]. Errors [BW74, BM78b, CS79b, DH75a, DH77, DL72, Mor75a, Pie71, PK79, Pre74a, Saw70, Sin79, You72b, CS82, Mor75b, Sär71]. Established [Ray71]. Estimate [Leo74, Atk70]. Estimated [Dur75, Gar72a, Hos79, RW75]. Estimates [BGR76, Blo78, BW78, Cor72, FH70b, Ham70, MT77, PW75, Sin79, Smi73, Sow71, Sow72, TN79, Wed76, MH78]. Estimating [BW71, CH74, Dav75, ET76, Fre73a, FR72, God76a, Gra79, Han72, HT73, Hin76b, Kry72, LS77c, Sch75, Sch76, Seb70, Seb71, Sil78a, Sti70, Tik71a, Wat71, Dav76a]. Estimation [Adl77, Ahs71, Ali77, Ali79, Ano81, BAS70, BCK78, Bas72, Bec74, Bec79, BM75, Bes77, Bha76, Bin75, BL76, Bli70, Box71, Bre79, BP74, Bru78, BO78, CSW76, CT77, Cox76c, CS79b, DB69, DB70, DGS74, DGK75, Dig75, Dig77, DL72, ESF73, Far78, Far79, Fel74, Fre70, Fre72, FV79, Gar70, GJ79a, GW73, God78, GLGH79, Gou77, HT71, HP79, Har69, Har70, Hea77, Hey75, Hin78d, Jan75, Jon74a, Jon74b, Kei74a, Kei74b, Koh77, KM73, Lee79, Leo76, Lev72, Lew75, Lin74, Mac74a, ML74a, MW77, Mee79, MN74, MS76, MR70, Moh79, Mur77, Mur79, Nic73, Nic76, Nic77, Nic70, O’H73, OL76, O’N79, PHL75, PDT77, Pie71, Pie72, Pre74b, Pre70, Pru76, Que56, RS70, RS78a, Ros79, Rud74, Sat77, Sch78a, Sha77, Sie78, SE75, Sow72]. Estimation [Spr73, Sti71, Thi74, Ton71, Tua79, Wal71, dJS73, Ali78, Bin76, CS82, CH72, Mor75b]. Estimator [Bes74, DA70b, EH78b, EK72, God77, Hea78, Hol73, Pre71a, Pre79b, Ree72, See75, Tam78, Pre71b]. Estimators [Chm76, CK68, CK70, DC78, Fea77, Geo71, GK71, HBA70, Has71, Hei78, JOZD70, McK78, Mil73, PR77, RR71, RS78a, RK75, RF70, RW75, Tik70, Ton71, Hut71, McL77]. Evaluating [Web70]. Evaluation [Cox75a, Geo71]. Event [Mes73, RA79]. Exact [Alt71b, AH73, Ans79, Dan72, Dav70a, Dav72, HMM77, Mar79a, MK79, New74, NH79, PK78, PG69, PG70]. Example [See75]. Exceeding [Kap70]. Exchangeable [BTG77]. exciting [Haw71]. Existence [HPF70, HPF71, Wed76]. Expansion [Dom72, Ped79, DT72]. Expansions [Cra72]. Expectation [DS72, Mor75a]. Expectations [Daw73]. Expected [AL76, EH78b]. Experiment [Bro74c, Efr71]. Experimental [BW74, Box71, BP76, DGP77, Fre73a, PCdCM74, Woo78]. Experimentation [DS73]. Experiments [And71b, AH73, Atk72a, AF75a, AF75b, Ber75, Cot79, Dow76, Feu79, GL70, HPF70, HPF71, HJ79, HP74, Hol78, Joh73, LJ76, LB77, Pea71, PPK73, Vuc77, Wei78, Atk70, Mor75b, HW72]. Explanatory [Pre73]. Exploratory [Goo74]. Exponential [Bli70, Blo73, Bro74a, BY77, FH70b, JOZD70, Lwi75, Mac74a, MG72, MM76, Pre73, SFC72, Pet77]. Exponentiality [Dur75, GW78]. Extended [Bes73, JK73, Kry74, MDS74]. Extension [Aka79, BP76, DB69, DB70, EM72, Whi73a]. Extensions [Joh74b]. Extinction [Sti71]. Extraneous [Hem77]. Extrapolation [HC72]. Extremal [Ize76]. Extreme [AR72, Law74, Mui74, Ste77, AR81].

F [Boo75]. Facilitate [JK71a, JK71b, Joh74b]. Factor [Cha77, CS74, ER75, Jar77, LM73, Pat70]. Factorial [BGP77, Bai77, Cot79, Joh73, KP76, LJ76, LB77]. Factorials [BEH76]. Factorization [JW71, God76b]. Factors [GD74, JW71, Mes73]. Failure [BD76, MG72, PB78]. Familial [Cla78]. Families [Alt76, De 77, HPF70, HPF71, KMN75, dBP77, Wal67, Wal70a]. Family [Atk79, Bli70, JOZD70, Lwi75, Tar79a, Saw78]. Farlie [SPB78]. Fast [MRR^{+}53]. Fedorov [HW72]. Field [DGP77]. Fields [Bes77]. Finch [Hea79, Kem79]. Finding [Hol71b]. Fingerprint [DR73]. Finite [AN79, Bli73, CSW76, Cox75b, De 77, FGS74, Han72, Hol71b, dBP77, Roy70, Roy76, Wyn76, DTN77]. First [MV76, Phi78, Ree72, Rud74, Sow72]. First-Order [Phi78, Ree72, Sow72, Rud74]. Fisher [Pfa74a, Dal78, EH78b, Kem75, MEA76, MS70, Nai74, Pfa74b, Sch78a, Ste75b, Sti73, SS73]. Fit [Ait75, Dah71, DG72, FS71, Gas75, HP72, Hol72b, Hol72a, Koe79, LB78, Moo73, PK79, SFC72, Sla73, Sri70, Ste77, Ste79a, Tar79b, Tar90]. Fitting [Aka79, BM78b, Buk72, Cha73a, Cha73b, JK71a, JK71b, PDT78, dJ76, Joh74b]. Fixed [DS68, Gar70, Kap70, NS72, PS73, DS76a]. Flash [EB78]. Following [Mee79, Sie78]. Forcing [Efr71]. Forensic [Lin77, Seh78]. Forest [Cox76c]. forgotten [Mac77a]. Form [BDJS78, God75, Goo69, Goo70, Kha78, Rob78b, Sha70b, Kem75]. Forms [HP72, JW79, Mul70, Sty70, JW81]. Formula [Koe79]. Formulae [Lee71]. Foundations [Cox72a, Sto72]. Four [LGGB79, Pat70, PP75]. Fraction [Mac70]. Fractions [LJ76]. Francia [Sar75]. Free [BMW78, BMW79, CR72, Koc79, O’N79, TW77, Van70, Van73a]. Freedom [Dav75, GL70, Pre74a, Sie79, Dav76a]. Frequency [BL76, Eng74, Eng75, Mes73, MV76, Nic73, Joh74b]. Frequentist [Woo76]. Front [Ano70d, Ano70e, Ano70f, Ano71d, Ano71e, Ano71f, Ano72d, Ano72e, Ano72f, Ano73d, Ano73e, Ano73f, Ano74d, Ano74e, Ano74f, Ano75d, Ano75e, Ano75f, Ano76d, Ano76e, Ano76f, Ano77d, Ano77e, Ano77f, Ano78d, Ano78e, Ano78f, Ano79d, Ano79e, Ano79f]. Function [DLB72, Dav73, GM78, Lac75, Lev72, LS76a, LB79, Mac74a, Mor75a, New74, NH79, PK78, RT73, Roy77, Sch75, Ste79a, TW79, Wic73, Dav79, Sch76]. Functional [DL72, Dol76, DC77]. Functionals [Wat71]. Functions [Bar65, Bro77b, Cox71, Gra79, HW78, Jon74a, KS72, Mui74, Mur77, Mur79, Roy76, Spj72, Wed74]. Fundamental [Som71a, Som76, HS72a]. Further [Kay79, LO78, Dav79]. Future [AR72, AR81].

G [Cox70a]. Galton [CH72, Qui70, Sti70]. Galton-Watson [CH72]. Gamma [BB78, Ize76, JK72, Pre74b, Pre74d, You71, Pet78]. Gaps [Cre78, Cre79]. Gart [VB78]. Gauges [HS78a]. Gauss [SH76, Wed74]. Gauss-Newton [Wed74]. Gaussian [Lou76, Aka73, BM75, Bes77, God77, God78, Lou75, PDT77]. Geary [DR74, HM75a, SBL79]. General [Afo72, And76a, DL72, Jör70, Koe79, Pea70, Pea72, Whi73a, Wil70a, Wor75, Wyn75, Dut75, MH78]. Generalization [BD77, HM76, Som71a, Som76, Nai74]. Generalized [Bre70, Bro78, CSW76, Chi74, Dav70a, DF79, GSW75, HS72b, JH78, Joh73, Kem75, Lat77, Lee74, Pet78, SF79, Tik71b, Wed76, Dav72, Wed74]. Generalizing [Gra79]. Generated [FH70a, HS78a]. Generation [Atk79]. genetic [Mar70b]. Geometric [Eng75, Her76, ST73, Vit74]. Geometrical [Tit75]. Given [You70a, You72a]. Goal [Chi74, SE75]. Godambe [Cox72a, Sto72]. Goodness [Ait75, Dah71, FS71, Gas75, HP72, Hol72b, Hol72a, Koe79, Moo73, PK79, SFC72, Sla73, Sri70, Ste77, Tar79b, Tar90]. Goodness-Of-Fit [Dah71, FS71, HP72, Gas75, Hol72b, Hol72a, Koe79, Moo73, SFC72, Sla73]. Gram [DT72]. Gram-Charlier [DT72]. Graph [Gra79]. Graphic [Gab71]. Graphical [Cox79, DS76b, GL70]. Group [Dey77, Fre76, HT71, HT73, Hwa76, Poc77, SE75, Spj72]. Grouped [Alt76]. Groupings [GL70]. Groups [Cox76a, Cox77, HS78c, LS75a, SY74, Sla73, DC77]. Growth [BCK78, DG76, Fea75, Fea77, Hin79]. Guide [Rod77]. Gumbel [SPB78].

Haenszel [LSG83, BK74, LSG79]. Half [Sha70a]. Hannan [Koh77]. Harmonic [Ran74, Wal71]. Hastings [Hit03]. Having [BT74]. Hazard [Koc79]. Heavy [LS77b]. Henshaw [HM75a]. Hermite [Phi71]. Heteroscedasticity [HM75a]. Heuristic [Fel74]. High [Cre76, Cre78]. High-Order [Cre76, Cre78]. Higher [Cre79, Zen75]. Higher-Order [Zen75]. Highly [PR77]. Hinkley [BN78, Jam78, Rob78a, Spr78]. Histograms [Leo73, Sco79]. historical [HS72a, Sti74]. History [Cra72, Hit03, Rab70, She70, She71b, Sti75, DP72, Goo79a, HS72a, Ken71, Ken72, Mac77a, Pla72, Sär71, Sea77, Sti73, Sti74]. Hollander [LS79, Sch78b]. Homogeneity [Fuj79, Mor73a, SHA78a, Vit74, Wis68, Wis90]. Homoscedasticity [Kor72]. Hopf [Atk74]. Horizon [UL76]. Hotelling [Dav72, Dav70a, HS72b, McK74, Tik71b]. Household [Bec79]. Hurst [AL76]. Hybrid [Ton71]. Hypotheses [Cha74a, De 77, Hay79, HS78a, JMM75, Joa72, Mor70b, MDS74, MH71, dBP77, Sha75, Wal67, Wal70a, Joa75, Pre71b]. Hypothesis [Dav77, EB78, Hay75, NP73, Saw70, Sär71, Saw78].

Identifiable [Goo74]. Identification [Aka73, BGP77, Bri77, Han79].

Identities [Jos73]. Identity [Phi71]. II [Goo79a, JK71b]. III [DW71]. Implications [SD72]. Importance [UL76]. Imprecise [AL79]. Improper [SD72]. Improved [FS71, McL77, Wil76, Wil78, Fus73]. Improvements [BAS70]. Improving [Hin78d]. Inadequacy [BO75, JM78]. Inadequate [Atk72a]. Incidence [Cla78, PP79, Rog77, Rog78]. Incident [Yan78, Yan80]. Inclusion [Sin78a]. Incomplete [Bho79, CS74, CS75, Ekb76b, Fre76, HM72, JH78, LS74, Lit76, Mur79, Nai75, PW76, Rob70, Sin78b, Tit77, Blo76, CS76, Fie72, Pre79a]. Incompletely [Gol74, Sri70, Wat74]. Increased [Cha74b]. Independence [Dar71, DR73, GD74, MK79, Mee78, Sim77]. Independent [AL76, DH75a, DH77, Gai74, GW75, Geb70, RW75, Wal71]. Index [GJ79a, GJ79b]. Indicant [AB70]. Indicator [OW78]. Individual [RS78a]. Individuals [Alt76]. Induced [HS79]. Inefficiency [BW75]. Inequality [The75, Wyn77, HS72a, The76]. Infection [Kry72]. Infectious [BAS70, Bec72, Gou77]. Inference [BB78, Coo79, Cox75b, Cox75c, DG76, FR79, Gol74, HBA70, HS79, Har74, Hin70b, HH70, HH71b, Hin71, Hin77a, Hin78b, Hin78e, Hin79, HM72, Joh74a, Lit76, Mac70, MEA76, PP78, Rob78b, Rub76, Saw70, Sha76a, Smi75, SD72, Tam78, Wat74, Wil77a, Atk74, Hin78c, Sha78b, Smi76, Cox72a, Cox72b, Sto72]. influential [Har10]. Information [Ano70j, Ano71j, Ano72j, Ano73j, Ano74j, Ano75g, Ano76g, Ano77g, Ano78g, Ano79g, EH78b, Hem77, HM72, Oak77, PT71, Shi76, Sil78b]. Initial [GM78, Han72]. Initially [Nea70]. Inquadmissible [See75]. Integer [Dar71]. Integer-Valued [Dar71]. Integral [BT74, Dut73, KC76, Web70, Atk74, Boh73, Boo75, Dut75]. Integrals [Dut73, Olk72, Sis76, Wyn75, Wyn77]. Integrated [God75, Hea77, Roy77, Wic73]. Inter [PT71]. Inter-Block [PT71]. Interaction [BM75, Dar74, Gar72b, Gar73, HWB^{+}77, Hir78, McK78, RS78b, VB78]. Interactions [Cot79, Zen75]. Intercept [LS77c]. Interclass [Ros79]. Internal [GL70]. Internally [Mar77]. Interpolation [HC72]. Interpretation [Joh70]. Interpreting [HW78]. Intersecting [BW71]. Interval [Bre79, Gar70, Law72, Ton71, Vit73]. Intervals [AKH70, BF72, Buh77, DG72, FH70a, Jar79, Law74, LMH67, LMH71, Mar79a, Odé73, Rob75, Rob77, Rob78c, Sha70c, Spj72, Tar79a, Wal70b, Lew72]. Intraclass [Alt71b, Rot79]. Invariance [Cor77]. Invariant [Daw77, KMN75]. Invariants [Pee78]. Inverse [And76a, Bes74, DS72, MS76, Odé73, Sat77, Sha73, SS77, Sis72, You71]. Inversion [Dav73, Mar77]. Invertibility [And75d]. Investigation [You72b]. Involving [Jos73, MH71]. Items [SY74]. Iterated [BT74]. Iterative [MT77, Wor75].

J [HW72]. J. [Bar75b, BN75, Bir75, HS72a, McL75, Sch78b]. Jackknife [Efr79, Far78, GSW75, Hin77b, Hin78d, Jon74a, Mil74, Sha76b, Tho76, Hin78a].

Jenkins [McL77]. Johnson [Cox70a]. Joint [O’H76, Spj72]. Jonckheere [Ode72]. Joshi [Cox70a]. Judges [HS78c]. Judging [Har10].

Kalbfleisch [Bar75b, BN75, Bir75, McL75]. Kendall [Bes73, Joh79, Sim77, Woo70]. Kernel [AA76, Tit77]. Key [Bro73, Bro77b]. Keys [GP75]. Kiefer [Joh70]. Klimko [HW72]. Know [ET76]. Known [Hin77a, Hin78b, JK71b, MV76, Hin78c]. Kolmogorov [Dur75, MD71, MM76, Woo78, WA78]. Kolmogorov-Smirnov [Woo78, WA78]. Kotz [Cox70a]. Krawtchouk [SH76]. Kruskal [Bre70]. Kuiper [BS73a]. Kurtosis [Bow73, Mar70a].

L [Cox70a]. Labelled [Mac70]. Lack [LB78]. Lagrangian [ST74]. Lambda [JK73]. Lancaster [Zen75]. Laplace [Mar77, Sti73, Sti75]. Large [D’A71, Daw73, JMM75, MW77, Pat78, Rob74, Sla73, Spr73, Whi78, WA78, vK73, Tuk58, Woo78]. Large-Sample [JMM75, Rob74, WA78, Woo78]. Largest [Ofo73, Ofo75, HH71a]. Latent [BAS70, Gil79, Goo74, Gou77, Mui74]. Lattice [BM75, Mar79b]. Law [DF79, She70]. Laws [PHL75, She71b]. Layout [Spj72]. Leading [Cro74]. Leads [Fea77]. Learning [Pru76]. Least [AL71, BW75, Bro78, Cor72, DW71, Gea70, Kno75b, Mil76, Pie71, Pie72, Wat72, Pla72]. Lecture [Efr79]. Left [Hyd77, MV76]. Lehmann [Bro74b, Pet72, You73a, ZP77]. Lehmann-Type [Pet72]. Length [BE72, Wil75a]. Lengths [Col72]. Level [BL76, Ked78, Lin74]. Levels [Cox75a, Pat70, DTN77]. Levin [Wal75]. Lewis [Cor77]. Life [BD76, BD78, Cla78, HP75, KP73, LS79, Mor78, Tar75, HP76, HP80]. Light [LS77b]. Likelihood [Aka73, Ans79, BW78, Cop75, CD78, Cox75d, Dav71, DC78, DL72, Edw73, EH78b, FK74, God76a, God77, God78, HBA70, Ham70, HP79, Hay75, Hin78e, Hin79, HS75, KLC76, Lee72a, Lev72, LB79, MT77, MB73, NP73, New74, NH79, Pee71, PK78, Pre74b, Pru76, Ree72, Ros79, Roy76, Sha75, Wed76, Wil76, Woo79, deC72, Dav79, Wed74, Wil78]. Likelihood-Ratio [Lee72a]. Likelihoods [Cox76b, Cro74, KP73, Spr73]. Limit [AR72, Mat70, SW73, AR81]. Limits [BMW78, BMW79, Hin77b, Hin78a, Van70, Van73a]. Line [Gar72a, LS77c, Ram79]. Linear [Bec70, BM78b, BF72, Bro74c, Bro77a, Bro78, BJ79, CK68, CK70, Cox75c, Cox76a, Cox77, DH75a, DH77, DC78, Edw73, FL76, GM78, GW73, GK71, Hil78, JMM75, Jen79, Jos72, KMN75, Krz79, Lac75, Lee74, LS76a, LO78, Mar71, Mar76b, Mor75a, MDS74, Nic73, Nic76, Nic77, Nic70, NS79, Odé73, Pre78, Pre83, Pru76, Rob74, Ros73, Roy70, RW75, Saw70, Shi77, Spj72, Wed76, Atk72b, DC77, MH78, Wed74, TA71]. Linear-Circular [LO78, Mar76b]. Lines [BW71, Col72]. Literature [Rab70]. Local [BR78, Hay75]. Location [DTL76, HM75b, Hin78e, KMN75, Law72, Mar79a, Moh79]. Location-Scale [DTL76]. Log [Cro74, FL76, JW79, LS76a, Moh79, Pre74b, JW81]. Logarithmic [Kem75]. Logarithms [Cre76]. Logistic [And72, And75a, And79, AKH70, Far79, Law72, McC77, MG78, PP79, RS70, Ste79a, Atk72b]. Look [Aka78, Efr79, Har10]. Lower [BMW78, BMW79, CC77]. Lumped [Sha76a, Sha78b].

M [Goo79a, HW72]. M. [Sch78b]. Machines [MRR^{+}53]. Main [LJ76]. Mantel [LSG83, BK74, LSG79]. Mantel-Haenszel [BK74]. Many [ET76]. Marginal [Ahs71, KP73, Lev72, O’H76, Spr75, Pla77]. Marginals [Gar70]. Markov [Sha78b, Bas72, Dev73, Han72, Has70, Ked76, LSZ78, Mat70, MW77, Pes73, Sha76a, Sha75]. Markovian [Pre71c]. Matched [Alt71a, HP74, Hol78, Cop73]. Matching [Car77]. Mathematical [Cra72, CH74, Mac77a]. Matrices [Ait69, Ait71, Gab71, Gil79, Lay74, Sil78b, SS77, Ste75a, Thi74]. Matrix [And76a, Cha76, Mui74, NP73, Sha73, Sis72, Whi73b]. Matter [Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. Maximal [Pet72]. Maximum [AB70, Aka73, AR72, BW78, CS75, DPF74a, DPF74b, DC78, DL72, EH78b, FK74, God77, God78, HBA70, Ham70, HP79, HS75, Mes73, MT77, Pre74b, Pru76, Ree72, Ros79, Wed76, deC72, AR81, BT73, CS76]. May [DGS74]. Mean [AAK72, Ano48a, Ano48b, Bai79, Bin75, BL76, BGKO66, BGKO71, BMW78, BMW79, DA70b, FH70b, Gea36, Hea78, Hin77a, Hin78b, HP75, Jan75, Jos72, KM73, Lac75, LS79, Leo74, Man70a, Mil73, MB73, Ofo73, Ofo75, Pre79b, Ran74, SFC72, Sha70a, SBL79, Sie77, Sin79, Bin76, Hin78c, HP76, HP80]. Means [Ahs71, Bho78, Chi74, Ekb76a, Ekb76b, Jon74a, LS75b, LS74, Lin74, Lit76, Lou75, Mar76a, MG73, Mor73b, Nai75, Sis76, Spj72, Tik71a, Ton71, Wil70b, Wil77a, dJS73, Lou76]. Measure [Gor79a, Gor79b, LB78, Rot79, Tjø78, Wal75]. Measurement [GW78, You72b]. Measures [Hil78, Mar70a, Moh79, Rob78b, Sil78b]. Median [Fre70, Van70]. Meijer [Dav79]. Mendenhall [Wil75a]. Menten [BM78b]. Method [AA76, AH73, BP76, CL76, Cox76c, Dom72, EM72, Fre76, GLGH79, Gra79, Jör70, Leo73, LS77c, Ran74, Tit77, Web70, Wor75, Pla72, Wed74]. Methods [Cop72, Dav75, DPA77, Dig75, Efr79, Feu79, FR72, Has70, Lai78, Leo72, MT77, Nic77, Poc77, Whi78, Dav76a]. Metropolis [BBLS00, BS00, Hit03]. Michaelis [BM78b]. Migration [Ren72]. Minimal [Afo72]. Minimax [JOZD70, SS75, DS75]. Minimizing [Rea73]. Minimum [Aka79, Blo78, Jan75, Kno75b, Thi74]. Mining [Jar79]. Mises [Hil77, MEA76, Sch78a, Hil76, HR77, KG76, MS75, Pet76a, PS76, Pet77, Pet78, SG74, Upt73, ZP77]. Mises-Lehmann [ZP77]. Missing [Bho78, Ekb76a, MG73, MT77, MB73, Mor73b, Rub76]. Mitoses [Mac70]. Mixed [And75d, Ans79, FR72, New74, Ped79]. Mixture [Bec70, Beh70, BS73c, Cha76, MS75, Whi73b]. Mixtures [Cox71, Krz79, PP78]. Modal [Ste75b]. Modality [Hol70]. Model [Aka79, Atk78, Bec70, BD77, Bis72a, Bis72b, BO75, Blo73, Box71, Bro78, Cla78, CK68, CK70, CS73, Cox75c, DB69, DB70, DN79, DF79, DG76, Eng75, Far77, Fea77, Gei74, God79, Hin79, JM78, KP73, KR76, Kry74, Lee74, Mar71, McC77, MI77, Pie71, Pol75, Pre74b, Shi76, Shi77, Sow71, Sow72, Str75, MH78, TA71]. Models [Aka73, Ali77, And71b, AN79, Atk72a, AF75a, AF75b, Bai79, Bro73, Cha73a, Cha73b, CH74, Daw77, DC78, Eng74, Far79, FL76, Goo74, Goo79b, HW78, HP79, Hil78, Hos79, Jen79, KM79, Kay79, LW78, LSZ78, Lay75, LB78, LB79, McC78, MW77, Nic73, Nic76, Nic77, NH79, dBP78, PK78, Pie72, PK79, Pre75, PB78, PP79, Pru76, Ram79, Roy70, Smi73, Tua79, Wed76, Whi73a, WL77, Ali78, Atk72b, Cop73, Wed74]. Moderate [D’A71]. Modes [MS75, O’H76]. Modification [Leo74]. Modifications [Pre74a]. Modified [DC78, God78, Goo73, Gre79, HM75b, JW79, KP78, PS76, Ray71, JW81]. Moivre [DP72]. Moment [Adl77, Lin74, Qui70, RF70, You71]. Moments [Ano48a, FGS74, GP70, Gea36, Jen79, Koe79, MR70, MV76, Saw58a, Saw58b, Saw79, Sis72, SH76]. Monotone [Geb70]. Monotonically [Wil77a]. Monotonicity [Olk72]. Monte [Bar65, DL73, Has70, Hut71, Lay74, LDG75, Mor75b, Pes73, Tik70, Yat72]. Monte-Carlo [DL73, Yat72]. Mood [HM75b]. Moran [SS70]. Morgenstern [SPB78]. Mortal [Bec74]. Most [Mul77, RA79, Har10]. Movement [Yan76]. Moving [Aka73, Ali77, And75b, And75d, And76a, And76b, Ans79, AN79, Bha76, CT77, DPF74a, DPF74b, God75, God77, God78, Han79, HP79, Hos79, Koh77, LB79, McC73, New74, NH79, PK78, PDT78, Pie71, Roy77, Sha73, Tua79, Wic73, Ali78, And75c]. MR [AR81, BR81, CS82, Har79, HP80, JW81]. MR0497517 [Pre83]. Multi [ER75]. Multi-Factor [ER75]. Multiarmed [GJ79b]. Multidimensional [Goo73]. Multifactor [Atk73]. Multinomial [Gai74, GW75, Olk72, RA79, Sto74, SH76, You71]. Multiparameter [Leo76]. Multiple [BT77, BP76, FL76, Lee72b, Mac74b, MT67, McK77, Sli70, Spj77a, Spj78, TP74, You72b, You74, Fie72, MT72, Sea77]. Multiple-Stage [You72b, You74]. Multiplicative [Dar74, Goo79b]. Multirespnse [GL70]. Multitrial [DPA77]. Multitype [Qui70]. Multivariate [AA76, And79, BT74, BP75, BP74, CS78, DB69, DB70, DB71, Dav76b, Dut73, FK72, HM76, Joh71, JK73, Kor72, KLC76, Lee71, Lit76, Mar70a, Mar71, MB73, MDS74, Mur79, NH79, PK78, PDT78, Saw70, Tar79a, The75, Tit77, Web70, Boh73, Dal78, Dut75, SS73, The76]. Mutation [CH74]. mutually [Haw71].

N [Boo75, Cox70a, DC77]. Naive [Dev73]. Napoleonic [Sti75]. Nature [And75d, Cox76b]. Nearness [Rob78b]. Neat [Hai71]. Necessary [Kha78]. Necessity [Rob78b]. Negative [Bin75, Bis72a, Bis72b, You70b, You73b, Bin76]. Nested [SD79]. Newton [Wed74]. Neyman [Wal70b]. Neyman-Shortest [Wal70b]. Nicholas [BBLS00]. Nile [Cob78]. no [BR81, CS76, CS82, Dav76a, DC77, HM75b, Hin78c, HP80, JW81, LS77a, Lou76, Pre83, Sha78b, The76, Wil78]. Noisy [Dev73]. Non [JP69, PT70, Sub69, Sub70, Has74, Yat72]. non-additivity [Yat72]. Non-Central [JP69, PT70]. Non-Normal [Sub69, Sub70]. non-normality [Has74]. Nonadditivity [Pat70]. Noncentral [Han75, JP70, Sie79]. Noncircular [Phi78, Ram72]. Nondeterministic [Bha76]. Nonergodic [FR79]. Nonformation [BN76]. Nonhomogeneous [LS76a]. Nonlinear [Box71, Bri77, Cha73a, Cha73b, DL72, Whi73a]. Nonnegative [BMW78, BMW79, Dar71]. Nonnormal [Pre74a, Saw70, Sin70, Sin72, Wat76, You74, Yat72]. Nonnormality [Mar71]. Nonnull [Joh79]. Nonorthogonal [ER77, JW71]. Nonparametric [Bre79, CH72, Dow76, DTL76, EK72, GK70, GG71, Hol71a, Mil73, Mor78, Rot79, Sim77, Sli70]. Nonrandomized [Wal70b]. Nonrandomness [HH79]. Nonstandard [Cha74a, Mol77]. Nonstationary [Mac74a]. Nonsymmetric [LS76b, LS77a]. Nonzero [McK78]. Normal [Ahs71, AAK72, AL76, Ano48a, BT74, Beh70, BGKO66, BGKO71, BS73c, Cha76, Chi74, DLB72, DA70b, Dut73, FR72, Gea36, GNP73, Han75, HS79, Hea78, Hin69, Hin70a, Hin77a, Hin78b, Joh72, Kap70, KC76, KLC76, LC76, Leo74, Man70a, Mar76a, Moh79, MB73, Mur79, Ofo75, PR77, Pre70, Pre71a, Pre79b, Sam70, Saw58a, Saw58b, Saw79, Sha70c, She71a, Shi78, Sie77, ST71, Spr73, Sty70, Sub69, Sub70, Tik70, Tik71a, Ton71, TW79, Web70, Wil77a, DP72, Hin78c, Ofo73, She70]. Normality [Bow73, BS75, CS78, D’A70a, D’A71, D’A72, DR74, Dye74, GO77, Hai71, Hol72b, Hol72a, LS76b, LS77b, PDB77, Pet76a, Pre76, Spi77, DP73, DP74, Has74, LS77a, Woo78]. Normalizing [Pre74c]. Normally [Car77]. Normed [Loy79]. not-quite [Tuk58]. Note [Aba74, And71a, CT77, Cor72, Cox71, Cox75a, Cox75b, Cox75c, Cox76b, Cox79, Cro74, De 77, Dig77, Dom72, Dye74, Eng75, Han79, HM75a, HMM77, Hil78, Hol71b, Jar79, Joh70, Jon74b, KR76, Koh77, Lee74, Loy79, Mar79a, Moo73, Mor70a, Mur77, Pea45, dBP77, Qui70, SS70, Tik71b, Wal70b, Zen75, And75c, BS73a, DP72, HS72a, Sti74]. Notes [BS73b, JK71a, PT70, Que56, Sty70]. Nuisance [Dav77, Hem77]. Null [D’A70a, DT73, LS79, Mul77]. Number [ET76, Gue72, Ked78, Man70b, Mes73, Sla73, dC71]. Numbers [Geo71, Luc77]. Numerical [Cha73a, Cha73b, Dav73, Fre73b].

Obituary [BBLS00]. Observational [Mes73]. Observations [AR72, Bho78, Bro74a, Cla76, CCS70, Daw73, DGS74, DL72, EM72, Far77, Gut71, MB73, SY74, ST71, AR81]. Observed [BL76, EH78b]. Obtaining [MT77, Som71b]. Occupied [Str77]. Odds [Cla74, Cla76, Gar70, McK78]. Omnibus [BS75, D’A71]. One [BF72, CR72, Ekb76a, Geb70, Har69, Har70, KHB70a, MD71, MT67, Mor73b, Pus70, Ran74, Shi78, Sli70, Smi73, Spj72, Tar79a, Van73a, Wyn75, Har10, MT72, MH78]. One- [MT67, MT72]. One-Parameter [Tar79a]. One-Sample [CR72, MD71]. One-Sided [BF72, Shi78, Sli70, Van73a, Wyn75]. One-Way [Har69, Har70, KHB70a, Ran74, Smi73, Spj72]. Only [Dav77, Har74]. Operator [JW71]. Optimal [AF75b, Bel77, Bel79, Ber74, Bro72, Bro77a, CH75, CCS70, Cre79, Fre70, Gla78, GK71, Gut71, Hem77, HW72, HC72, Hil78, HJ79, JMM75, Joh71, Kie75, Lay75, Lou75, Lou76, Mor70b, Rob78c, Sco79, SY74, ST73, ST74, Sil78b, ST71, Tit75]. Optimality [OW78, Tit75]. Optimum [God76a, MG72, New71, Pes73, Pes75, Pes76]. Order [Atk73, BD77, Cor72, Cre76, Cre78, Cre79, GNP73, HS79, HP72, Jos73, KMN75, Ked76, McK78, MN74, Pee78, Pes75, Phi78, Pre74d, Ree72, Shi76, Sin72, Sow72, Ste75a, Sub69, Sub70, You70a, You70b, You71, You72a, You73b, Zen75, Pes76, Rud74]. Ordered [Cha74b, Cla74, Cla76, Dav70b, Fuj79, Goo79b, Hin72, Hir78, MPG76, Mar76a, McC77, McC78, Nel77, Sis76, Spj77b, Ton71, Wil77a, Ode72]. Ordering [Spj77b]. ordinates [Boo75]. Orientation [Dow72]. Orthogonal [Bru78, PPK73, Wor75]. Orthogonality [ER75, ER77]. Orthonormal [Pus70]. Other [Sti71]. Outlier [AB79, DGK75, Jos72]. Outliers [Bro75]. Overlapping [Mac72].

P [Cox70a, Cox72a, Sto72]. P. [Hea79, Kem79]. Packing [AH76]. Pair [HP74, Hol78, Mor79]. Paired [BEH76, DB69, DB70, DB71, DS73, Ekb76a, Ekb76b, FL76, FK74, Gue72, HS75, LB77, McC77, dC71, deC72]. Paired-Comparison [HS75]. Paper [Bar75b, BN75, BN78, Bir75, Hea79, Jam78, Kem79, McL75, Rob78a, Sch78b, SBL79, Spr78]. Parameter [Dav77, Hem77, Jon74b, Lwi75, ML74a, OL76, Pre71c, Ree72, Rud74, Sch78a, Sow71, Sow72, Tar79a]. Parameters [And75d, AKH70, BCK78, Box71, CT77, Dur75, Has71, Hea77, Hin78e, Hos79, Law72, Law74, Lev72, MW77, MN74, PW75, PHL75, PDT77, RS70, RS78a, Ros73, Sha75, Spj77b, Sti71, Tua79, Wal70b]. Parametric [Bec74, KS72, Kno75a, McC78, Pre75, Ram79]. Pareto [DF79]. Partial [Bil77, Cox75d, HW78, Joh79, Shi77]. Partially [Cox75c, LSZ78, OW78, Sin78b]. Particle [LS77c, Nic70, Wat71]. Particular [Bri77, Krz79, Mul70]. Partitioning [SY74, ST71]. Past [Geo71]. Path [Goo73]. Patient [UL76]. Patil [Cox70a]. Patterns [BGP77, Bai77, Bre70, Cor77, CL76, Dig77]. Pearson [BS73b, BS79, Dah71, DG72, Koe79, MV76, Sla73]. Penalties [GG71]. Percentage [AP70, BDJS78, BS79, Har64, HS72b, JP69, JP70, Lee72b, MK79, Wei74, HH71a, Har65, Har79]. Percentile [Buk72, Dav70b, Dav71, Has71, Sli70]. Percentiles [Chm76, ZP77]. Perception [EB78]. Performance [Kie75, Ray71, YD73]. Periodic [Mac74b, Tro79]. Periodicities [Mac77b]. Periodogram [Bri74]. Periods [BAS70, Gou77]. Personal [McL75]. Phase [MI77, New71, dJS73]. Pierce [DTN77]. pioneer [Mac77a]. Placed [HIR72]. Planar [Mar79b]. Plane [BR78]. Planning [Atk72a]. Plans [Wyn76]. Plantation [BH74]. Plasma [Bar65]. Play [NS72, PS73, SW70, Fus73]. Play-the-Winner [NS72, PS73, SW70, Fus73]. Plot [Rob70]. Plots [DFA77]. Plotting [DS76b, Sma78, Spj77a, Spj78]. Point [Dig77, Gar70, HH79, HH71b, Hin71, Lew75, Smi75, Haw71, HS72a, Hin70b, HH70, Smi76]. Points [AP70, BDJS78, BS79, BR78, Buk72, Cox76b, DT71, D’A72, Har64, HS72b, JP69, JP70, Lee72b, Luc77, MK79, Mor79, Wei74, BT73, HH71a, Har65, Har79]. Poisson [Cox75b, ESF73, GP70, Hol70, LS76a, Mac74a, Mac72, Pal72, PP78, Vit73, Wal70b]. Polychotomous [AB70]. Polynomial [BO75, Cox71, Mac74a]. Polynomials [Phi71, You77]. Population [Ano81, AG74, Bli73, BO78, DA70b, FGS74, Fre72, GLGH79, Hay79, Hol73, Hol71b, Juo70, Man70a, MW77, Ofo73, Ofo75, PP75, Pre70, Pre71a, Roy70, Roy76, Saw58a, Saw58b, Saw79, TW79, Wat76, Wyn76, Yan76, Yue74, Mar70b]. Populations [CSW76, Cha76, Chi74, CH74, Dav76b, DS76b, HSW72, Kap70, KLC76, LC76, Lou75, Ofo73, Ofo75, Sin70, Sin72, SW70, ST71, TY74, Ton71, You74, vK73, Fie72, Fus73, Lou76]. Portmanteau [DN79, DTN77]. positive [DT72]. Posterior [AAK72, Atk78, Daw73, Goo73, O’H76]. Potthoff [Lee74]. Power [Bow73, Cre78, DR74, DLB72, DN79, Hin75, JK72, Lee71, LDG75, MB73, Pet72, You70a, You72a, Hin76a, Ode72]. Powers [Mar76a, PDB77]. Practice [Fin79a]. Precise [Box71]. Precision [DGP77, Pea71, RS78a]. Preconceived [CD78]. Prediction [Ait75, Ali77, Ali79, Bai79, Blo72, Bro74c, Cam78, Dun76, Jan75, KMN75, Rob78c, Sto74, Vit73, Wat72, You77, Ali78]. Predictive [BD76, BD78, Gei74]. Predictor [Cha77]. Predictors [God75]. Preliminary [DA70b]. Preposterior [SS77]. Prescott [Loy79]. Presence [BR75, BW74, LGGB79, MG73, BR81]. Present [Dav77]. Press [Boo75]. Previous [CCS70, Gut71]. Principal [Gab71]. Prior [BTG77, BS76b, Gol74, Wat74, DS75]. Probabilistic [Gil79]. Probabilities [Ano81, Atk78, Buh77, BO78, CC77, Han72, Mic75, Sin78a, Som71b]. Probability [Ano48b, Ano70g, Ano70h, Ano70i, Ano71h, Ano71g, Ano71i, Ano72g, Ano72h, Ano72i, Ano73g, Ano73h, Ano73i, Ano74g, Ano74h, Ano74i, BE72, Cra72, DT71, D’A72, DS72, Dut73, Fel74, Gil79, GG71, Koc79, LC76, Mur77, Rab70, She70, She71b, Sis76, Sti71, Sti75, TW79, Wyn77, Boh73, Boo75, DP72, Dut75, Goo79a, HS72a, Ken71, Ken72, Mac77a, Pla72, Sär71, Sea77, Sti73, Sti74]. Probable [RA79]. Problem [Adl77, Ait77, Ait79, Bro74b, Cob78, DS68, GJ79b, Lat77, Lin77, Mar79b, MS70, NS72, Pfa74a, Pre71c, Seh78, Som71b, WAK77, Yan76, DS76a, Nai74, Pfa74b, Pre79a, SS73]. Problems [AB79, Ber74, Che71, Dom72, Fed74, Gol76, HHvdB74, HM72, MM76, SD72, Swi76, Atk74, Dal78]. Procedure [Aka78, Aka79, Bec79, Bil77, DGS74, HSW72, Ofo73, Ofo75, RA79, Vuc77, Wil70a, Fus73]. Procedures [BEH76, Col72, DB69, DB70, FV71, Koh77, Law72, MPG76, McK77, Nic73, Sow72, Swi76, TY74, Van73b, Wil77a, You72b, Dal78, DS75]. Process [And76a, Ans79, Bha76, CT77, Cox75b, ESF73, God77, Joh70, Ked78, Kei74a, Kei74b, Koh77, Lew75, Mac74a, New74, PW75, PDT77, Qui70, Roy77, Rud74, San71, Sha73, Vit73, Wic73, CH72, HS72a, Sti70]. Processes [Ali79, And75b, And75d, And76b, BS76a, BB78, Bec74, Bes72, BM75, Bri74, CS79a, DPF74a, DPF74b, FR79, God75, Hey75, LS76a, Mar79b, PDT78, Ren72, Sti71, Swe78, vK73, And75c, Haw71]. Product [SS77]. Products [Wil77b]. Programmes [Bec72]. Programming [DH75b]. Progressively [DF79]. Properties [BD77, Chm76, Joh79, LS75a, LSG79, LSG83, Mor73a, Nic70, Olk72, Pat78, Ste75a, Tho76, Woo76, BK74, Blo76]. Property [Blo78]. Proportional [GLGH79, Sin78a]. Proportions [Alt71a, Bec70, Kap70, New71]. Proschan [LS79]. Proton [Bar65]. Proton-Electron [Bar65]. Prove [Hai71]. Pseudolikelihood [Bes77]. Pseudoquadratic [God78]. Publication [Ano71g]. Publications [Ano70g, Ano70h, Ano70i, Ano71h, Ano71i, Ano72g, Ano72h, Ano72i, Ano73g, Ano73h, Ano73i, Ano74g, Ano74h, Ano74i].

Quadratic [And75a, BDJS78, Goo69, Goo70, HP72, Har69, Har70, HS75, Kha78, Mul70, O’H73, Sha70b, Sty70]. Quadrature [SH76]. Quality [GK70]. Quantal [AB70, Sow71, Sow72]. Quantitative [Pat70]. Quartiles [Van73a]. Quasi [DR73, Pes75, Wed74, Pes76]. Quasi- [Pes75, Pes76]. Quasi-Independence [DR73]. Quasi-likelihood [Wed74]. Quenouille [Pat73]. queueing [Atk74]. Quick [RS78b]. quite [Tuk58]. Quotient [Ize76].

R [DC77]. R. [SBL79]. Rabbinic [Rab70]. Radial [Bar65]. Radii [Sma78]. Random [AH76, Atk79, BMW78, BMW79, BP76, Col72, Coo79, DG72, DGP77, FGS74, Gei74, GNP73, Hin69, Hin70a, Kin72, Mor79, She71a, Smi75, SW73, Wei78, Wil77b, Yan78, Yan80, Hin70b, Smi76, TA71]. Randomization [Cop73]. Randomized [KHB70b, Woo78]. Randomly [HIR72, Ros73, KG76]. Randomness [AS72, BR78, Kno75a, Kno77]. Range [AL76, BGKO66, BGKO71, Bro74d, DLB72, Sin70, TP74, Lew72, Pre71b]. range-type [Pre71b]. Rank [Cro74, Joh79, JW79, MT67, MT72, Pet72, Pre78, You73a, JW81, Pet76b, Pre83]. Ranking [Chi74, FK74, HS75, Som71a, Som76, deC72]. Rankings [Gor79a, Gor79b, LS75a, Pre79a]. Rao [Fea77, Lee74]. Rare [Bec72, CS79b, CS82]. Rate [ESF73, Kry72, LS76a, Mac74a, Mar70b, McL74, McL76, Seb71]. Rates [CH74, Koc79, Seb70, Seb71]. Ratio [AP70, Ano48a, Buh77, Cla74, Cla76, CL76, Dav71, Gar70, Gea36, HM76, Hay75, Hin69, Hin70a, KLC76, Lee72a, McK78, MB73, NP73, Pee71, RR71, Sha75, Wil76, Woo79, BT73, Dav79, Hut71, Wil78]. Rational [PDT77]. Ratios [Dav70b, Mul70]. Recapture [Fre73b, Fre73c, Pol75, Fie72]. Reciprocals [SM72]. Reconciliation [BCK78]. Reconstructing [Dev73]. Recovery [PT71]. Recursive [And75d, Wil70a, dJ76]. Reduction [Lee74, Has74]. Reference [DR73, Hin78d, Hol78, Koc79, Krz79, MPG76, PCdCM74]. Regions [And76b, Atk73, DT72, Juo70]. Regression [AL71, Atk72a, AC77, Atk78, BO75, BF72, Bro72, Bro74c, Bro74a, Bro77a, BJ79, CSW76, CK68, CK70, CCS70, DB71, DH75a, DH77, DGK71, DW71, Edw73, Fed74, Gar72a, Gea70, Geb70, Gol76, Gut71, HP79, Jos72, KP73, Kay79, LM73, Lay75, LMH67, LMH71, Mil76, Mor78, Odé73, dBP78, Pie71, PK79, Ros73, Roy70, Spj77a, Tam78, Spj78]. Regressions [NS79]. Regular [Dey77, FR79, Fre76]. Regularity [BR78]. Related [Bri74, Cam78, MM76, Mor78, Wol76]. Relation [Cox76a, Cox77, Dan72, Dol76, Eni73, GSW75, Pat78, PP75, Rob74, Spr73, Yan78, Yan80, DC77]. Relations [Dol76, You71, DC77]. Relationship [DB71, DL72, PT70, TW76]. Relationships [Han75]. Relative [Gea70, BK74]. Reliability [MG72]. Remark [CD78, Gut71, Kha78, MG78]. Remarks [Hey75]. Remote [Fre73a]. Removal [GLGH79]. Repeated [Cop72, FH70a, Sie77, Woo79]. Replacement [Ber74]. Replicated [DL72]. Replications [Gue72, dC71]. Reply [EH78a, Fin79b, HS78b, Kal75a]. Reporting [Seb70, Seb71]. Representation [BT74, Bha76, Dut73, FL76, God77, Pus70]. Representations [Wil77b]. Requirement [Bow72]. Requirements [KHB70a, KHB70b]. Rescaled [AL76]. Research [HJ79]. Residual [AN79, Gea70, HP75, JW71, LS79, HP76, HP80]. Residuals [And71c, Bro75, CS71, Loy79, Wal71]. Resolvable [PW76, Sin78b]. Response [AB70, Cox71, Ekb76a, HJ79, KM73, Mor73b, Sow71, Sow72, Atk70]. Responses [Bho78, Bho79, Ekb76b]. Restraints [And75d]. Result [Wyn75]. Results [Arm70, CSW76, Cre78, Far79, KP76, Lee72a, LO78, RR71, Tro79, You73b, DP73, DP74, WA78]. Retrospective [Far79, PB78]. Returns [Seb70, Seb71]. Review [Cox70a, Cox72a, Cox72b, HW72, Mil74, Sto72]. Rho [Woo70]. Rhythmometry [Mar76b]. Ridge [Vuc77]. Ridge-Type [Vuc77]. Rietz [Efr79]. Right [Hyd77, Pre78, Pre83]. Risk [Hol78, McL74, Mes73, Rea73, Wal75, BK74]. Risks [TN79, Sea77]. Rival [AF75a, Hil78]. Robin [Dan69, Dan70]. Robust [AC77, BD75, Cox76c, DGK75, Dig75, Dig77, Lew75, MP78, Mar79a, DL73, MH78]. Robustness [Bro74d, FH70b, GW78, HM75a, Mar71, PP75, Sha70c]. Root [SHA78a]. Roots [Mui74, Wat76]. Rotatable [Cor72]. Roughness [GG71]. Round [Dan69, Dan70]. Round-Robin [Dan69, Dan70]. Routine [SD72]. Rows [SD79]. Roy [Lee74]. Rule [Dev73]. Rules [Ray71, WAK77]. Run [BE72]. Runs [Str77].

S [Cox70a, Cox72b]. Saddlepoint [Ped79, Phi78]. Sample [And72, AN79, AG74, Bli70, Bow72, Bro74b, CR72, Cha74b, D’A72, De 77, GK70, Gar72a, Ize76, JMM75, KHB70a, KHB70b, Kon78, Lat77, LDG75, LSG79, LSG83, MD71, Man70b, Mui74, NS72, Ofo73, Ofo75, Pat78, dBP77, PR77, Pol75, PS73, RR71, Rob74, Roy77, Saw58a, Saw58b, Saw79, SS75, Spr73, Swi76, Tua79, Upt73, Wat76, Whi78, Wic73, Yan78, Yan80, You70a, You72a, You73a, Yue74, BK74, Lew72, Ode72, Pet76b, Woo78, WA78, YD73, ZP77]. Samples [Ano48a, BGKO66, BGKO71, Bre70, D’A71, DT71, DLB72, GM78, Gea36, Geo71, Hea78, MN74, Mul77, Pet76a, Pre70, Pre79b, Saw58a, Saw58b, Saw79, Sin70, Sin78a, Tik70, Tik71a, VB78, Boo75, DTN77, Tuk58]. Sampling [BR75, Bel77, BTG77, Bel79, Ber74, Bli73, BS73b, Col72, CS79b, DPA77, FGS74, Has70, Hei78, LS75b, Mar79b, MS76, Mul70, NS72, New71, Pes73, PS73, Rao73b, Roy70, Roy76, Sat77, SD76, SW70, TY74, Tam78, Wyn76, dJS73, BR81, CS82, Rao73a]. Savage [KP78]. Scalar [Blo73]. Scale [DTL76, Hin78e, KMN75, Law72, MP78, MM76]. Scaling [HG76]. Scandinavian [Sär71]. Scarce [Nea70]. Schemes [BS76b, SD76]. school [Sär71]. Science [Lin77, Seh78]. scientists [Har10]. Scores [BY77, Dan69, Dan70, Pre70, Pre71a]. Screening [And71b, Cot79, You74]. Search [FK74, Pal71]. Seasonal [HW78]. Second [Atk73, Cor72, Gar72a, Ked76, McK78, Pee78, Pes75, Sis72, Pes76]. Second-Order [Cor72, Ked76, McK78, Pee78, Pes75, Pes76]. Selected [BD77, ZP77]. Selecting [Bro77b, GP75, LC76, Ofo73, Ofo75, RA79, SW70, TY74, TW79, Fus73]. Selection [And71a, Chi74, DS68, McK77, NS72, OW78, Shi76, Sin78a, Som71a, Som71b, Som76, Wei78, You72b, DS76a]. self [Haw71]. self-exciting [Haw71]. Semi [LSZ78]. Semi-Markov [LSZ78]. Sensitivity [PP78, Pre76]. Separate [And72, De 77, dBP77, dBP78, Wal67, Wal70a]. Sequence [DH75b, Gor73, HH71b, Smi75, Hin70b, HH70, Smi76]. Sequence-Comparison [Gor73]. Sequences [Kin72]. Sequential [AH76, BD78, BM79, Bil77, Bin75, Bin76, Efr71, ESF73, FL71, Fre70, Fre72, Fre73b, Fre73c, Hay79, Joa72, Joa75, Jon74b, JW79, JW81, KP78, LS75b, Lou75, Man70b, Mee78, MI77, OW78, Poc77, RA79, Rea73, Rob78c, Sie78, SWH75, TY74, UL76, Wei78, Whi78, WJ79, Woo76, Lou76]. Sequentially [And71b, AH73]. Serial [CH75, DW71, KP76, Mor70a, Ram72]. Series [AB79, Blo72, Blo73, BT77, Bri77, Bru78, Cam78, DN79, Dom72, Eng75, God79, HW78, Hos79, Kem75, KC76, LB78, Mac74b, Mac77b, MG72, MR70, Nea70, Nic76, Nic77, Pie72, Ree72, Tia72, TW76, TH78, Wal67, Wal70a, Wal71, Wis68, DT72, Que49, Wis90]. Set [Fuj79, OW78, Pus70, SY74, ST71, Yan78, Yan80]. Sethuraman [Sch78b]. Sets [Ket71]. Setting [Saw70]. Several [AF75b, Cox76a, Cox77, Ket71, LC76, Mac77b, Ofo73, Ofo75, DC77, Zel71]. Shakespeare [ET76, Wil75a]. Shannon [GJ79a]. Shape [HM75b, MG78, PP75]. Shapiro [Sar75]. Shapiro-Francia [Sar75]. Share [Mac77b]. Sharp [BF72]. Sharpening [Sha76b]. Sharper [Sat77]. Short [Bro75]. Short-Cut [Bro75]. Shortest [Wal70b]. Shrinkage [AAK72]. Sided [BF72, Shi78, Sli70, Van73a, Wyn75]. Sign [Gea70]. Signals [CH79]. Significance [And71c, Cox75a, DTN77, Rog77, Rog78, Sie77]. Significant [Mes73]. Silvey [Gut71]. Simple [AH73, Bas72, Bes77, FGS74, Hin79, Kry72, Kry74, McN72, PP75, PR77, Tik70, HS72a, Pre71b]. Simplified [KMN75, RS70]. Simulation [DT71, Joh70, LS76a, Mor75b]. Simultaneous [Dal78, LMH67, LMH71, McK77, MDS74, Odé73, RS78a]. Single [Jos72, Upt73]. Single-Sample [Upt73]. Singular [Gil79, Sil78b, Sty70]. Singularities [Mul70, RS78c]. Situation [GK70]. Situations [Mol77]. Six [FK72]. Size [Ano81, Bow72, BO78, Cha74b, D’A71, DS68, Fre72, Hol73, Hol71b, KHB70a, KHB70b, Mul77, NS72, Nic70, PS73, Sin78a, Wat71, Yan78, Yan80, DS76a]. Sizes [Bis72a, Bis72b, PT71, ZP77]. Skewness [Mar70a]. Slippage [Jos72, You73b]. slope [Atk70]. Small [BM78a, DT71, D’A72, Geo71, LSG79, RR71, VB78, BK74, Boo75, LSG83]. small-sample [BK74]. Smirnov [Dur75, MD71, MM76, Woo78, WA78]. Smoothing [BM72]. Sobel [Fus73]. Sobel-Weiss [Fus73]. Solution [Cob78, Gue72]. Solutions [BS73c]. Some [AB79, Ait69, Ait71, Bel77, Bel79, Bes72, BD77, Blo76, CSW76, Che71, Chm76, Cla74, CS75, CS76, Cro74, DN79, De 77, Far79, Feu79, FR72, Goo73, Han75, HMM77, HC72, Hol70, HM72, Joh71, JK71a, Jos72, Kal74, Kay79, Kor72, Krz79, Lee72a, LDG75, Leo76, LS75a, LO78, Mar76a, Mar71, McL75, PT70, dBP77, Pre75, RT73, RK75, Rob75, Rob77, SHA78a, Sha77, She71b, Tho76, TH78, Tik70, Tit75, Tro79, Wal67, Wal70a, Wil77a, Wil77b, You70a, You72a, You73a, You73b, Atk74, Haw71, Hut71, Saw78]. Space [BS73c]. Spacings [Cre76, Dur75]. Spatial [Ali79, BM75, CS73, CS79a, Cor77, CL76, Dig77, HH79, Pal72, RS78b, Tjø78]. Spatial-Temporal [Ali79]. Spearman [Chm76, Woo70]. Spearman-Type [Chm76]. Special [Hin78d, Hol78, MPG76, PCdCM74, Saw70]. Species [ET76, Eng74, Eng75]. Specific [Seb70, Seb71]. Specification [DN79]. Specified [Gol74, Sri70, Wat74]. Specifying [NP73]. Spectra [Haw71]. Spectral [Adl77, Lin74, MR70, Nea70, PDT77, Sha77]. Spectrum [Blo73]. Sphere [BM78a, Mor79, Saw78]. Spheres [HIR72]. Spherical [Kin72, MP78]. Sphericity [Joh72, NS79]. Split [Rob70]. Splitting [Cox75a]. Springer [Joh70]. Sprott [Cox72a, Sto72]. Spurious [DGS74]. Square [Goo79b, Har64, Jan75, McC78, Sha70a, Har65, Har79]. Squared [Bai79, Bri74, Dah71, DG72, Goo69, Goo70, Han75, Hea77, Kha78, KM73, Mol77, Sie79, Sin79, Sla73, Sma78]. Squares [AL71, BW75, Bro78, Cor72, DW71, Gea70, Kno75b, Mil76, Pie71, Pie72, Wat72, Pla72]. Srinivasan [Moo73]. Stabilizing [Woo70]. Stable [BB78, PHL75, Wil77b]. Stage [BTG77, DGS74, Fea77, HSW72, Som71b, You72b, You74]. Stand [Cox76c]. Standard [Ano48a, Cop72, FH70b, Gea36, Gue72, Hea78, Pre70, Pre71a, Pre79b, Tik71a, Mor75b]. Standardized [SBL79, You71]. State [HM72, Ked76, MRR^{+}53]. States [Han72, Sha76a, Sha78b]. Stationarity [And75d]. Stationary [Ali79, Bes72, Bha76, Bri74, Cam78, Ked78, LB79, Nea70, PW75, PDT77, Wal71]. Statistic [Bow73, Bri74, Bro74d, BY77, CR72, Cor77, Cre79, Dah71, Gor73, HH79, HS72b, Joh72, Koe79, KP78, LS79, Lee71, You70a, You72a, BS73a, DTN77, KG76, Lep73, Pet76b, SG74]. Statistical [AB76, Ait77, Ait79, AL79, CS79a, Coo79, Cox72a, Cox72b, Dav76b, GSW75, HS79, Joh74a, Mac70, Pee78, Sha76a, Sha78b, She71b, SD72, Sto72, Goo79a, Sär71]. Statistically [Mes73]. Statistics [Afo72, Ano70g, Ano70h, Ano70i, Ano71h, Ano71g, Ano71i, Ano72g, Ano72h, Ano72i, Ano73g, Ano73h, Ano73i, Ano74g, Ano74h, Ano74i, AG74, Cam78, Cla74, CS71, Cra72, Dow72, DGK71, Fin79a, GP70, GNP73, HS79, HP72, Jos73, KMN75, Ked76, KLC76, Lat77, LS76b, LS77b, MD71, Mar77, MN74, PP75, Pet76a, Pre74d, Rab70, She70, She71b, Sin72, Ste75a, Sti75, Sub69, Sub70, Tho76, Upt74, You70b, You71, You73a, You73b, Blo76, DP72, Goo79a, HS72a, HR77, Ken71, Ken72, Lep71, LS77a, Mac77a, PS76, Pet77, Pet78, Pla72, Sär71, Sea77, Sti73, Sti74, WA78]. Stein [EM72]. step [MH78]. Stochastic [Bar75a, DG76, Kry72, Kry74, McN72, PW75]. Stopping [Ber74]. Straight [BW71, Col72]. Stratification [LGGB79, Rao73b, Rao73a]. Stratified [New71]. Stratum [Jon74a]. Strauss [KR76]. Structural [Cox76a, Cox77, Dol76, Rob74, DC77]. Structure [BT74, Bes72, Goo74, Kie75, Lee79, Qui70, Saw70, Wor75, SPB78]. Structured [Far78]. Structures [Jör70, KLC76]. Studden [HW72]. Student [Hin78a, BDJS78, HH71a, Hin77b, Mic75, Pre74c, SS70]. Studentized [Bro74d, DLB72, Mar77]. Studies [Cla78, Cra72, DN79, DP72, Goo79a, HS72a, HP74, Joh70, Ken71, Ken72, Mac77a, Mes73, Pla72, PB78, PP79, Rab70, Sär71, Sea77, She70, She71b, Sti73, Sti74, Sti75, Wil75a]. Study [LDG75, Tik70, DL73]. Stylized [Pre76]. Subclass [Mar79b]. subdivided [Mar70b]. Subject [Bre70, Fed74, OL76, You72b]. Subsample [FH70a]. Subset [Box71, DS68, DS76a]. Subset-Size [DS68, DS76a]. Successive [HPF70, HPF71, Sha70a]. Sufficiency [Kal75b, Ked78, McL75, Pee78, Spr75, Sti73]. Sufficient [Afo72, Ked76, Kha78]. Sum [Dar71, Hin71, KM79, LW78, MT67, She71a, SS77, WL77, MT72]. Summands [AL76]. Sums [And75b]. Supercritical [Aba74]. Superimposed [HW73]. Supplementary [GK70]. Surface [Mor79, Atk70]. Surfaces [KM73, Pre76]. Survey [Yan78, Yan80]. Surveys [Ram79, Rao73b, SS75, Rao73a]. Survival [Cox79, DF79, HMM77, HP74, Hyd77, Kal74, KM77, KM79, Kay79, KP78, LW78, LGGB79, Oak77, Pol75, Seb70, Seb71, TW77, WL77]. Survivals [Pre73]. Switching [Atk79]. Symmetric [Hin76b, Sch75, Spi77, Sch76]. Symmetrical [Buk72]. Symmetrically [Hea78]. Symmetry [DFA77, Hin75, Hol71a, Juo70, Kin72, HR77, Hin76a, SG74]. Synchrony [San71]. Synthesis [Dol76, DC77]. System [Ait77, Bri77, HM72, MG72, MV76, Thi74]. Systematic [BR75, BR81, BM78b, Hei78]. Systems [Gai74, GW75].

Table [Alt71b, AP70, Boo75, HWB^{+}77, Har64, Tar75, VB78, HH71a, Har65, Har79, Lep73, Pla77, ZP77]. Tables [Bes73, Bow72, Cla78, Dar74, FK72, Gar70, Gar73, Goo73, Goo79b, Gue72, GD74, Ham70, JP69, JP70, JK71a, JK71b, Joh74b, Lai78, Lee72b, McC78, Mee78, MT77, Mor78, Mul77, Nel77, SD76, Van70, Van73a, Whi78, Wil76, Boo75, Cop73, Dav79, Fie72, Gar72b, Wil78, Zel71]. Tag [Pol75]. Tag-Recapture [Pol75]. Tagging [Fre73b]. Tail [Mic75]. Tails [LS77b]. Tau [Bes73, Sim77, Woo70]. Tears [FGS74]. Technique [BM72, Som71b]. Techniques [Cha73a, Cha73b, GL70]. Teller [Har10]. Temporal [Ali79, Tia72, TW76]. Tendency [Cla78]. Terminals [JK71b]. Terms [Pus70]. Terry [Atk72b]. Tes [Sla73]. Test [AL71, BS75, Bre70, Bro75, CR72, CS71, Cro74, D’A71, D’A72, DR74, Dah71, DG72, DB69, DB70, DN79, DA70b, Gra79, Gre70, Gre79, HM75b, Hol71a, Lee71, LS75a, LSG79, Mac77b, Man70a, MG73, Mol77, Moo73, MB73, Mor73b, PP75, Pee71, Rog77, Rog78, Sam70, SA76, SFC72, Shi78, Sim77, Sis76, Sli70, Spi77, Ste75b, TP74, VB78, Atk72b, BK74, HS72a, Lew72, Ode72, Sar75, Woo78, Yat72, ZP77, LSG83]. Testing [BD78, BM75, Bho78, Bho79, Cha74b, CS78, Dav77, DW71, God79, HWB^{+}77, HS78a, HS78c, Hwa76, Hyd77, Joh72, Kno75a, Kno77, KP78, LJ76, LS76b, LS77b, Mac74a, MPG76, MK79, NP73, Nai75, Nel77, Pet76a, PK79, Sie79, SE75, Sri70, Tar79b, Tar90, Vit74, Wis68, Wol76, Yan78, Yan80, LS77a, SG74, Wis90]. Tests [Ait69, Ait71, AS72, And71b, And71c, AH73, BD76, BS76a, Bow73, Bro77b, Buh77, Cha74a, Cho77, Cre78, DP73, DP74, DR73, DLB72, De 77, DFA77, Dow76, DTL76, Dur75, Dye74, FS71, Fuj79, GW78, Gar73, Gas75, GO77, GP75, HM75a, Hay79, HR77, Hin71, HP75, HP76, HP80, Hol72b, Hol72a, JMM75, Joa72, Joh71, JW79, Jos72, Kno77, Kor72, KLC76, Lay74, Lee71, LDG75, LGGB79, Lou75, Mac74b, Man70b, Mar76a, Mar71, MM76, Mee78, Mee79, Mor70b, Mor73a, Mor78, MDS74, MH71, PDB77, dBP77, dBP78, Pet72, Pet77, Pre74a, Pre78, Pre76, RT73, Rea73, RS78b, SHA78a, Sha75, Shi77, Sie77, Sie78, Ste79a, Swe78, Tar75, TW77, Upt73, Wal67, Wal70a, Wil76, Woo76, Woo79, You70a, DL73, Gar72b, Joa75, JW81, Lou76, Pre83, Pre71b]. tests [Que49, Saw78, Wil78]. Tetrachoric [Ham70]. Their [Has70, Koc79, Spr73]. Theorem [Mat70, Som71a, Som76, Whi73a]. Theory [Bro72, Cox72b, Dan72, KP76, Pal71, Rob74, Rob75, Rob77, Roy70, Roy76, ST73, Spr73, Sär71, HW72]. There [DGS74]. Thinned [BH74]. Three [AH76, GW75, Hay79, Hil78, MI77, MV76, PPK73]. Three-Dimensional [PPK73]. Three-Phase [MI77]. Threshold [EB78]. Time [AB79, Blo72, Blo73, BT77, Bri77, Cam78, DN79, Far77, God79, HW78, Hin72, Hos79, KM77, LB78, Mac74b, Mac77b, MG72, MR70, Nea70, Nic76, Nic77, O’N79, PDT77, Pie72, PB78, RT73, Seb70, Tia72, TW76, TH78, TN79, Wal67, Wal70a, Wal71, Que49]. Time-Censored [Far77]. Time-Dependent [KM77]. Time-Ordered [Hin72]. time-series [Que49]. Time-Specific [Seb70]. Tolerance [Mil73, Sha70c]. Total [Sim77]. totals [Pla77]. Tournament [Dan69, Dan70]. Transect [Ram79]. Transfer [Ait77, RT73]. Transform [Mar77]. Transformation [D’A70a, GW73, Woo70]. Transformations [And71a, BM78b, Hin75, JK72, Krz79, Hin76a, Pre74c]. Transforms [Dut73]. Transient [CH79]. Transition [BW71, Han72]. Trapping [GLGH79]. Treatment [BEH76, Edw73, FL71]. Treatments [Fre73a, Gla78, Sli70]. Trend [BR75, BY77, Tar75, BR81]. Trends [Rog77, Rog78]. Trial [Gla78, TN79, Yat72]. Trials [BM79, FL71, Gai74, GW75, MI77, Poc77, SWH75, UL76, WJ79]. Triangular [Sin78b]. Trimmed [Yue74, YD73]. Trinomials [HS78a]. Truncated [Bil77]. Truncation [Hyd77, Lwi75]. Tukey [BP76, JK73]. Turing [Goo79a]. Turning [Cox76b, HS72a]. Twentieth [Har10]. Twin [HP74]. Two [Adl77, AH76, And76b, BW71, BCK78, Bel77, BTG77, Bel79, Bes72, Bro74b, Buh77, Cha76, Dav75, Dav76a, DGS74, DS76b, DG76, Fea77, Gai74, Gla78, Hay79, HPF70, HPF71, Her76, Hin69, Hin70a, HSW72, HS78c, HM72, Jon75, Jos73, Kap70, Ked76, Koc79, KP78, Lai78, Lat77, LDG75, Lew75, LS75a, Lou75, MS75, MT67, New71, Ofo73, Ofo75, Ran74, SM72, SY74, SD76, SS77, Sin78a, Smi73, SW70, Som71b, Swi76, TY74, TW76, Tik71a, Ton71, Van73a, WAK77, Whi73b, You70a, You72a, You73a, Yue74, dJS73, AF75a, Fus73, Gre79, Lou76, MT72, Pet76b, TA71, YD73]. Two-Armed [Jon75, WAK77]. Two-Dimensional [Bes72, DG76, Lew75]. Two-Phase [New71, dJS73]. Two-Population [Hay79]. Two-Sample [Bro74b, Lat77, LDG75, Ofo73, Ofo75, Swi76, You72a, You73a, Yue74, Pet76b, YD73]. Two-Sided [Van73a]. Two-Stage [BTG77, DGS74, Fea77, HSW72, Som71b]. Two-State [HM72, Ked76]. Two-Way [Her76, Lai78, MT67, Ran74, SM72, SD76, Smi73, MT72]. Type [AR81, Chm76, MD71, MM76, Pet72, PDT78, SH76, Vuc77, Dal78, Pre71b, SG74, AR72, PR77, Rod77].

Ultrastructural [Bro78, Dol76, Pat78, DC77]. Un-Bayesian [SD72]. Unbiased [Cha77, Har69, Har70, MN74, MS76, Sat77, Wal70b]. Uncertainty [OL76]. Unclassified [GM78]. Unconditional [God76a]. Underlying [HM75b]. Unequal [Bre70, Fre73a, PT71, Pol75, Yue74]. Unidentifiable [Goo74]. Uniformity [Sie79]. unimodal [DT72]. Unimodality [Cop75]. Unique [PW75]. Uniqueness [Wed76]. Units [Sin78a]. Univ [Boo75]. Univariate [Bru78, DB69, DB70, Dav76b, Dal78]. Unknown [DA70b, Gra79, MM76, Ofo73, Ofo75, SFC72, Spj72]. Unlimited [LMH67, LMH71]. unmatched [Cop73]. Unreplicated [Pat78]. Unsen [ET76]. Upper [AR72, BT73, CC77, Lee72b, Loy79, AR81]. Use [CK68, CK70, CS79a, Cop72, Hem77, LS76b, LS77b, Pre71b, LS77a]. Used [Cra72, HP72, Joh72, SW73]. Using [Ber75, Bro75, Buk72, Cor72, Cox76c, Dig75, Gar72a, God78, Goo74, Har74, Has70, Hin77b, Hin78a, KP78, LS77c, Mur79, Nea70, Pes73, Pre70, Pre76, RW75, You77, dJS73, Pet77].

V [BN78, Cox72a, HW72, Jam78, Rob78a, Spr78, Sto72]. Vaccination [Bec72]. Validation [Sto74, Sto77]. Value [AL76, AR72, CD78, Gai74, GW75, Gil79, Kap70, Law74, Sam70, Ste77, AR81, HH71a]. Valued [Dar71]. Values [BY77, MG73, Sis76, Sla73]. variability [Mar70b]. Variable [Alt76, BMW78, BMW79, Cha76, Far77, LW78, McK77, SW73, WL77]. Variable-Sum [LW78, WL77]. Variables [Atk79, BP76, Coo79, Dar71, DH75a, DH77, Fed74, Geb70, Gil79, Goo73, GNP73, Hin69, Hin70a, HH71b, JK72, Ket71, Krz79, Nai75, Pre73, She71a, Smi75, Sty70, TW76, Tjø78, Wil77b, Yan78, Yan80, Hin70b, HH70, Smi76]. Variance [Afo72, AP70, Bes74, Blo78, Cam78, Chm76, DA70b, Daw77, Har69, Har70, Har74, Has74, Hei78, JW71, KHB70a, Kor72, MPG76, MS76, MDS74, MH71, Nel77, Olk72, Ran74, RW75, SHA78a, Sat77, See75, Wil70a, Woo70, Wor75]. Variances [Ahs71, Bho79, Fuj79, Ofo73, Ofo75, Pre74d, RK75, SA76, Spj72, Thi74, Yue74]. Variate [You72b]. Variates [Bho78, Bho79, BDJS78, Mor73b, Wil70b, HH71a]. Variation [Hin77a, Hin78b, Kie75, Hin78c]. Vary [Ano81, BO78]. Varying [Bis72a, Bis72b, Luc77, RK75]. Vector [Bai79, EM72, MB73, Nic76, Nic77, Ste75b, Ste79b]. Version [GK71, Fus73]. Versus [EH78b, Sli70]. via [AC77, Mar77]. Views [BCK78]. Volume [Ano70j, Ano71j, Ano72j, Ano73j, Ano74j, Ano75g, Ano76g, Ano77g, Ano78g, Ano79g].

W [Cox70a, HW72]. Waite [DR73]. Wald [Man70a]. Wallis [Bre70]. War [Goo79a]. Watson [CH72, HM75a, Qui70, Sti70]. Way [Hai71, Har69, Har70, Her76, KHB70a, Lai78, MT67, Ran74, SM72, SD76, Smi73, Spj72, MT72]. Weather [Feu79]. Weibull [Ber75, Has71, Law74]. Weights [RW75]. Weiss [Fus73]. Whether [Mac77b]. which [Fea77]. Whittaker [Bri74]. Whose [MV76]. Width [Sil78a]. Wiener [Atk74]. Wiener-Hopf [Atk74]. Wilcoxon [JW81, Gre79, JW79, Lat77, Lep71, Lep73]. Wilks [PG70, Dav79, HM76, Lee72a, PG69]. Window [Sil78a]. Winner [NS72, PS73, SW70, Fus73]. Wishart [Sis72]. Wishart-Matrix [Sis72]. witch [Sti74]. Within [Mac74a]. Without [FGS74, Ber74, Jen79]. Wold [God76b]. Wolfowitz [Joh70]. Word [Wil75a]. Word-Length [Wil75a]. Words [ET76]. Work [Sti75, Goo79a, Ken71, Ken72]. Works [Wil75a]. World [Goo79a].

XII [Rod77]. XXIII [She70]. XXIV [Rab70]. XXIX [Pla72]. XXV [She71b]. XXVI [Ken72, Ken71]. XXVII [Sär71]. XXVIII [Cra72]. XXX [DP72]. XXXI [HS72a]. XXXII [Sti73]. XXXIII [Sti74]. XXXIV [Sti75]. XXXV [Sea77]. XXXVI [Mac77a]. XXXVII [Goo79a].

Yule [FR79].

Zacks [Cox72b]. Zero [Lac75, Sie79]. Zero-Mean [Lac75].

References

Aitchison:1976:MBD

[AA76] John Aitchison and C. G. G. Aitken. Multivariate binary discrimination by the kernel method. Biometrika, 63(3):413–420, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335719.

Andrews:1972:SPM

[AAK72] R. L. Andrews, J. C. Arnold, and R. G. Krutchkoff. Shrinkage of the posterior mean in the normal case. Biometrika, 59(3):693–695, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334825.

Aitchison:1970:PQR

[AB70] J. Aitchison and J. A. Bennett. Polychotomous quantal response by maximum indicant. Biometrika, 57(2):253–262, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334834.

Aitchison:1976:SDW

[AB76] John Aitchison and C. B. Begg. Statistical diagnosis when basic cases are not classified with certainty. Biometrika, 63(1):1–12, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335078.

Abraham:1979:BAS

[AB79] Bovas Abraham and George E. P. Box. Bayesian analysis of some outlier problems in time series. Biometrika, 66(2):229–236, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335653.

Abakuks:1974:NSC

[Aba74] Andris Abakuks. A note on supercritical carrier-borne epidemics. Biometrika, 61(2):271–275, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334354.

Atkinson:1977:RRD

[AC77] A. C. Atkinson and D. R. Cox. Robust regression via discriminant analysis. Biometrika, 64(1):15–19, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335764.

Adler:1977:SME

[Adl77] Robert J. Adler. A spectral moment estimation problem in two dimensions. Biometrika, 64(2):367–373, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335705.

Atkinson:1975:DED

[AF75a] A. C. Atkinson and V. V. Fedorov. The design of experiments for discriminating between two rival models. Biometrika, 62(1):57–70, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334487.

Atkinson:1975:ODE

[AF75b] A. C. Atkinson and V. V. Fedorov. Optimal design: Experiments for discriminating between several models. Biometrika, 62 (2):289–303, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335364.

Afonja:1972:MSS

[Afo72] Biyi Afonja. Minimal sufficient statistics for variance components for a general class of designs. Biometrika, 59(2):295–302, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334573.

Arnold:1974:BDB

[AG74] Barry C. Arnold and Richard A. Groeneveld. Bounds for deviations between sample and population statistics. Biometrika, 61 (2):387–389, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334372.

Andrews:1973:SMC

[AH73] D. F. Andrews and Agnes M. Herzberg. A simple method for constructing exact tests for sequentially designed experiments. Biometrika, 60(3):489–497, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334997.

Akeda:1976:RSP

[AH76] Yoshiaki Akeda and Motoo Hori. On random sequential packing in two and three dimensions. Biometrika, 63(2):361–366, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335631.

Ahsanullah:1971:EMB

[Ahs71] M. Ahsanullah. On the estimation of means in a bivariate normal distribution with equal marginal variances. Biometrika, 58(1):230– 233, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334338.

Aitkin:1969:STC

[Ait69] Murray A. Aitkin. Some tests for correlation matrices. Biometrika, 56(2):443–446, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334438. See correction [Ait71].

Aitkin:1971:CST

[Ait71] Murray A. Aitkin. Correction: “Some Tests for Correlation Matrices”. Biometrika, 58(1):245, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334343. See [Ait69].

Aitchison:1975:GPF

[Ait75] James Aitchison. Goodness of prediction fit. Biometrika, 62(3): 547–554, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335509.

Aitchison:1977:CPS

[Ait77] John Aitchison. A calibration problem in statistical diagnosis: The system transfer problem. Biometrika, 64(3):461–472, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2345322.

Aitchison:1979:CPS

[Ait79] J. Aitchison. A calibration problem in statistical diagnosis: The clinical amalgamation problem. Biometrika, 66(2):357–366, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335671.

Akaike:1973:MLI

[Aka73] Hirotugu Akaike. Maximum likelihood identification of Gaussian autoregressive moving average models. Biometrika, 60(2):255–265, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334537.

Akaike:1978:NLB

[Aka78] Hirotugu Akaike. A new look at the Bayes procedure. Biometrika, 65(1):53–59, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335276.

Akaike:1979:BEM

[Aka79] Hirotugu Akaike. A Bayesian extension of the minimum AIC procedure of autoregressive model fitting. Biometrika, 66(2):237–242, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335654.

Antle:1970:CIP

[AKH70] Charles Antle, Lawrence Klimko, and William Harkness. Confidence intervals for the parameters of the logistic distribution. Biometrika, 57(2):397–402, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334848.

Abrahamse:1971:NTA

[AL71] A. P. J. Abrahamse and A. S. Louter. On a new test for autocorrelation in least squares regression. Biometrika, 58(1):53–60, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334316.

Anis:1976:EVA

[AL76] A. A. Anis and E. H. Lloyd. The expected value of the adjusted rescaled Hurst range of independent normal summands. Biometrika, 63(1):111–116, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335090.

Aitchison:1979:SDI

[AL79] J. Aitchison and I. J. Lauder. Statistical diagnosis from imprecise data. Biometrika, 66(3):475–483, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335166.

Ali:1977:AAM

[Ali77] Mukhtar M. Ali. Analysis of autoregressive-moving average models: Estimation and prediction. Biometrika, 64(3):535–545, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345330. See

correction [Ali78].

Ali:1978:ACA

[Ali78] M. M. Ali. Amendments and corrections: “Analysis of autoregressive-moving average models: estimation and prediction” (Biometrika 64 (1977), 535–545). Biometrika, 65(3):677, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335932. See

[Ali77].

Ali:1979:ASS

[Ali79] Mukhtar M. Ali. Analysis of stationary spatial-temporal processes: Estimation and prediction. Biometrika, 66(3):513–518, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335170.

Altham:1971:AMP

[Alt71a] Patricia M. E. Altham. The analysis of matched proportions. Biometrika, 58(3):561–576, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334391.

Altham:1971:EBA

[Alt71b] Patricia M. E. Altham. Exact Bayesian analysis of an intraclass 2 × 2 table. Biometrika, 58(3):679–680, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334407.

Altham:1976:DVA

[Alt76] Patricia M. E. Altham. Discrete variable analysis for individuals grouped into families. Biometrika, 63(2):263–269, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335618.

Ansley:1979:FSD

[AN79] Craig F. Ansley and Paul Newbold. On the finite sample distribution of residual autocorrelations in autoregressive-moving average models. Biometrika, 66(3):547–553, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335174.

Andrews:1971:NSD

[And71a] D. F. Andrews. A note on the selection of data transformations. Biometrika, 58(2):249–254, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334514.

Andrews:1971:SDE

[And71b] D. F. Andrews. Sequentially designed experiments for screening out bad models with F tests. Biometrika, 58(3):427–432, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334379.

Andrews:1971:STB

[And71c] D. F. Andrews. Significance tests based on residuals. Biometrika, 58(1):139–148, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334324.

Anderson:1972:SSL

[And72] J. A. Anderson. Separate sample logistic discrimination. Biometrika, 59(1):19–35, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334611.

Anderson:1975:QLD

[And75a] J. A. Anderson. Quadratic logistic discrimination. Biometrika, 62(1):149–154, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334497.

Anderson:1975:BSA

[And75b] O. D. Anderson. Bounding sums for the autocorrelations of moving average processes. Biometrika, 62(3):706–707, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335539.

Anderson:1975:NDA

[And75c] O. D. Anderson. A note on differencing autoregressive moving average (p, q) processes. Biometrika, 62(2):521–523, August 1975.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335395.

Anderson:1975:RNS

[And75d] O. D. Anderson. The recursive nature of the stationarity and invertibility restraints on the parameters of mixed autoregressivemoving average processes. Biometrika, 62(3):704–706, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335538.

Anderson:1976:IAM

[And76a] O. D. Anderson. On the inverse of the autocovariance matrix for a general moving average process. Biometrika, 63(2):391–394, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335634.

Anderson:1976:TCA

[And76b] O. D. Anderson. On two convex autocorrelation regions for moving average processes. Biometrika, 63(3):681–683, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335755.

Anderson:1979:MLC

[And79] J. A. Anderson. Multivariate logistic compounds. Biometrika, 66(1):17–26, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335237.

Anonymous:1948:CMR

[Ano48a] Anonymous. Corrigenda: “Moments of the Ratio of the Mean Deviation to the Standard Deviation for Normal Samples”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332365. See [Gea36].

Anonymous:1948:CPM

[Ano48b] Anonymous. Corrigenda: “The Probability of the Mean Deviation”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332366. See [Gea36, Pea45].

Anonymous:1970:BMa

[Ano70a] Anonymous. Back matter. Biometrika, 57(1):??, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334968.

Anonymous:1970:BMb

[Ano70b] Anonymous. Back matter. Biometrika, 57(2):??, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334864.

Anonymous:1970:BMc

[Ano70c] Anonymous. Back matter. Biometrika, 57(3):i–vi, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334798.

Anonymous:1970:FMa

[Ano70d] Anonymous. Front matter. Biometrika, 57(1):??, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334931.

Anonymous:1970:FMb

[Ano70e] Anonymous. Front matter. Biometrika, 57(2):??, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334831.

Anonymous:1970:FMc

[Ano70f] Anonymous. Front matter. Biometrika, 57(3):??, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334764.

Anonymous:1970:NPSa

[Ano70g] Anonymous. New publications in statistics and probability. Biometrika, 57(1):227, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334967.

Anonymous:1970:NPSb

[Ano70h] Anonymous. New publications in statistics and probability. Biometrika, 57(2):469, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334863.

Anonymous:1970:NPSc

[Ano70i] Anonymous. New publications in statistics and probability. Biometrika, 57(3):684, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334797.

Anonymous:1970:VI

[Ano70j] Anonymous. Volume information. Biometrika, 57(3):i–ix, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334763.

Anonymous:1971:BMa

[Ano71a] Anonymous. Back matter. Biometrika, 58(1):??, April 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334346.

Anonymous:1971:BMb

[Ano71b] Anonymous. Back matter. Biometrika, 58(2):??, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334531.

Anonymous:1971:BMc

[Ano71c] Anonymous. Back matter. Biometrika, 58(3):i–vi, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334415.

Anonymous:1971:FMa

[Ano71d] Anonymous. Front matter. Biometrika, 58(1):??, April 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334312.

Anonymous:1971:FMb

[Ano71e] Anonymous. Front matter. Biometrika, 58(2):??, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334513.

Anonymous:1971:FMc

[Ano71f] Anonymous. Front matter. Biometrika, 58(3):??, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334376.

Anonymous:1971:NPSb

[Ano71g] Anonymous. New publication in statistics and probability. Biometrika, 58(2):401–402, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334530.

Anonymous:1971:NPSa

[Ano71h] Anonymous. New publications in statistics and probability. Biometrika, 58(1):247–248, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334345.

Anonymous:1971:NPSc

[Ano71i] Anonymous. New publications in statistics and probability. Biometrika, 58(3):688, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334414.

Anonymous:1971:VI

[Ano71j] Anonymous. Volume information. Biometrika, 58(3):i–viii, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334375.

Anonymous:1972:BMa

[Ano72a] Anonymous. Back matter. Biometrika, 59(1):??, April 1972. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334649.

Anonymous:1972:BMb

[Ano72b] Anonymous. Back matter. Biometrika, 59(2):??, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334607.

Anonymous:1972:BMc

[Ano72c] Anonymous. Back matter. Biometrika, 59(3):i–vi, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334830.

Anonymous:1972:FMa

[Ano72d] Anonymous. Front matter. Biometrika, 59(1):??, April 1972. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334608.

Anonymous:1972:FMb

[Ano72e] Anonymous. Front matter. Biometrika, 59(2):??, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334568.

Anonymous:1972:FMc

[Ano72f] Anonymous. Front matter. Biometrika, 59(3):??, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334800.

Anonymous:1972:NPSa

[Ano72g] Anonymous. New publications in statistics and probability. Biometrika, 59(1):238, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334648.

Anonymous:1972:NPSb

[Ano72h] Anonymous. New publications in statistics and probability. Biometrika, 59(2):499, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334606.

Anonymous:1972:NPSc

[Ano72i] Anonymous. New publications in statistics and probability. Biometrika, 59(3):700, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334829.

Anonymous:1972:VI

[Ano72j] Anonymous. Volume information. Biometrika, 59(3):i–ix, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334799.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. Biometrika, 60(1):??, April 1973. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334930.

Anonymous:1973:BMb

[Ano73b] Anonymous. Back matter. Biometrika, 60(2):??, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334567.

Anonymous:1973:BMc

[Ano73c] Anonymous. Back matter. Biometrika, 60(3):??, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335024.

Anonymous:1973:FMa

[Ano73d] Anonymous. Front matter. Biometrika, 60(1):??, April 1973. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334899.

Anonymous:1973:FMb

[Ano73e] Anonymous. Front matter. Biometrika, 60(2):??, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334532.

Anonymous:1973:FMc

[Ano73f] Anonymous. Front matter. Biometrika, 60(3):??, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334991.

Anonymous:1973:NPSa

[Ano73g] Anonymous. New publications in statistics and probability. Biometrika, 60(1):215, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334929.

Anonymous:1973:NPSb

[Ano73h] Anonymous. New publications in statistics and probability. Biometrika, 60(2):437, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334566.

Anonymous:1973:NPSc

[Ano73i] Anonymous. New publications in statistics and probability. Biometrika, 60(3):671, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335023.

Anonymous:1973:VI

[Ano73j] Anonymous. Volume information. Biometrika, 60(3):i–viii, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334990.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Biometrika, 61(1):??, April 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334311.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Biometrika, 61(2):??, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334374.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. Biometrika, 61(3):??, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334762.

Anonymous:1974:FMa

[Ano74d] Anonymous. Front matter. Biometrika, 61(1):??, April 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334279.

Anonymous:1974:FMb

[Ano74e] Anonymous. Front matter. Biometrika, 61(2):??, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334347.

Anonymous:1974:FMc

[Ano74f] Anonymous. Front matter. Biometrika, 61(3):??, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334719.

Anonymous:1974:NPSa

[Ano74g] Anonymous. New publications in statistics and probability. Biometrika, 61(1):206, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334310.

Anonymous:1974:NPSb

[Ano74h] Anonymous. New publications in statistics and probability. Biometrika, 61(2):391, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334373.

Anonymous:1974:NPSc

[Ano74i] Anonymous. New publications in statistics and probability. Biometrika, 61(3):649, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334761.

Anonymous:1974:VI

[Ano74j] Anonymous. Volume information. Biometrika, 61(3):i–viii, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334718.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. Biometrika, 62(1):??, April 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334512.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. Biometrika, 62(2):??, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335399.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. Biometrika, 62(3):??, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335543.

Anonymous:1975:FMa

[Ano75d] Anonymous. Front matter. Biometrika, 62(1):??, April 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334480.

Anonymous:1975:FMb

[Ano75e] Anonymous. Front matter. Biometrika, 62(2):??, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335353.

Anonymous:1975:FMc

[Ano75f] Anonymous. Front matter. Biometrika, 62(3):??, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335507.

Anonymous:1975:VI

[Ano75g] Anonymous. Volume information. Biometrika, 62(3):i–ix, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335506.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Biometrika, 63(1):??, April 1976. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335116.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Biometrika, 63(2):??, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335645.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Biometrika, 63(3):??, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335760.

Anonymous:1976:FMa

[Ano76d] Anonymous. Front matter. Biometrika, 63(1):??, April 1976. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335077.

Anonymous:1976:FMb

[Ano76e] Anonymous. Front matter. Biometrika, 63(2):??, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335612.

Anonymous:1976:FMc

[Ano76f] Anonymous. Front matter. Biometrika, 63(3):??, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335718.

Anonymous:1976:VI

[Ano76g] Anonymous. Volume information. Biometrika, 63(3):i–ix, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335717.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. Biometrika, 64(1):??, April 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335798.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. Biometrika, 64(2):??, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335716.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. Biometrika, 64(3):??, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345359.

Anonymous:1977:FMa

[Ano77d] Anonymous. Front matter. Biometrika, 64(1):??, April 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335761.

Anonymous:1977:FMb

[Ano77e] Anonymous. Front matter. Biometrika, 64(2):??, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335682.

Anonymous:1977:FMc

[Ano77f] Anonymous. Front matter. Biometrika, 64(3):??, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345317.

Anonymous:1977:VI

[Ano77g] Anonymous. Volume information. Biometrika, 64(3):i–ix, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345316.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Biometrika, 65(1):??, April 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335308.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Biometrika, 65(2):??, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335233.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Biometrika, 65(3):??, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335936.

Anonymous:1978:FMa

[Ano78d] Anonymous. Front matter. Biometrika, 65(1):??, April 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335269.

Anonymous:1978:FMb

[Ano78e] Anonymous. Front matter. Biometrika, 65(2):??, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335201.

Anonymous:1978:FMc

[Ano78f] Anonymous. Front matter. Biometrika, 65(3):??, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335892.

Anonymous:1978:VI

[Ano78g] Anonymous. Volume information. Biometrika, 65(3):i–xi, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335891.

Anonymous:1979:BMa

[Ano79a] Anonymous. Back matter. Biometrika, 66(1):??, April 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335268.

Anonymous:1979:BMb

[Ano79b] Anonymous. Back matter. Biometrika, 66(2):??, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335681.

Anonymous:1979:BMc

[Ano79c] Anonymous. Back matter. Biometrika, 66(3):??, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335200.

Anonymous:1979:FMa

[Ano79d] Anonymous. Front matter. Biometrika, 66(1):??, April 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335234.

Anonymous:1979:FMb

[Ano79e] Anonymous. Front matter. Biometrika, 66(2):??, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335646.

Anonymous:1979:FMc

[Ano79f] Anonymous. Front matter. Biometrika, 66(3):??, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335157.

Anonymous:1979:VI

[Ano79g] Anonymous. Volume information. Biometrika, 66(3):i–ix, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335156.

Anonymous:1981:ACE

[Ano81] Anonymous. Amendments and corrections: “Estimation of the Size of a Closed Population When Capture Probabilities Vary Among Animals”. Biometrika, 68(1):345, April 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335844. See [BO78].

Ansley:1979:AEL

[Ans79] Craig F. Ansley. An algorithm for the exact likelihood of a mixed autoregressive-moving average process. Biometrika, 66(1):59–65, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335242.

Amos:1970:CAT

[AP70] D. E. Amos and E. S. Pearson. Corrections and additions to a table of the percentage points of the variance ratio F. Biometrika, 57(1):211–212, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334954.

Antle:1972:UCL

[AR72] Charles E. Antle and Fred Rademaker. An upper confidence limit on the maximum of m future observations from a Type 1 extreme value distribution. Biometrika, 59(2):475–477, August 1972. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334595. See correction

[AR81].

Antle:1981:ACU

[AR81] Charles E. Antle and Fred Rademaker. Amendments and corrections: “An upper confidence limit on the maximum of m future observations from a Type I extreme value distribution” [Biometrika 59 (1972), 475–477; MR 49 #4143]. Biometrika, 68(3):738, December 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335466. See

[AR72].

Armitage:1970:CAR

[Arm70] P. Armitage. The combination of assay results. Biometrika, 57 (3):665–666, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334786.

Anderson:1972:TRD

[AS72] T. W. Anderson and M. A. Stephens. Tests for randomness of directions against equatorial and bimodal alternatives. Biometrika, 59(3):613–621, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334812.

Atkinson:1970:DEE

[Atk70] A. C. Atkinson. The design of experiments to estimate the slope of a response surface. Biometrika, 57(2):319–328, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334839.

Atkinson:1972:PED

[Atk72a] A. C. Atkinson. Planning experiments to detect inadequate regression models. Biometrika, 59(2):275–293, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334572.

Atkinson:1972:TLL

[Atk72b] A. C. Atkinson. A test of the linear logistic and Bradley-Terry models. Biometrika, 59(1):37–42, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334612.

Atkinson:1973:MSO

[Atk73] A. C. Atkinson. Multifactor second order designs for cuboidal regions. Biometrika, 60(1):15–19, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334901.

Atkinson:1974:WHI

[Atk74] C. Atkinson. A Wiener-Hopf integral equation arising in some inference and queueing problems. Biometrika, 61(2):277–283, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334355.

Atkinson:1978:PPC

[Atk78] A. C. Atkinson. Posterior probabilities for choosing a regression model. Biometrika, 65(1):39–48, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335274.

Atkinson:1979:FSA

[Atk79] A. C. Atkinson. A family of switching algorithms for the computer generation of beta random variables. Biometrika, 66(1):141–145, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335253.

Bailey:1977:PCF

[Bai77] R. A. Bailey. Patterns of confounding in factorial designs. Biometrika, 64(3):597–603, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345338.

Baillie:1979:APM

[Bai79] Richard T. Baillie. Asymptotic prediction mean squared error for vector autoregressive models. Biometrika, 66(3):675–678, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335193.

Barker:1965:MCC

[Bar65] A. A. Barker. Monte Carlo calculations of the radial distribution functions for a proton-electron plasma. Australian Journal of Physics, 18(??):119–133, ???? 1965. CODEN AUJPAS. ISSN 0004-9506 (print), 1446-5582 (electronic).

Barbour:1975:DCS

[Bar75a] Andrew D. Barbour. The duration of the closed stochastic epidemic. Biometrika, 62(2):477–482, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335390.

Barnard:1975:CPJ

[Bar75b] G. A. Barnard. Comments on paper by J. D. Kalbfleisch. Biometrika, 62(2):260–261, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335357. See [Kal75b].

Bailey:1970:IEL

[BAS70] Norman T. J. Bailey and Cynthia Alff-Steinberger. Improvements in the estimation of the latent and infectious periods of a contagious disease. Biometrika, 57(1):141–153, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334944.

Basawa:1972:EAC

[Bas72] I. V. Basawa. Estimation of the autocorrelation coefficient in simple Markov chains. Biometrika, 59(1):85–89, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334618.

Basawa:1978:IGS

[BB78] I. V. Basawa and P. J. Brockwell. Inference for gamma and stable processes. Biometrika, 65(1):129–133, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335287.

Balazs:2000:ONC

[BBLS00] Nandor L. Balazs, John C. Browne, James D. Louck, and Daniel S. Strottman. Obituary: Nicholas Constantine Metropolis. Physics Today, 53(10):100, October 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.aip.org/

pt/vol-53/iss-10/p100.html.

Baksalary:1978:RTD

[BCK78] J. K. Baksalary, L. C. A. Corsten, and R. Kala. Reconciliation of two different views on estimation of growth curve parameters. Biometrika, 65(3):662–665, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335926.

Box:1975:RD

[BD75] George E. P. Box and Norman R. Draper. Robust designs. Biometrika, 62(2):347–352, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335371.

Bancroft:1976:PDL

[BD76] G. A. Bancroft and I. R. Dunsmore. Predictive distributions in life tests under competing causes of failure. Biometrika, 63(1):195–198, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335103.

Bhansali:1977:SPO

[BD77] R. J. Bhansali and D. Y. Downham. Some properties of the order of an autoregressive model selected by a generalization of Akaike’s EPF criterion. Biometrika, 64(3):547–551, December 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345331.

Bancroft:1978:PSL

[BD78] G. A. Bancroft and I. R. Dunsmore. Predictive sequential life testing. Biometrika, 65(3):609–613, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335913.

Bishop:1978:PPQ

[BDJS78] Thomas A. Bishop, Edward J. Dudewicz, June M. Juritz, and Michael A. Stephens. Percentage points of a quadratic form in Student t variates. Biometrika, 65(2):435–439, August 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335227.

Brook:1972:APD

[BE72] D. Brook and D. A. Evans. An approach to the probability distribution of Cusum run length. Biometrika, 59(3):539–549, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334805.

Becker:1970:MDM

[Bec70] N. G. Becker. Mixture designs for a model linear in the proportions. Biometrika, 57(2):329–338, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334840.

Becker:1972:VPR

[Bec72] Niels Becker. Vaccination programmes for rare infectious diseases. Biometrika, 59(2):443–453, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334589.

Becker:1974:PEM

[Bec74] Niels Becker. On parametric estimation for mortal branching processes. Biometrika, 61(3):393–399, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334720.

Becker:1979:EPH

[Bec79] Niels Becker. An estimation procedure for household disease data. Biometrika, 66(2):271–277, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335658.

Behboodian:1970:MND

[Beh70] Javad Behboodian. On a mixture of normal distributions. Biometrika, 57(1):215–217, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334956.

Bradley:1976:TCP

[BEH76] Ralph A. Bradley and Abdalla T. El-Helbawy. Treatment contrasts in paired comparisons: Basic procedures with application to factorials. Biometrika, 63(2):255–262, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335617.

Bellhouse:1977:SOD

[Bel77] D. R. Bellhouse. Some optimal designs for sampling in two dimensions. Biometrika, 64(3):605–611, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345339. See correction [Bel79].

Bellhouse:1979:ACS

[Bel79] D. R. Bellhouse. Amendments and corrections: “Some Optimal Designs for Sampling in Two Dimensions”. Biometrika, 66(2):402, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335680.

See [Bel77].

Berry:1974:COS

[Ber74] Donald A. Berry. A class of optimal stopping problems for sampling without replacement. Biometrika, 61(2):361–368, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334366.

Berry:1975:DCE

[Ber75] G. Berry. Design of carcinogenesis experiments using the Weibull distribution. Biometrika, 62(2):321–328, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335367.

Besag:1972:CSS

[Bes72] J. E. Besag. On the correlation structure of some two-dimensional stationary processes. Biometrika, 59(1):43–48, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334613.

Best:1973:ETK

[Bes73] D. J. Best. Extended tables for Kendall’s Tau. Biometrika, 60 (2):429–430, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334561.

Best:1974:VIB

[Bes74] D. J. Best. The variance of the inverse binomial estimator. Biometrika, 61(2):385–386, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334371.

Besag:1977:EPE

[Bes77] Julian Besag. Efficiency of pseudolikelihood estimation for simple Gaussian fields. Biometrika, 64(3):616–618, December 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345341.

Bohrer:1972:SOS

[BF72] Robert Bohrer and George K. Francis. Sharp one-sided confidence bounds for linear regression over intervals. Biometrika, 59(1):99– 107, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334620.

Bland:1966:DRM

[BGKO66] R. P. Bland, R. D. Gilbert, C. H. Kapadia, and D. B. Owen. On the distributions of the range and mean range for samples from a normal distribution. Biometrika, 53(1/2):245–248, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334072. See cor-

rection [BGKO71].

Bland:1971:CDR

[BGKO71] R. P. Bland, R. D. Gilbert, C. H. Kapadia, and D. B. Owen. Correction: “On the Distributions of the Range and Mean Range for Samples from a Normal Distribution”. Biometrika, 58(1):245, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334344. See

[BGKO66].

Bailey:1977:IEC

[BGP77] R. A. Bailey, F. H. L. Gilchrist, and H. D. Patterson. Identification of effects and confounding patterns in factorial designs. Biometrika, 64(2):347–354, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335703.

Barnett:1976:CCE

[BGR76] V. Barnett, P. J. Green, and A. Robinson. Concomitants and correlation estimates. Biometrika, 63(2):323–328, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335626.

Brown:1974:TP

[BH74] Susannah Brown and P. Holgate. The thinned plantation. Biometrika, 61(2):253–261, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334352.

Bhansali:1976:EMA

[Bha76] R. J. Bhansali. Estimation of the moving average representation of a stationary nondeterministic process. Biometrika, 63(2):408–410, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335639.

Bhoj:1978:TEM

[Bho78] Dinesh S. Bhoj. Testing equality of means of correlated variates with missing observations on both responses. Biometrika, 65(1):225–228, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335301.

Bhoj:1979:TEV

[Bho79] Dinesh S. Bhoj. Testing equality of variances of correlated variates with incomplete data on both responses. Biometrika, 66 (3):681–683, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335195.

Billard:1977:TPS

[Bil77] L. Billard. A truncated partial sequential procedure. Biometrika, 64(3):567–572, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345334.

Binns:1975:SEM

[Bin75] Michael Binns. Sequential estimation of the mean of a negative binomial distribution. Biometrika, 62(2):433–440, August 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335384. See correction

[Bin76].

Binns:1976:CAS

[Bin76] M. R. Binns. Corrections and amendments: “Sequential estimation of the mean of a negative binomial distribution” (Biometrika 62 (1975), 433–440). Biometrika, 63(1):218, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335112. See [Bin75].

Binder:1978:BCA

[Bin78] D. A. Binder. Bayesian cluster analysis. Biometrika, 65(1):31–38, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335273.

Birnbaum:1975:CPJ

[Bir75] Allan Birnbaum. Comments on paper by J. D. Kalbfleisch. Biometrika, 62(2):262–264, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335359. See [Kal75b].

Bissell:1972:ANB

[Bis72a] A. F. Bissell. Another negative binomial model with varying element sizes. Biometrika, 59(3):691–693, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334824.

Bissell:1972:NBM

[Bis72b] A. F. Bissell. A negative binomial model with varying element sizes. Biometrika, 59(2):435–441, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334588.

Buckley:1979:LRC

[BJ79] Jonathan Buckley and Ian James. Linear regression with censored data. Biometrika, 66(3):429–436, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335161.

Bennett:1974:SSP

[BK74] B. M. Bennett and C. Kaneshiro. On the small-sample properties of the Mantel-Haenszel test for relative risk. Biometrika, 61 (2):233–236, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334350.

Bjornham:1976:FEC

[BL76] Ȧke Björnham and Georg Lindgren. Frequency estimation from crossings of an observed mean level. Biometrika, 63(3):507–512, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335728.

Blight:1970:ECS

[Bli70] B. J. N. Blight. Estimation from a censored sample for the exponential family. Biometrika, 57(2):389–395, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334847.

Blight:1973:SAF

[Bli73] B. J. N. Blight. Sampling from an autocorrelated finite population. Biometrika, 60(2):375–385, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334551.

Bloomfield:1972:EPT

[Blo72] P. Bloomfield. On the error of prediction of a time series. Biometrika, 59(3):501–507, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334801.

Bloomfield:1973:EMS

[Blo73] P. Bloomfield. An exponential model for the spectrum of a scalar time series. Biometrika, 60(2):217–226, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334533.

Blom:1976:SPI

[Blo76] Gunnar Blom. Some properties of incomplete U-statistics. Biometrika, 63(3):573–580, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335738.

Blom:1978:PMV

[Blo78] Gunnar Blom. A property of minimum variance estimates. Biometrika, 65(3):642–643, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335918.

Bennett:1972:CEB

[BM72] G. Kemble Bennett and Harry F. Martz, Jr. A continuous empirical Bayes smoothing technique. Biometrika, 59(2):361–368, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334580.

Besag:1975:ETS

[BM75] J. E. Besag and P. A. P. Moran. On the estimation and testing of spatial interaction in Gaussian lattice processes. Biometrika, 62(3):555–562, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335510.

Bingham:1978:SCD

[BM78a] Christopher Bingham and K. V. Mardia. A small circle distribution on the sphere. Biometrika, 65(2):379–389, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335218.

Blunck:1978:SEF

[BM78b] M. Blunck and T. P. Mommsen. Systematic errors in fitting linear transformations of the Michaelis–Menten equation. Biometrika, 65(2):363–368, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335216.

Begg:1979:SAC

[BM79] Colin B. Begg and Cyrus R. Mehta. Sequential analysis of comparative clinical trials. Biometrika, 66(1):97–103, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335248.

Breth:1978:DFL

[BMW78] M. Breth, J. S. Maritz, and E. J. Williams. On distributionfree lower confidence limits for the mean of a nonnegative random variable. Biometrika, 65(3):529–534, December 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335904. See correction

[BMW79].

Breth:1979:ACD

[BMW79] M. Breth, J. S. Maritz, and E. J. Williams. Amendments and corrections: “On Distribution-Free Lower Confidence Limits for the Mean of a Nonnegative Random Variable”. Biometrika, 66(3):689, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335198.

See [BMW78].

Barndorff-Nielsen:1973:A

[BN73] O. Barndorff-Nielsen. On M-ancillarity. Biometrika, 60(3):447– 455, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2334993.

Barndorff-Nielsen:1975:CPJ

[BN75] Ole Barndorff-Nielsen. Comments on paper by J. D. Kalbfleisch. Biometrika, 62(2):261–262, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335358. See [Kal75b].

Barndorff-Nielsen:1976:N

[BN76] O. Barndorff-Nielsen. Nonformation. Biometrika, 63(3):567–571, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335737.

Barndorff-Nielsen:1978:CPB

[BN78] Ole Barndorff-Nielsen. Comments on paper by B. Efron and D. V. Hinkley. Biometrika, 65(3):482–483, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335894. See [EH78b].

Blight:1975:BAM

[BO75] B. J. N. Blight and L. Ott. A Bayesian approach to model inadequacy for polynomial regression. Biometrika, 62(1):79–88, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334489.

Burnham:1978:ESC

[BO78] K. P. Burnham and W. S. Overton. Estimation of the size of a closed population when capture probabilities vary among animals. Biometrika, 65(3):625–633, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335915. See correction [Ano81].

Bohrer:1973:MPI

[Boh73] Robert Bohrer. A multivariate t probability integral. Biometrika, 60(3):647–654, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335016.

Boomsma:1975:TEE

[Boo75] A. Boomsma. Table erratum: “An error in F. N. David’s tables of the correlation coefficient” (Tables of the ordinates and probability integral of the distribution of the correlation coefficient in small samples. Cambridge Univ. Press, 1938). Biometrika, 62(3):711, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335542.

Bowman:1972:TSS

[Bow72] K. O. Bowman. Tables of the sample size requirement. Biometrika, 59(1):234, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2334643.

Bowman:1973:PKS

[Bow73] K. O. Bowman. Power of the kurtosis statistic, b_{2}, in tests of departures from normality. Biometrika, 60(3):623–628, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335013.

Box:1971:EDC

[Box71] M. J. Box. An experimental design criterion for precise estimation of a subset of the parameters in a nonlinear model. Biometrika, 58(1):149–153, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334325.

Brunk:1974:EDM

[BP74] H. D. Brunk and Donald A. Pierce. Estimation of discrete multivariate densities for computer-aided differential diagnosis of disease. Biometrika, 61(3):493–499, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334731.

Bock:1975:MCA

[BP75] R. Darrell Bock and Anne C. Petersen. A multivariate correction for attenuation. Biometrika, 62(3):673–678, December 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335527.

Bryant:1976:ETM

[BP76] J. L. Bryant and A. S. Paulson. An extension of Tukey’s method of multiple comparisons to experimental designs with random concomitant variables. Biometrika, 63(3):631–638, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335745.

Blight:1974:CBB

[BR74] B. J. N. Blight and P. V. Rao. The convergence of Bhattacharyya bounds. Biometrika, 61(1):137–142, April 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334295.

Bellhouse:1975:SSP

[BR75] D. R. Bellhouse and J. N. K. Rao. Systematic sampling in the presence of a trend. Biometrika, 62(3):694–697, December 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335534. See correction

[BR81].

Brown:1978:RLR

[BR78] D. Brown and P. Rothery. Randomness and local regularity of points in a plane. Biometrika, 65(1):115–122, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335285.

Bellhouse:1981:ACS

[BR81] D. R. Bellhouse and J. N. K. Rao. Amendments and corrections: “Systematic sampling in the presence of a trend” [Biometrika 62 (1975), no. 3, 694–697; MR 52 #15763]. Biometrika, 68(2):576, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335609.

See [BR75].

Breslow:1970:GKW

[Bre70] Norman Breslow. A generalized Kruskal–Wallis test for comparing K samples subject to unequal patterns of censorship. Biometrika, 57(3):579–594, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334776.

Breth:1979:NBI

[Bre79] M. Breth. Nonparametric Bayesian interval estimation. Biometrika, 66(3):641–644, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335186.

Brillinger:1974:ADW

[Bri74] David R. Brillinger. The asymptotic distribution of the Whittaker periodogram and a related chi-squared statistic for stationary processes. Biometrika, 61(3):419–422, December 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334722.

Brillinger:1977:IPN

[Bri77] David R. Brillinger. The identification of a particular nonlinear time series system. Biometrika, 64(3):509–515, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345326.

Brooks:1972:DTA

[Bro72] R. J. Brooks. A decision theory approach to optimal regression designs. Biometrika, 59(3):563–571, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334807.

Brown:1973:ADB

[Bro73] P. J. Brown. Aspects of design for binary key models. Biometrika, 60(2):309–318, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334542.

Brooks:1974:ERC

[Bro74a] Margaret A. Brooks. Exponential regression with correlated observations. Biometrika, 61(1):109–115, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334291.

Brooks:1974:BAT

[Bro74b] R. J. Brooks. Bayesian analysis of the two-sample problem under the Lehmann alternatives. Biometrika, 61(3):501–507, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334732.

Brooks:1974:CEP

[Bro74c] R. J. Brooks. On the choice of an experiment for prediction in linear regression. Biometrika, 61(2):303–311, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334359.

Brown:1974:RSR

[Bro74d] R. A. Brown. Robustness of the Studentized range statistic. Biometrika, 61(1):171–175, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334300.

Brown:1975:SCT

[Bro75] B. M. Brown. A short-cut test for outliers using residuals. Biometrika, 62(3):623–629, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335519.

Brooks:1977:ORD

[Bro77a] R. J. Brooks. Optimal regression design for control in linear regression. Biometrika, 64(2):319–325, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335699.

Brown:1977:FST

[Bro77b] P. J. Brown. Functions for selecting tests in diagnostic key construction. Biometrika, 64(3):589–596, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345337.

Brown:1978:GLS

[Bro78] G. H. Brown. Generalized least squares applied to the linear ultrastructural model. Biometrika, 65(2):441–444, August 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335228.

Brunk:1978:UDE

[Bru78] H. D. Brunk. Univariate density estimation by orthogonal series. Biometrika, 65(3):521–528, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335903.

Barr:1973:NKS

[BS73a] D. R. Barr and R. H. Shudde. A note on Kuiper’s V_{n} statistic. Biometrika, 60(3):663–664, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335018.

Bowman:1973:NDS

[BS73b] K. O. Bowman and L. R. Shenton. Notes on the distribution of \sqrt{b}_{1} in sampling from Pearson distributions. Biometrika, 60(1):155–167, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334917.

Bowman:1973:SSN

[BS73c] K. O. Bowman and L. R. Shenton. Space of solutions for a normal mixture. Biometrika, 60(3):629–636, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335014.

Bowman:1975:OTC

[BS75] K. O. Bowman and L. R. Shenton. Omnibus test contours for departures from normality based on \sqrt{b_{1}} and b_{2}. Biometrika, 62 (2):243–250, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335355.

Basawa:1976:ETB

[BS76a] I. V. Basawa and D. J. Scott. Efficient tests for branching processes. Biometrika, 63(3):531–536, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335731.

Booth:1976:BAS

[BS76b] N. B. Booth and A. F. M. Smith. Batch acceptance schemes based on an autoregressive prior. Biometrika, 63(1):133–136, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335093.

Bowman:1979:APP

[BS79] K. O. Bowman and L. R. Shenton. Approximate percentage points for Pearson distributions. Biometrika, 66(1):147–151, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335254.

Beichl:2000:MA

[BS00] Isabel Beichl and Francis Sullivan. The Metropolis algorithm. Computing in Science and Engineering, 2(1):65–69, January/ February 2000. CODEN CSENFA. ISSN 1521-9615 (print),

1558-366X (electronic). URL http://dlib.computer.org/cs/
books/cs2000/pdf/c1065.pdf;   http://www.computer.org/
cse/cs1999/c1065abs.htm.

Beckman:1973:UPM

[BT73] R. J. Beckman and G. L. Tietjen. Upper 10% and 25% points of the maximum F ratio. Biometrika, 60(1):213–214, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334928.

Bechhofer:1974:IIR

[BT74] Robert E. Bechhofer and Ajit C. Tamhane. An iterated integral representation for a multivariate normal integral having block covariance structure. Biometrika, 61(3):615–619, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334747.

Box:1977:CAM

[BT77] G. E. P. Box and G. C. Tiao. A canonical analysis of multiple time series. Biometrika, 64(2):355–365, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335704.

Bellhouse:1977:TSS

[BTG77] D. R. Bellhouse, M. E. Thompson, and V. P. Godambe. Twostage sampling with exchangeable prior distributions. Biometrika, 64(1):97–103, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335777.

Buhrman:1977:TCI

[Buh77] J. M. Buhrman. Tests and confidence intervals for the difference and ratio of two probabilities. Biometrika, 64(1):160–162, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335791.

Bukac:1972:FCU

[Buk72] J. Bukac. Fitting S_{B} curves using symmetrical percentile points. Biometrika, 59(3):688–690, December 1972. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334822.

Bacon:1971:ETB

[BW71] David W. Bacon and Donald G. Watts. Estimating the transition between two intersecting straight lines. Biometrika, 58(3):525–534, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334387.

Berenblut:1974:EDP

[BW74] I. I. Berenblut and G. I. Webb. Experimental design in the presence of autocorrelated errors. Biometrika, 61(3):427–437, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334724.

Bloomfield:1975:ILS

[BW75] Peter Bloomfield and Geoffrey S. Watson. The inefficiency of least squares. Biometrika, 62(1):121–128, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334493.

Bryant:1978:ABC

[BW78] Peter Bryant and John A. Williamson. Asymptotic behaviour of classification maximum likelihood estimates. Biometrika, 65 (2):273–281, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335205.

Burr:1977:CVE

[BY77] P. C. Burr and D. H. Young. Critical values for the exponential scores trend statistic. Biometrika, 64(3):640–642, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345349.

Cameron:1978:PVR

[Cam78] M. A. Cameron. The prediction variance and related statistics for stationary time series. Biometrika, 65(2):283–296, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335206.

Carpenter:1977:MWC

[Car77] R. G. Carpenter. Matching when covariables are normally distributed. Biometrika, 64(2):299–307, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335697.

Carlson:1977:AAB

[CC77] Roger Carlson and Michael A. Camp. Anomalies arising from a Bayesian analysis of Dempster’s upper and lower probabilities. Biometrika, 64(3):651–653, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345354.

Covey-Crump:1970:ORD

[CCS70] P. A. K. Covey-Crump and S. D. Silvey. Optimal regression designs with previous observations. Biometrika, 57(3):551–566, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334773.

Copas:1978:RLB

[CD78] J. B. Copas and J. M. Dickey. A remark on likelihood, Bayes and affinity for a preconceived value. Biometrika, 65(3):648–649, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335921.

Crump:1972:NEA

[CH72] Kenny S. Crump and Richard B. Howe. Nonparametric estimation of the age of a Galton-Watson branching process. Biometrika, 59(3):533–538, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334804.

Crump:1974:MME

[CH74] Kenny S. Crump and David G. Hoel. Mathematical models for estimating mutation rates in cell populations. Biometrika, 61 (2):237–252, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334351.

Chase:1975:SDE

[CH75] Gerald R. Chase and David G. Hoel. Serial dilutions: Error effects and optimal designs. Biometrika, 62(2):329–334, August 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335368.

Cameron:1979:TS

[CH79] M. A. Cameron and E. J. Hannan. Transient signals. Biometrika, 66(2):243–258, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335655.

Chambers:1973:CAF

[Cha73a] John M. Chambers. Corrections and amendments: “Fitting Nonlinear Models: Numerical Techniques”. Biometrika, 60(3):669, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335021.

See [Cha73b].

Chambers:1973:FNM

[Cha73b] John M. Chambers. Fitting nonlinear models: Numerical techniques. Biometrika, 60(1):1–13, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334900. See correction [Cha73a].

Chant:1974:ATC

[Cha74a] D. Chant. On asymptotic tests of composite hypotheses in nonstandard conditions. Biometrika, 61(2):291–298, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334357.

Chase:1974:TOA

[Cha74b] Gerald R. Chase. On testing for ordered alternatives with increased sample size for a control. Biometrika, 61(3):569–578, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334741.

Chang:1976:EAV

[Cha76] W. C. Chang. The effects of adding a variable in dissecting a mixture of two normal populations with a common covariance matrix. Biometrika, 63(3):676–678, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335752.

Chan:1977:UPF

[Cha77] N. N. Chan. On an unbiased predictor in factor analysis. Biometrika, 64(3):642–644, December 1977. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345350.

Chen:1971:SPC

[Che71] C. W. Chen. On some problems in canonical correlation analysis. Biometrika, 58(2):399–400, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334529.

Chiu:1974:RMN

[Chi74] W. K. Chiu. The ranking of means of normal populations for a generalized selection goal. Biometrika, 61(3):579–584, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334742.

Chmiel:1976:SPS

[Chm76] Joseph J. Chmiel. Some properties of Spearman-type estimators of the variance and percentiles in bioassay. Biometrika, 63 (3):621–626, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335743.

Choi:1977:TED

[Cho77] S. C. Choi. Tests of equality of dependent correlation coefficients. Biometrika, 64(3):645–647, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345351.

Clemmer:1968:UEB

[CK68] Bennie A. Clemmer and Richard G. Krutchkoff. The use of empirical Bayes estimators in a linear regression model. Biometrika, 55(3):525–534, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334257. See correction [CK70].

Clemmer:1970:CUE

[CK70] B. A. Clemmer and R. G. Krutchkoff. Correction: “The Use of Empirical Bayes Estimators in a Linear Regression Model”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334860. See [CK68].

Cox:1976:CDR

[CL76] Trevor F. Cox and Toby Lewis. A conditioned distance ratio method for analyzing spatial patterns. Biometrika, 63(3):483–491, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335725.

Clayton:1974:SOR

[Cla74] D. G. Clayton. Some odds ratio statistics for the analysis of ordered categorical data. Biometrika, 61(3):525–531, December 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334735.

Clayton:1976:ORC

[Cla76] D. G. Clayton. An odds ratio comparison for ordered categorical data with censored observations. Biometrika, 63(2):405–408, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335638.

Clayton:1978:MAB

[Cla78] D. G. Clayton. A model for association in bivariate life tables and its application in epidemiological studies of familial tendency in chronic disease incidence. Biometrika, 65(1):141–151, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335289.

Cobb:1978:PNC

[Cob78] George W. Cobb. The problem of the Nile: Conditional solution to a changepoint problem. Biometrika, 65(2):243–251, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335202.

Coleman:1972:SPL

[Col72] Rodney Coleman. Sampling procedures for the lengths of random straight lines. Biometrika, 59(2):415–426, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334586.

Cooke:1979:SIB

[Coo79] Peter Cooke. Statistical inference for bounds of random variables. Biometrika, 66(2):367–374, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335672.

Copas:1972:EBM

[Cop72] J. B. Copas. Empirical Bayes methods and the repeated use of a standard. Biometrika, 59(2):349–360, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334579.

Copas:1973:RMM

[Cop73] J. B. Copas. Randomization models for the matched and unmatched 2 × 2 tables. Biometrika, 60(3):467–476, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334995.

Copas:1975:ULC

[Cop75] J. B. Copas. On the unimodality of the likelihood for the Cauchy distribution. Biometrika, 62(3):701–704, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335537.

Cornell:1972:NEL

[Cor72] John A. Cornell. A note on the equality of least squares estimates using second-order equiradial rotatable designs. Biometrika, 59(3):686–687, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334821.

Cormack:1977:ICL

[Cor77] R. M. Cormack. The invariance of Cox and Lewis’s statistic for the analysis of spatial patterns. Biometrika, 64(1):143–144, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335785.

Cotter:1979:SDF

[Cot79] Sarah C. Cotter. A screening design for factorial experiments with interactions. Biometrika, 66(2):317–320, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335664.

Cox:1970:BRB

[Cox70a] D. R. Cox. Book review: A Dictionary and Bibliography of Discrete Distributions, by G. P. Patil; S. W. Joshi; Discrete Distributions. by N. L. Johnson; S. Kotz. Biometrika, 57(2):465–466, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334859.

Cox:1970:CC

[Cox70b] D. R. Cox. The continuity correction. Biometrika, 57(1):217–219, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334957.

Cox:1971:NPR

[Cox71] D. R. Cox. A note on polynomial response functions for mixtures. Biometrika, 58(1):155–159, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334326.

Cox:1972:BRBa

[Cox72a] D. R. Cox. Book review: Foundations of Statistical Inference, by V. P. Godambe; D. A. Sprott. Biometrika, 59(1):235, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334644.

Cox:1972:BRBb

[Cox72b] D. R. Cox. Book review: The Theory of Statistical Inference, by S. Zacks. Biometrika, 59(1):237, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334646.

Cox:1975:NDS

[Cox75a] D. R. Cox. A note on data-splitting for the evaluation of significance levels. Biometrika, 62(2):441–444, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335385.

Cox:1975:NEB

[Cox75b] D. R. Cox. A note on empirical Bayes inference in a finite Poisson process. Biometrika, 62(3):709–711, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335541.

Cox:1975:NPB

[Cox75c] D. R. Cox. A note on partially Bayes inference and the linear model. Biometrika, 62(3):651–654, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335523.

Cox:1975:PL

[Cox75d] D. R. Cox. Partial likelihood. Biometrika, 62(2):269–276, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335362.

Cox:1976:LSR

[Cox76a] N. R. Cox. The linear structural relation for several groups of data. Biometrika, 63(2):231–237, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335614. See correction [Cox77].

Cox:1976:NDN

[Cox76b] N. R. Cox. A note on the determination of the nature of turning points of likelihoods. Biometrika, 63(1):199–201, April 1976. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335104.

Cox:1976:RED

[Cox76c] Trevor F. Cox. The robust estimation of the density of a forest stand using a new conditioned distance method. Biometrika, 63(3):493–499, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335726.

Cox:1977:CAL

[Cox77] N. R. Cox. Corrections and amendments: “The Linear Structural Relation for Several Groups of Data”. Biometrika, 64(2):427, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335715. See

[Cox76a].

Cox:1979:NGA

[Cox79] D. R. Cox. A note on the graphical analysis of survival data. Biometrika, 66(1):188–190, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335265.

Carnal:1972:OSD

[CR72] H. Carnal and H. Riedwyl. On a one-sample distribution-free test statistic V. Biometrika, 59(2):465–467, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334592.

Cramer:1972:SHP

[Cra72] Harald Cramér. Studies in the history of probability and statistics. XXVIII. On the history of certain expansions used in mathematical statistics. Biometrika, 59(1):205–207, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334631.

Cressie:1976:LHO

[Cre76] Noel Cressie. On the logarithms of high-order spacings. Biometrika, 63(2):343–355, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335629.

Cressie:1978:PRT

[Cre78] Noel Cressie. Power results for tests based on high-order gaps. Biometrika, 65(1):214–218, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335298.

Cressie:1979:OSB

[Cre79] Noel Cressie. An optimal statistic based on higher order gaps. Biometrika, 66(3):619–627, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335184.

Crowley:1974:NSR

[Cro74] John Crowley. A note on some recent likelihoods leading to the log rank test. Biometrika, 61(3):533–538, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334736.

Cox:1971:TSC

[CS71] D. R. Cox and E. J. Snell. On test statistics calculated from residuals. Biometrika, 58(3):589–594, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334393.

Clifford:1973:MSC

[CS73] Peter Clifford and Aidan Sudbury. A model for spatial conflict. Biometrika, 60(3):581–588, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335008.

Conniffe:1974:EFC

[CS74] D. Conniffe and Joan Stone. The efficiency factor of a class of incomplete block designs. Biometrika, 61(3):633–636, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334752.

Conniffe:1975:SIB

[CS75] D. Conniffe and Joan Stone. Some incomplete block designs of maximum efficiency. Biometrika, 62(3):685–686, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335530. See correction

[CS76].

Conniffe:1976:CAS

[CS76] D. Conniffe and J. Stone. Corrections and amendments: “Some incomplete block designs of maximum efficiency” (Biometrika 62 (1975), no. 3, 685–686). Biometrika, 63(3):686, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335759. See [CS75].

Cox:1978:TMN

[CS78] D. R. Cox and N. J. H. Small. Testing multivariate normality. Biometrika, 65(2):263–272, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335204.

Clifford:1979:UBS

[CS79a] Peter Clifford and Aidan W. Sudbury. On the use of bounds in the statistical analysis of spatial processes. Biometrika, 66(3):495–503, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335168.

Cox:1979:SER

[CS79b] D. R. Cox and E. J. Snell. On sampling and the estimation of rare errors. Biometrika, 66(1):125–132, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335251. See correction [CS82].

Cox:1982:ACS

[CS82] D. R. Cox and E. J. Snell. Amendments and corrections: “On sampling and the estimation of rare errors” [Biometrika 66 (1979), no. 1, 125–132; MR 81f:62015]. Biometrika, 69(2):491, August 1982.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335430. See [CS79b].

Cassel:1976:SRG

[CSW76] Claes M. Cassel, Carl E. Särndal, and Jan H. Wretman. Some results on generalized difference estimation and generalized regression estimation for finite populations. Biometrika, 63(3):615–620, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335742.

With comments by T. M. F. Smith and a reply by the authors.

Cooper:1977:NEP

[CT77] D. M. Cooper and R. Thompson. A note on the estimation of the parameters of the autoregressive-moving average process. Biometrika, 64(3):625–628, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345344.

DAgostino:1970:TNN

[D’A70a] Ralph B. D’Agostino. Transformation to normality of the null distribution of g_{1}. Biometrika, 57(3):679–681, December 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334794.

Davis:1970:EPT

[DA70b] R. L. Davis and J. C. Arnold. An efficient preliminary test estimator for the variance of a normal population when the mean is unknown. Biometrika, 57(3):674–677, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334792.

DAgostino:1971:OTN

[D’A71] Ralph B. D’Agostino. An omnibus test of normality for moderate and large size samples. Biometrika, 58(2):341–348, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334522.

DAgostino:1972:SSP

[D’A72] Ralph B. D’Agostino. Small sample probability points for the D test of normality. Biometrika, 59(1):219–221, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334638.

Dahiya:1971:PCS

[Dah71] Ram C. Dahiya. On the Pearson chi-squared goodness-of-fit test statistic. Biometrika, 58(3):685–686, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334410.

Dalal:1978:SCP

[Dal78] S. R. Dalal. Simultaneous confidence procedures for univariate and multivariate Behrens-Fisher type problems. Biometrika, 65(1):221– 225, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335300.

Daniels:1969:RRT

[Dan69] H. E. Daniels. Round-robin tournament scores. Biometrika, 56 (2):295–299, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2334422. See correction [Dan70].

Daniels:1970:CRR

[Dan70] H. E. Daniels. Correction: “Round-Robin Tournament Scores”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334862. See [Dan69].

Daniels:1972:ERT

[Dan72] H. E. Daniels. An exact relation in the theory of carrier-borne epidemics. Biometrika, 59(1):211–213, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334634.

Darroch:1971:DIB

[Dar71] J. N. Darroch. A definition of independence for bounded-sum, nonnegative, integer-valued variables. Biometrika, 58(2):357–368, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334524.

Darroch:1974:MAI

[Dar74] J. N. Darroch. Multiplicative and additive interaction in contingency tables. Biometrika, 61(2):207–214, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334348.

Davis:1970:EDH

[Dav70a] A. W. Davis. Exact distributions of Hotelling’s generalized T_{0}^{2}. Biometrika, 57(1):187–191, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334948. See correction [Dav72].

Davis:1970:PAO

[Dav70b] A. W. Davis. Percentile approximations for ordered F ratios. Biometrika, 57(2):457–459, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334856.

Davis:1971:PAC

[Dav71] A. W. Davis. Percentile approximations for a class of likelihood ratio criteria. Biometrika, 58(2):349–356, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334523.

Davis:1972:CAE

[Dav72] A. W. Davis. Corrections and amendments: “Exact distributions of Hotelling’s generalized T_{0}^{2}” (Biometrika 57 (1970), 187– 191). Biometrika, 59(2):498, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334603. See [Dav70a].

Davies:1973:NIC

[Dav73] R. B. Davies. Numerical inversion of a characteristic function. Biometrika, 60(2):415–417, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334555.

Davenport:1975:TME

[Dav75] James M. Davenport. Two methods of estimating the degrees of freedom of an approximate F. Biometrika, 62(3):682–684, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335529. See

corrections and amendments [Dav76a].

Davenport:1976:CAT

[Dav76a] James M. Davenport. Corrections and amendments: “Two methods of estimating the degrees of freedom of an approximate F ” (Biometrika 62 (1975), no. 3, 682–684). Biometrika, 63(3):685, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335758.

See [Dav75].

Davis:1976:SDU

[Dav76b] A. W. Davis. Statistical distributions in univariate and multivariate Edgeworth populations. Biometrika, 63(3):661–670, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335749.

Davies:1977:HTW

[Dav77] R. B. Davies. Hypothesis testing when a nuisance parameter is present only under the alternative. Biometrika, 64(2):247–254, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335690.

Davis:1979:DEM

[Dav79] A. W. Davis. On the differential equation for Meijer’s G_{p, p}^{p,0} function, and further tables of Wilks’s likelihood ratio criterion. Biometrika, 66(3):519–531, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335171.

Dawid:1973:PEL

[Daw73] A. P. Dawid. Posterior expectations for large observations. Biometrika, 60(3):664–667, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335019.

Dawid:1977:IDA

[Daw77] A. P. Dawid. Invariant distributions and analysis of variance models. Biometrika, 64(2):291–297, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335696.

Davidson:1969:MPC

[DB69] Roger R. Davidson and Ralph A. Bradley. Multivariate paired comparisons: The extension of a univariate model and associated estimation and test procedures. Biometrika, 56(1):81–95, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334693. See cor-

rections [DB70].

Davidson:1970:CMP

[DB70] Roger R. Davidson and Ralph A. Bradley. Corrections: “Multivariate Paired Comparisons: The Extension of a Univariate Model and Associated Estimation and Test Procedures”. Biometrika, 57(1): 225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334961.

See [DB69].

Davidson:1971:RRM

[DB71] Roger R. Davidson and Ralph A. Bradley. A regression relationship for multivariate paired comparisons. Biometrika, 58(3):555–560, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334390.

Pereira:1977:NCF

[dBP77] Basilio de B. Pereira. A note on the consistency and on the finite sample comparisons of some tests of separate families of hypotheses. Biometrika, 64(1):109–113, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335779. See correction [De 77].

Pereira:1978:TES

[dBP78] Basilio de B. Pereira. Tests and efficiencies of separate regression models. Biometrika, 65(2):319–327, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335210.

deCani:1971:NRP

[dC71] John S. de Cani. On the number of replications of a paired comparison. Biometrika, 58(1):169–175, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334328.

Dolby:1977:CAU

[DC77] G. R. Dolby and N. R. Cox. Corrections and amendments: “The ultrastructural relation: a synthesis of the functional and structural relations” (Biometrika 63 (1976), no. 1, 39–50) by Dolby; “The linear structural relation for several groups of data” (Biometrika 63 (1976), no. 2, 231–237) by N. R. Cox. Biometrika, 64(2):427, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335714.

See [Dol76].

Dent:1978:MML

[DC78] Warren T. Dent and Shirley Cassing. On modified maximum likelihood estimators of the autocorrelation coefficient in linear models. Biometrika, 65(1):211–213, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335297.

DeBPereira:1977:CAN

[De 77] B. De B. Pereira. Corrections and amendments: “A Note on the Consistency and on the Finite Sample Comparisons of Some Tests of Separate Families of Hypotheses”. Biometrika, 64(3):655, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345357.

See [dBP77].

deCani:1972:BBA

[deC72] John S. deCani. A branch and bound algorithm for maximum likelihood paired comparison ranking. Biometrika, 59(1):131–135, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334623.

Denton:1972:DCB

[Den72] Gillian M. Denton. On Downton’s carrier-borne epidemic. Biometrika, 59(2):455–461, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334590.

Devore:1973:NRR

[Dev73] Jay L. Devore. The naive rule for reconstructing a noisy Markov chain. Biometrika, 60(2):227–233, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334534.

Dey:1977:CRG

[Dey77] A. Dey. Construction of regular group divisible designs. Biometrika, 64(3):647–649, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2345352.

Davis:1979:GPL

[DF79] Henry T. Davis and Michael L. Feldstein. The generalized Pareto law as a model for progressively censored survival data. Biometrika, 66(2):299–306, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335662.

Doksum:1977:PTS

[DFA77] Kjell A. Doksum, Grete Fenstad, and Rolf Aaberge. Plots and tests for symmetry. Biometrika, 64(3):473–487, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345323.

Dahiya:1972:PCS

[DG72] Ram C. Dahiya and John Gurland. Pearson chi-squared test of fit with random intervals. Biometrika, 59(1):147–153, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334625.

Downham:1976:ITD

[DG76] D. Y. Downham and D. H. Green. Inference for a two-dimensional stochastic growth model. Biometrika, 63(3):551–554, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335734.

Draper:1971:DCR

[DGK71] N. R. Draper, I. Guttman, and H. Kanemasu. The distribution of certain regression statistics. Biometrika, 58(2):295–298, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334517.

Devlin:1975:REO

[DGK75] Susan J. Devlin, R. Gnanadesikan, and J. R. Kettenring. Robust estimation and outlier detection with correlation coefficients. Biometrika, 62(3):531–545, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335508.

Duby:1977:PDE

[DGP77] Camille Duby, Xavier Guyon, and Bernard Prum. The precision of different experimental designs for a random field. Biometrika, 64(1):59–66, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335771.

Desu:1974:TSE

[DGS74] Mahamunulu M. Desu, Edmund A. Gehan, and Norman C. Severo. A two-stage estimation procedure when there May be spurious observations. Biometrika, 61(3):593–599, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334744.

Davies:1975:EEI

[DH75a] R. B. Davies and B. Hutton. The effect of errors in the independent variables in linear regression. Biometrika, 62(2):383–391, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335377. See

correction [DH77].

Delcoigne:1975:SCD

[DH75b] A. Delcoigne and P. Hansen. Sequence comparison by dynamic programming. Biometrika, 62(3):661–664, December 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335525.

Davies:1977:CAE

[DH77] R. B. Davies and B. Hutton. Corrections and amendments: “The Effect of Errors in the Independent Variables in Linear Regression”. Biometrika, 64(3):655, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345355. See [DH75a].

Diggle:1975:RDE

[Dig75] Peter J. Diggle. Robust density estimation using distance methods. Biometrika, 62(1):39–48, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334485.

Diggle:1977:NRD

[Dig77] Peter J. Diggle. A note on robust density estimation for spatial point patterns. Biometrika, 64(1):91–95, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335776.

deJong:1976:RFA

[dJ76] Piet de Jong. The recursive fitting of autoregressions. Biometrika, 63(3):525–530, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335730.

deGraft-Johnson:1973:EDM

[dJS73] K. T. deGraft Johnson and J. Sedransk. Estimation of domain means using two-phase sampling. Biometrika, 60(2):387–393, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334552.

Dolby:1972:MLE

[DL72] G. R. Dolby and S. Lipton. Maximum likelihood estimation of the general nonlinear functional relationship with replicated observations and correlated errors. Biometrika, 59(1):121–129, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334622.

Duncan:1973:MCS

[DL73] G. T. Duncan and M. W. J. Layard. A Monte-Carlo study of asymptotically robust tests for correlation coefficients. Biometrika, 60(3):551–558, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335004.

David:1972:PFR

[DLB72] H. A. David, P. A. Lachenbruch, and H. P. Brandis. The power function of range and Studentized range tests in normal samples. Biometrika, 59(1):161–168, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334627.

Davies:1979:SPS

[DN79] Neville Davies and Paul Newbold. Some power studies of a portmanteau test of time series model specification. Biometrika, 66 (1):153–155, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335255.

Dolby:1976:URS

[Dol76] G. R. Dolby. The ultrastructural relation: A synthesis of the functional and structural relations. Biometrika, 63(1):39–50, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335082. See cor-

rection [DC77].

Domb:1972:NSE

[Dom72] C. Domb. A note on the series expansion method for clustering problems. Biometrika, 59(1):209–211, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334633.

Downs:1972:OS

[Dow72] Thomas D. Downs. Orientation statistics. Biometrika, 59(3): 665–676, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334817.

Downton:1976:NTB

[Dow76] F. Downton. Nonparametric tests for block experiments. Biometrika, 63(1):137–141, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335094.

Daw:1972:SHP

[DP72] R. H. Daw and E. S. Pearson. Studies in the history of probability and statistics. XXX. Abraham de Moivre’s 1733 derivation of the normal curve: a bibliographical note. Biometrika, 59(3):677–680, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334818.

DAgostino:1973:TDN

[DP73] Ralph D’Agostino and E. S. Pearson. Tests for departure from normality. Empirical results for the distributions of b_{2} and √b_{1}. Biometrika, 60(3):613–622, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335012. See correction [DP74].

DAgostino:1974:CAT

[DP74] R. D’Agostino and E. S. Pearson. Corrections and amendments: “Tests for departure from normality. Empirical results for the distributions of b_{2} and √b_{1}” (Biometrika 60 (1973), 613–622). Biometrika, 61(3):647, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334757. See [DP73].

Deshpande:1977:MSM

[DPA77] M. N. Deshpande and S. G. Prabhu-Ajgaonkar. On multitrial sampling methods. Biometrika, 64(2):422–424, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335712.

Davies:1974:CAM

[DPF74a] N. Davies, M. B. Pate, and M. G. Frost. Corrections and amendments: “Maximum Autocorrelations for Moving Average Processes”. Biometrika, 61(3):647, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334758. See [DPF74b].

Davies:1974:MAM

[DPF74b] N. Davies, M. B. Pate, and M. G. Frost. Maximum autocorrelations for moving average processes. Biometrika, 61(1):199–200, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334307. See

correction [DPF74a].

Darroch:1973:TIR

[DR73] J. N. Darroch and D. Ratcliff. Tests of F -independence with reference to quasi-independence and Waite’s fingerprint data. Biometrika, 60(2):395–401, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334553.

DAgostino:1974:PGT

[DR74] Ralph B. D’Agostino and Bernard Rosman. The power of Geary’s test of normality. Biometrika, 61(1):181–184, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334302.

Desu:1968:FSS

[DS68] M. Mahamunulu Desu and Milton Sobel. A fixed subset-size approach to the selection problem. Biometrika, 55(2):401–410, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334883. See cor-

rection [DS76a].

Dawid:1972:ECI

[DS72] A. P. Dawid and M. Stone. Expectation consistency of inverse probability distributions. Biometrika, 59(2):486–489, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334599.

Davidson:1973:BAP

[DS73] Roger R. Davidson and Daniel L. Solomon. A Bayesian approach to paired comparison experimentation. Biometrika, 60(3):477–487, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334996.

DeRouen:1975:MPC

[DS75] T. A. DeRouen and Y. R. Sarma. G_{1}-minimax procedures for the case of prior distributions in discriminant analysis. Biometrika, 62(2):403–406, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335380.

Desu:1976:CAF

[DS76a] M. M. Desu and Milton Sobel. Corrections and amendments: “A fixed subset-size approach to the selection problem” (Biometrika 55 (1968), 401–410). Biometrika, 63(3):685, December 1976. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335756. See [DS68].

Doksum:1976:PCG

[DS76b] Kjell A. Doksum and Gerald L. Sievers. Plotting with confidence: Graphical comparisons of two populations. Biometrika, 63(3):421–434, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335720.

DAgostino:1971:SPP

[DT71] Ralph B. D’Agostino and Gary L. Tietjen. Simulation probability points of b_{2} for small samples. Biometrika, 58(3):669–672, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334404.

Draper:1972:RPU

[DT72] Norman R. Draper and David E. Tierney. Regions of positive and unimodal series expansion of the Edgeworth and Gram-Charlier approximations. Biometrika, 59(2):463–465, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334591.

DAgostino:1973:AND

[DT73] Ralph B. D’Agostino and Gary L. Tietjen. Approaches to the null distribution of \sqrt{b}_{1}. Biometrika, 60(1):169–173, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334918.

Duran:1976:CLS

[DTL76] B. S. Duran, W. S. Tsai, and T. O. Lewis. A class of locationscale nonparametric tests. Biometrika, 63(1):173–176, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335099.

Davies:1977:SLB

[DTN77] N. Davies, C. M. Triggs, and P. Newbold. Significance levels of the Box-Pierce portmanteau statistic in finite samples. Biometrika, 64(3):517–522, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345327.

Dunsmore:1976:APA

[Dun76] I. R. Dunsmore. Asymptotic prediction analysis. Biometrika, 63 (3):627–630, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335744.

Durbin:1975:KST

[Dur75] James Durbin. Kolmogorov–Smirnov tests when parameters are estimated with applications to tests of exponentiality and tests on spacings. Biometrika, 62(1):5–22, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334482.

Dutt:1973:RMN

[Dut73] John E. Dutt. A representation of multivariate normal probability integrals by integral transforms. Biometrika, 60(3):637–645, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335015.

Dutt:1975:CPI

[Dut75] John E. Dutt. On computing the probability integral of a general multivariate t. Biometrika, 62(1):201–205, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334504.

Durbin:1971:TSC

[DW71] J. Durbin and G. S. Watson. Testing for serial correlation in least squares regression. III. Biometrika, 58(1):1–19, April 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334313.

Dyer:1974:CTN

[Dye74] Alan R. Dyer. Comparisons of tests for normality with a cautionary note. Biometrika, 61(1):185–189, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334303.

Eaves:1978:THF

[EB78] David Eaves and A. H. Burr. The threshold hypothesis for flash perception. Biometrika, 65(3):503–509, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335901.

Edwards:1973:LTL

[Edw73] A. W. F. Edwards. The likelihood treatment of linear regression. Biometrika, 60(1):73–77, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334907.

Efron:1971:FSE

[Efr71] Bradley Efron. Forcing a sequential experiment to be balanced. Biometrika, 58(3):403–417, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334377.

Efron:1979:RLB

[Efr79] B. Efron. The 1977 Rietz Lecture: Bootstrap methods: Another look at the jackknife. Annals of Statistics, 7(1):1–26, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://www.jstor.org/stable/2958830. See

[Que49, Que56, Tuk58, Mil74].

Efron:1978:RC

[EH78a] B. Efron and D. V. Hinkley. Reply to comments. Biometrika, 65(3):487, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335898.

Efron:1978:AAM

[EH78b] Bradley Efron and David V. Hinkley. Assessing the accuracy of the maximum likelihood estimator: Observed versus expected Fisher information. Biometrika, 65(3):457–482, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335893. See comments [BN78,

Jam78, Rob78a, Spr78] and reply [EH78a].

Ergle:1972:NEB

[EK72] William D. Ergle and Richard G. Krutchkoff. A nonparametric empirical Bayes estimator. Biometrika, 59(2):483–485, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334598.

Ekbohm:1976:CMPa

[Ekb76a] Gunnar Ekbohm. Comparing means in the paired case with missing data on one response. Biometrika, 63(1):169–172, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335098.

Ekbohm:1976:CMPb

[Ekb76b] Gunnar Ekbohm. On comparing means in the paired case with incomplete data on both responses. Biometrika, 63(2):299–304, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335623.

Elston:1975:CBC

[Els75] R. C. Elston. On the correlation between correlations. Biometrika, 62(1):133–140, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334495.

Efron:1972:EBV

[EM72] Bradley Efron and Carl Morris. Empirical Bayes on vector observations: An extension of Stein’s method. Biometrika, 59(2):335–347, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334578.

Engen:1974:SFM

[Eng74] Steinar Engen. On species frequency models. Biometrika, 61 (2):263–270, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334353.

Engen:1975:NGS

[Eng75] Steinar Engen. A note on the geometric series as a species frequency model. Biometrika, 62(3):697–699, December 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335535.

Enis:1973:R

[Eni73] Peter Enis. On the relation E(X) = E{E(X|Y)}. Biometrika, 60(2):432–433, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334563.

Eccleston:1975:COM

[ER75] J. Eccleston and K. Russell. Connectedness and orthogonality in multi-factor designs. Biometrika, 62(2):341–345, August 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335370.

Eccleston:1977:AON

[ER77] J. A. Eccleston and K. G. Russell. Adjusted orthogonality in nonorthogonal designs. Biometrika, 64(2):339–345, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335702.

El-Sayyad:1973:BSE

[ESF73] G. M. El-Sayyad and P. R. Freeman. Bayesian sequential estimation of a Poisson process rate. Biometrika, 60(2):289–296, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334540.

Efron:1976:ENU

[ET76] Bradley Efron and Ronald Thisted. Estimating the number of unsen species: How many words did Shakespeare know? Biometrika, 63(3):435–447, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335721.

Farewell:1977:MBV

[Far77] V. T. Farewell. A model for a binary variable with time-censored observations. Biometrika, 64(1):43–46, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335768.

Farewell:1978:JES

[Far78] V. T. Farewell. Jackknife estimation with structured data. Biometrika, 65(2):444–447, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335229.

Farewell:1979:SRE

[Far79] V. T. Farewell. Some results on the estimation of logistic models based on retrospective data. Biometrika, 66(1):27–32, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335238.

Fearn:1975:BAG

[Fea75] T. Fearn. A Bayesian approach to growth curves. Biometrika, 62(1):89–100, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334490.

Fearn:1977:TSM

[Fea77] T. Fearn. A two-stage model for growth curves which leads to Rao’s covariance adjusted estimators. Biometrika, 64(1):141–143, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335784.

Fedorov:1974:RPC

[Fed74] V. V. Fedorov. Regression problems with controllable variables subject to error. Biometrika, 61(1):49–56, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334284.

Fellner:1974:HEP

[Fel74] William H. Fellner. Heuristic estimation of probability densities. Biometrika, 61(3):485–492, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334730.

Feuerverger:1979:SMA

[Feu79] Andrey Feuerverger. On some methods of analysis for weather experiments. Biometrika, 66(3):655–658, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335188.

Finucan:1974:MTS

[FGS74] H. M. Finucan, R. F. Galbraith, and M. Stone. Moments without tears in simple random sampling from a finite population. Biometrika, 61(1):151–154, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334297.

Forsythe:1970:ECI

[FH70a] Alan Forsythe and J. A. Hartigan. Efficiency of confidence intervals generated by repeated subsample calculations. Biometrika, 57(3):629–639, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334781.

Fryer:1970:RSE

[FH70b] J. G. Fryer and D. Holt. On the robustness of the standard estimates of the exponential mean to contamination. Biometrika, 57(3):641–648, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334782.

Fienberg:1972:MRC

[Fie72] Stephen E. Fienberg. The multiple recapture census for closed populations and incomplete 2^{k} contingency tables. Biometrika, 59(3):591–603, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334810.

Finch:1979:DAP

[Fin79a] Peter D. Finch. Description and analogy in the practice of statistics. Biometrika, 66(2):195–206, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2335647. See comments [Hea79, Kem79] and reply [Fin79b].

Finch:1979:RC

[Fin79b] Peter D. Finch. Reply to comments. Biometrika, 66(2):208, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335650. See

[Fin79a].

Freeman:1972:TMS

[FK72] Harold Freeman and Arnold M. Kuzmack. Tables of multivariate t in six and more dimensions. Biometrika, 59(1):217–219, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334637.

Flueck:1974:BSA

[FK74] John A. Flueck and James F. Korsh. A branch search algorithm for maximum likelihood paired comparison ranking. Biometrika, 61(3):621–626, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334748.

Flehinger:1971:STA

[FL71] B. J. Flehinger and T. A. Louis. Sequential treatment allocation in clinical trials. Biometrika, 58(3):419–426, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334378.

Fienberg:1976:LLR

[FL76] Stephen E. Fienberg and Kinley Larntz. Log linear representation for paired and multiple comparisons models. Biometrika, 63 (2):245–254, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335616.

Fryer:1972:CSM

[FR72] J. G. Fryer and C. A. Robertson. A comparison of some methods for estimating mixed normal distributions. Biometrika, 59(3):639–648, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334815.

Feigin:1979:AAI

[FR79] Paul D. Feigin and Benjamin Reiser. On asymptotic ancillarity and inference for Yule and regular nonergodic processes. Biometrika, 66(2):279–283, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335659.

Freeman:1970:OBS

[Fre70] P. R. Freeman. Optimal Bayesian sequential estimation of the median effective dose. Biometrika, 57(1):79–89, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334938.

Freeman:1972:SES

[Fre72] P. R. Freeman. Sequential estimation of the size of a population. Biometrika, 59(1):9–17, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334610.

Freeman:1973:EDU

[Fre73a] G. H. Freeman. Experimental designs with unequal concurrences for estimating direct and remote effects of treatments. Biometrika, 60(3):559–563, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335005.

Freeman:1973:NCB

[Fre73b] P. R. Freeman. A numerical comparison between sequential tagging and sequential recapture. Biometrika, 60(3):499–508, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334998.

Freeman:1973:SR

[Fre73c] P. R. Freeman. Sequential recapture. Biometrika, 60(1):141–153, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334916.

Freeman:1976:CMC

[Fre76] G. H. Freeman. A cyclic method of constructing regular group divisible incomplete block designs. Biometrika, 63(3):555–558, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335735.

Finkelstein:1971:IGF

[FS71] J. M. Finkelstein and R. E. Schafer. Improved goodness-of-fit tests. Biometrika, 58(3):641–645, December 1971. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334400.

Fujino:1979:THS

[Fuj79] Yoritake Fujino. Tests for the homogeneity of a set of variances against ordered alternatives. Biometrika, 66(1):133–139, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335252.

Fushimi:1973:IVS

[Fus73] Masanori Fushimi. An improved version of a Sobel-Weiss playthe-winner procedure for selecting the better of two binomial populations. Biometrika, 60(3):517–523, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335000.

Fisher:1971:ACP

[FV71] Lloyd Fisher and John W. Van Ness. Admissible clustering procedures. Biometrika, 58(1):91–104, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334320.

Fujikoshi:1979:EDC

[FV79] Y. Fujikoshi and L. G. Veitch. Estimation of dimensionality in canonical correlation analysis. Biometrika, 66(2):345–351, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335669.

Gabriel:1971:BGD

[Gab71] K. R. Gabriel. The biplot graphic display of matrices with application to principal component analysis. Biometrika, 58(3):453–467, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334381.

Gail:1974:VSC

[Gai74] Mitchell Gail. Value systems for comparing two independent multinomial trials. Biometrika, 61(1):91–100, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334289.

Gart:1970:PIE

[Gar70] John J. Gart. Point and interval estimation of the common odds ratio in the combination of 2 × 2 tables with fixed marginals.

Biometrika, 57(3):471–475, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334765.

Gardner:1972:UER

[Gar72a] M. J. Gardner. On using an estimated regression line in a second sample. Biometrika, 59(2):263–274, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334571.

Gart:1972:ITC

[Gar72b] John J. Gart. Interaction tests for 2 × s × t contingency tables. Biometrika, 59(2):309–316, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2334575. See corrections and amendments [Gar73].

Gart:1973:CAI

[Gar73] John J. Gart. Corrections and amendments: “Interaction Tests for 2×s×t Contingency Tables”. Biometrika, 60(2):435, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334565. See [Gar72b].

Gasser:1975:GFT

[Gas75] Theo Gasser. Goodness-of-fit tests for correlated data. Biometrika, 62(3):563–570, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335511.

Gunel:1974:BFI

[GD74] Erdogan Gûnel and James Dickey. Bayes factors for independence in contingency tables. Biometrika, 61(3):545–557, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334738.

Geary:1936:MRM

[Gea36] R. C. Geary. Moments of the ratio of the mean deviation to the standard deviation for normal samples. Biometrika, 28(3/4): 295–307, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333953. See comments [SBL79].

Geary:1970:REC

[Gea70] R. C. Geary. Relative efficiency of count of sign changes for assessing residual autoregression in least squares regression. Biometrika, 57(1):123–127, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334942.

Gebhardt:1970:AMR

[Geb70] Friedrch Gebhardt. An algorithm for monotone regression with one or more independent variables. Biometrika, 57(2):263–271, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334835.

Geisser:1974:PAR

[Gei74] Seymour Geisser. A predictive approach to the random effect model. Biometrika, 61(1):101–107, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334290.

George:1971:EEB

[Geo71] Stephen L. George. Evaluation of empirical Bayes estimators for small numbers of past samples. Biometrika, 58(1):244, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334342.

Good:1971:NRP

[GG71] I. J. Good and R. A. Gaskins. Nonparametric roughness penalties for probability densities. Biometrika, 58(2):255–277, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334515.

Gilula:1979:SVD

[Gil79] Zvi Gilula. Singular value decomposition of probability matrices: Probabilistic aspects of latent dichotomous variables. Biometrika, 66(2):339–344, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335668.

Gill:1979:BES

[GJ79a] Christine A. Gill and Derrick N. Joanes. Bayesian estimation of Shannon’s index of diversity. Biometrika, 66(1):81–85, April 1979.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335245.

Gittins:1979:DAI

[GJ79b] J. C. Gittins and D. M. Jones. A dynamic allocation index for the discounted multiarmed bandit problem. Biometrika, 66(3): 561–565, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335176.

Gabbert:1970:SSN

[GK70] James T. Gabbert and Richard G. Krutchkoff. Supplementary sample nonparametric empirical Bayes in a quality control situation. Biometrika, 57(1):67–69, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334936.

Griffin:1971:OLE

[GK71] Barry S. Griffin and Richard G. Krutchkoff. Optimal linear estimators: An empirical Bayes version with application to the binomial distribution. Biometrika, 58(1):195–201, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334331.

Gnanadesikan:1970:GTI

[GL70] R. Gnanadesikan and Elisa T. Lee. Graphical techniques for internal comparisons amongst equal degree of freedom groupings in multirespnse experiments. Biometrika, 57(2):229–237, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334832.

Glazebrook:1978:OAT

[Gla78] K. D. Glazebrook. On the optimal allocation of two or more treatments in a controlled clinical trial. Biometrika, 65(2):335–340, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335212.

Good:1979:PER

[GLGH79] I. J. Good, Byron C. Lewis, Raymond A. Gaskins, and L. W. Howell. Population estimation by the removal method assuming proportional trapping. Biometrika, 66(3):485–494, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335167.

Ganesalingam:1978:ELD

[GM78] S. Ganesalingam and G. J. McLachlan. The efficiency of a linear discriminant function based on unclassified initial samples. Biometrika, 65(3):658–662, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335925.

Gupta:1973:OSE

[GNP73] Shanti S. Gupta, Klaus Nagel, and S. Panchapakesan. On the order statistics from equally correlated normal random variables. Biometrika, 60(2):403–413, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334554.

Gastwirth:1977:CTN

[GO77] Joseph L. Gastwirth and M. E. B. Owens. On classical tests of normality. Biometrika, 64(1):135–139, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335783.

Godolphin:1975:DBF

[God75] E. J. Godolphin. A direct basic form for predictors of autoregressive integrated moving average processes. Biometrika, 62(2):483–496, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335391.

Godambe:1976:CLU

[God76a] V. P. Godambe. Conditional likelihood and unconditional optimum estimating equations. Biometrika, 63(2):277–284, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335620.

Godolphin:1976:CWF

[God76b] E. J. Godolphin. On the Cramér-Wold factorization. Biometrika, 63(2):367–380, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335632.

Godolphin:1977:DRM

[God77] E. J. Godolphin. A direct representation for the maximum likelihood estimator of a Gaussian moving average process. Biometrika, 64(2):375–384, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335706.

Godolphin:1978:MML

[God78] E. J. Godolphin. Modified maximum likelihood estimation of Gaussian moving averages using a pseudoquadratic convergence criterion. Biometrika, 65(1):203–206, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335295.

Godfrey:1979:TAT

[God79] L. G. Godfrey. Testing the adequacy of a time series model. Biometrika, 66(1):67–72, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335243.

Goldstein:1974:ABI

[Gol74] M. Goldstein. Approximate Bayesian inference with incompletely specified prior distributions. Biometrika, 61(3):629–631, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334750.

Goldstein:1976:BAR

[Gol76] M. Goldstein. Bayesian analysis of regression problems. Biometrika, 63(1):51–58, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335083.

Good:1969:CQF

[Goo69] I. J. Good. Conditions for a quadratic form to have a chi-squared distribution. Biometrika, 56(1):215–216, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334707. See correction [Goo70]

and remark [Kha78].

Good:1970:CCQ

[Goo70] I. J. Good. Corrections: “Conditions for a Quadratic Form to Have a Chi-Squared Distribution”. Biometrika, 57(1):225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334962. See [Goo69].

Goodman:1973:AMC

[Goo73] Leo A. Goodman. The analysis of multidimensional contingency tables when some variables are posterior to others: A modified path analysis approach. Biometrika, 60(1):179–192, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334920.

Goodman:1974:ELS

[Goo74] Leo A. Goodman. Exploratory latent structure analysis using both identifiable and unidentifiable models. Biometrika, 61(2):215–231, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334349.

Good:1979:SHP

[Goo79a] I. J. Good. Studies in the history of probability and statistics. XXXVII. A. M. Turing’s statistical work in World War II. Biometrika, 66(2):393–396, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335677.

Goodman:1979:MMS

[Goo79b] Leo A. Goodman. Multiplicative models for square contingency tables with ordered categories. Biometrika, 66(3):413–418, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335159.

Gordon:1973:SCS

[Gor73] A. D. Gordon. A sequence-comparison statistic and algorithm. Biometrika, 60(1):197–200, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334922.

Gordon:1979:AMA

[Gor79a] A. D. Gordon. Another measure of the agreement between rankings. Biometrika, 66(2):327–332, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335666.

Gordon:1979:MAB

[Gor79b] A. D. Gordon. A measure of the agreement between rankings. Biometrika, 66(1):7–15, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335236.

Gough:1977:ELI

[Gou77] K. J. Gough. The estimation of latent and infectious periods. Biometrika, 64(3):559–565, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345333.

Gart:1970:CMS

[GP70] John J. Gart and Hugh M. Pettigrew. On the conditional moments of the k-statistics for the Poisson distribution. Biometrika, 57(3):661–664, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334785.

Gower:1975:CDC

[GP75] J. C. Gower and R. W. Payne. A comparison of different criteria for selecting binary tests in diagnostic keys. Biometrika, 62(3):665–672, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335526.

Gratton:1979:GTG

[Gra79] Richard J. Gratton. Generalizing the test graph method for estimating functions of an unknown density. Biometrika, 66(3): 663–667, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335190.

Green:1970:TD

[Gre70] J. R. Green. A test for dependence. Biometrika, 57(1):205–207, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334952.

Green:1979:MWT

[Gre79] J. R. Green. Modified Wilcoxon test for two different distributions. Biometrika, 66(3):645–653, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335187.

Gray:1975:GJR

[GSW75] H. L. Gray, W. R. Schucany, and T. A. Watkins. On the generalized jackknife and its relation to statistical differentials. Biometrika, 62(3):637–642, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335521.

Guenther:1972:NRP

[Gue72] William C. Guenther. On the number of replications of a paired comparison: An easy solution with standard tables. Biometrika, 59(2):481–483, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334597.

Guttman:1971:ROR

[Gut71] Irwin Guttman. A remark on the optimal regression designs with previous observations of Covey-Crump and Silvey. Biometrika, 58(3):683–685, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334409.

Gleser:1973:ELT

[GW73] Leon Jay Gleser and Geoffrey S. Watson. Estimation of a linear transformation. Biometrika, 60(3):525–534, December 1973. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335001.

Gail:1975:VSC

[GW75] Mitchell Gail and James Ware. Value systems for comparing three or more independent multinomial trials. Biometrika, 62(2):379– 382, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335376.

Gail:1978:RME

[GW78] Mitchell Gail and James Ware. On the robustness to measurement error of tests for exponentiality. Biometrika, 65(2):305–309, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335208.

Haigh:1971:NWP

[Hai71] John Haigh. A neat way to prove asymptotic normality. Biometrika, 58(3):677–678, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334406.

Hamdan:1970:ETM

[Ham70] M. A. Hamdan. The equivalence of tetrachoric and maximum likelihood estimates of ρ in 2 × 2 tables. Biometrika, 57(1):212–215, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334955.

Handa:1972:CIS

[Han72] B. R. Handa. Choice of initial states in estimating transition probabilities of a finite ergodic Markov chain. Biometrika, 59(2):407–414, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334585.

Han:1975:SRB

[Han75] Chien Pai Han. Some relationships between noncentral chi-squared and normal distributions. Biometrika, 62(1):213–214, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334506.

Hannan:1979:NAM

[Han79] E. J. Hannan. A note on autoregressive-moving average identification. Biometrika, 66(3):672–674, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335192.

Harter:1964:NTP

[Har64] H. Leon Harter. A new table of percentage points of the chi-square distribution. Biometrika, 51(1/2):231–239, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334209. See corrigenda [Har65,

Har79].

Harter:1965:CNT

[Har65] H. Leon Harter. Corrigenda: “A new table of percentage points of the chi-square distribution”. Biometrika, 52(1/2):305, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333844. See [Har64].

Harville:1969:QUE

[Har69] David A. Harville. Quadratic unbiased estimation of variance components for the one-way classification. Biometrika, 56(2):313–326, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334424.

See corrections [Har70].

Harville:1970:CQU

[Har70] David A. Harville. Corrections: “Quadratic Unbiased Estimation of Variance Components for the One-Way Classification”. Biometrika, 57(1):226, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334964. See [Har69].

Harville:1974:BIV

[Har74] David A. Harville. Bayesian inference for variance components using only error contrasts. Biometrika, 61(2):383–385, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334370.

Harter:1979:ACN

[Har79] H. L. Harter. Amendments and corrections: “A new table of percentage points of the chi-square distribution” [Biometrika 51 (1964), 231–239; MR 30 #4315]. Biometrika, 66(3):689, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335199. See

[Har64].

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hastings:1970:MCS

[Has70] W. K. Hastings. Monte Carlo sampling methods using Markov chains and their applications. Biometrika, 57(1):97–109, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334940;
http://www.probability.ca/hastings/. This paper introduces

what is now known as the Metropolis–Hastings algorithm, a generalization of the work in [MRR^{+}53]. See [Hit03, page 255].

Hassanein:1971:PEP

[Has71] Khatab M. Hassanein. Percentile estimators for the parameters of the Weibull distribution. Biometrika, 58(3):673–676, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334405.

Hastings:1974:VRN

[Has74] W. Keith Hastings. Variance reduction and non-normality. Biometrika, 61(1):143–149, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334296.

Hawkes:1971:SSS

[Haw71] Alan G. Hawkes. Spectra of some self-exciting and mutually exciting point processes. Biometrika, 58(1):83–90, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334319.

Hayakawa:1975:LRC

[Hay75] Takesi Hayakawa. The likelihood ratio criterion for a composite hypothesis under a local alternative. Biometrika, 62(2):451–460, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335387.

Hayre:1979:TPS

[Hay79] L. S. Hayre. Two-population sequential tests with three hypotheses. Biometrika, 66(3):465–474, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335165.

Haas:1970:ICD

[HBA70] Gerald Haas, Lee Bain, and Charles Antle. Inference for the Cauchy distribution based on maximum likelihood estimators. Biometrika, 57(2):403–408, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334849.

Herzberg:1972:SOD

[HC72] Agnes M. Herzberg and D. R. Cox. Some optimal designs for interpolation and extrapolation. Biometrika, 59(3):551–561, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334806.

Heathcote:1977:ISE

[Hea77] C. R. Heathcote. The integrated squared error estimation of parameters. Biometrika, 64(2):255–264, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335691.

Healy:1978:MDE

[Hea78] M. J. R. Healy. A mean difference estimator of standard deviation in symmetrically censored normal samples. Biometrika, 65 (3):643–646, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335919.

Healy:1979:CPP

[Hea79] M. J. R. Healy. Comments on paper by P. D. Finch. Biometrika, 66(2):206, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335648. See [Fin79a].

Heilbron:1978:CEV

[Hei78] David C. Heilbron. Comparison of estimators of the variance of systematic sampling. Biometrika, 65(2):429–433, August 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335226.

Hemon:1977:OUE

[Hem77] Denis Hemon. Optimal use of extraneous information concerning a nuisance parameter. Biometrika, 64(1):51–57, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335770.

Herr:1976:GCC

[Her76] David G. Herr. A geometric characterization of connectedness in a two-way design. Biometrika, 63(1):93–100, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335088.

Heyde:1975:REE

[Hey75] C. C. Heyde. Remarks on efficiency in estimation for branching processes. Biometrika, 62(1):49–55, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334486.

Healy:1976:ASC

[HG76] M. J. R. Healy and H. Goldstein. An approach to the scaling of categorized attributes. Biometrika, 63(2):219–229, August 1976.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335613.

Hinkley:1970:IACb

[HH70] David V. Hinkley and Elizabeth A. Hinkley. Inference about the change-point in a sequence of binomial variables. Biometrika, 57 (3):477–488, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334766.

Hahn:1971:TPP

[HH71a] Gerald J. Hahn and Richard W. Hendrickson. A table of percentage points of the distribution of the largest absolute value of k Student t variates and its applications. Biometrika, 58(2):323–332, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334520.

Hinkley:1971:CAI

[HH71b] David V. Hinkley and Elizabeth A. Hinkley. Correction and amendments: “Inference about the Change-Point in a Sequence of Binomial Variables”. Biometrika, 58(3):687, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334411. See [Hin70b].

Hines:1979:ESD

[HH79] W. G. S. Hines and R. J. O’Hara Hines. The Eberhardt statistic and the detection of nonrandomness of spatial point distributions. Biometrika, 66(1):73–79, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335244.

Habbema:1974:CDA

[HHvdB74] J. D. F. Habbema, J. Hermans, and A. T. van der Burgt. Cases of doubt in allocation problems. Biometrika, 61(2):313–324, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334360.

Hill:1976:NAM

[Hil76] G. W. Hill. New approximations to the von Mises distribution. Biometrika, 63(3):673–676, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335751. See correction [Hil77].

Hill:1977:CAN

[Hil77] G. W. Hill. Corrections and amendments: “New Approximations to the von Mises Distribution”. Biometrika, 64(3):655, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345356. See

[Hil76].

Hill:1978:NEO

[Hil78] Peter D. H. Hill. A note on the equivalence of D-optimal design measures for three rival linear models. Biometrika, 65(3):666–667, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335928.

Hinkley:1969:RTC

[Hin69] D. V. Hinkley. On the ratio of two correlated normal random variables. Biometrika, 56(3):635–639, December 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334671. See correction

[Hin70a].

Hinkley:1970:CRT

[Hin70a] D. V. Hinkley. Correction: “On the Ratio of Two Correlated Normal Random Variables”. Biometrika, 57(3):683, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334796. See [Hin69].

Hinkley:1970:IACa

[Hin70b] David V. Hinkley. Inference about the change-point in a sequence of random variables. Biometrika, 57(1):1–17, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334932. See correction

[HH71b].

Hinkley:1971:IAC

[Hin71] D. V. Hinkley. Inference about the change-point from cumulative sum tests. Biometrika, 58(3):509–523, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334386.

Hinkley:1972:TOC

[Hin72] D. V. Hinkley. Time-ordered classification. Biometrika, 59(3): 509–523, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334802.

Hinkley:1975:PTS

[Hin75] David V. Hinkley. On power transformations to symmetry. Biometrika, 62(1):101–111, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334491. See correction [Hin76a].

Hinkley:1976:CAP

[Hin76a] David Hinkley. Corrections and amendments: “On power transformations to symmetry” (Biometrika 62 (1975), 101–111). Biometrika, 63(1):218, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335113. See [Hin75].

Hinkley:1976:ESD

[Hin76b] David Hinkley. On estimating a symmetric distribution. Biometrika, 63(3):680–681, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335754.

Hinkley:1977:CIA

[Hin77a] D. V. Hinkley. Conditional inference about a normal mean with known coefficient of variation. Biometrika, 64(1):105–108, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335778. See cor-

rections [Hin78b, Hin78c].

Hinkley:1977:JCL

[Hin77b] D. V. Hinkley. Jackknife confidence limits using Student t approximations. Biometrika, 64(1):21–28, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2335765. See amendments and corrections [Hin78a].

Hinkley:1978:ACJ

[Hin78a] D. V. Hinkley. Amendments and corrections: Jackknife Confidence Limits Using Student t Approximations. Biometrika, 65(3):677, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335933.

See [Hin77b].

Hinkley:1978:ACC

[Hin78b] David V. Hinkley. Amendments and corrections: “Conditional Inference About a Normal Mean with Known Coefficient of Variation”. Biometrika, 65(3):677, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335934. See [Hin77a].

Hinkley:1978:CCIc

[Hin78c] David V. Hinkley. Correction to: “Conditional inference about a normal mean with known coefficient of variation” (Biometrika 64 (1977), no. 1, 105–108). Biometrika, 65(3):677, 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). See [Hin77a].

Hinkley:1978:IJS

[Hin78d] David V. Hinkley. Improving the jackknife with special reference to correlation estimation. Biometrika, 65(1):13–21, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335271.

Hinkley:1978:LIA

[Hin78e] David V. Hinkley. Likelihood inference about location and scale parameters. Biometrika, 65(2):253–261, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335203.

Hinkley:1979:LIS

[Hin79] David V. Hinkley. Likelihood inference for a simple growth model. Biometrika, 66(3):659–662, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335189.

Holcomb:1972:CRP

[HIR72] D. F. Holcomb, M. Iwasawa, and F. D. K. Roberts. Clustering of randomly placed spheres. Biometrika, 59(1):207–209, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334632.

Hirotsu:1978:OAI

[Hir78] C. Hirotsu. Ordered alternatives for interaction effects. Biometrika, 65(3):561–570, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335908.

Hirotsu:1979:AA

[Hir79] C. Hirotsu. An F approximation and its application. Biometrika, 66(3):577–584, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335178.

Hitchcock:2003:HMH

[Hit03] David B. Hitchcock. A history of the Metropolis–Hastings algorithm. The American Statistician, 57(4):254–257, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s7/p254;
http://www.jstor.org/stable/30037292.

Hoel:1979:ODD

[HJ79] P. G. Hoel and R. I. Jennrich. Optimal designs for dose response experiments in cancer research. Biometrika, 66(2):307–316, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335663.

Howie:1972:SPI

[HM72] A. J. Howie and A. B. Macaulay. Some problems of inference for a two-state system with incomplete information. Biometrika, 59(2):395–406, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334584.

Harrison:1975:AHN

[HM75a] M. J. Harrison and B. P. M. McCabe. Autocorrelation with heteroscedasticity: A note on the robustness of the Durbin–Watson, Geary and Henshaw tests. Biometrika, 62(1):214–216, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334507.

Hettmansperger:1975:MMT

[HM75b] Thomas P. Hettmansperger and Janice S. Malin. A modified Mood’s test for location with no shape assumptions on the underlying distributions. Biometrika, 62(2):527–529, August 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335398.

Hart:1976:WMG

[HM76] M. L. Hart and A. H. Money. On Wilks’s multivariate generalization of the correlation ratio. Biometrika, 63(1):59–67, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335084.

Hensler:1977:NSE

[HMM77] G. L. Hensler, K. G. Mehrotra, and J. E. Michalek. A note on some exact efficiency calculations for survival distributions. Biometrika, 64(3):635–637, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345347.

Holgate:1970:MSC

[Hol70] P. Holgate. The modality of some compound Poisson distributions. Biometrika, 57(3):666–667, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334787.

Hollander:1971:NTB

[Hol71a] Myles Hollander. A nonparametric test for bivariate symmetry. Biometrika, 58(1):203–212, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334332.

Holst:1971:NFS

[Hol71b] Lars Holst. A note on finding the size of a finite population. Biometrika, 58(1):228–229, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334337.

Holst:1972:CAA

[Hol72a] L. Holst. Corrections and amendments: “Asymptotic Normality and Efficiency of Certain Goodness-of-Fit Tests”. Biometrika, 59(3):699, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334828. See [Hol72b].

Holst:1972:ANE

[Hol72b] Lars Holst. Asymptotic normality and efficiency for certain goodness-of-fit tests. Biometrika, 59(1):137–145, April 1972. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334624. See correction

[Hol72a].

Holgate:1973:ESA

[Hol73] P. Holgate. An estimator for the size of an animal population. Biometrika, 60(1):135–140, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334915.

Holt:1978:CRA

[Hol78] J. D. Holt. Competing risk analyses with special reference to matched pair experiments. Biometrika, 65(1):159–165, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335291.

Hosking:1979:ADR

[Hos79] J. R. M. Hosking. The asymptotic distribution of R^{2} for autoregressive-moving average time series models when parameters are estimated. Biometrika, 66(1):156–157, April 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335256.

Hartley:1972:QFO

[HP72] H. O. Hartley and R. C. Pfaffenberger. Quadratic forms in order statistics used as goodness-of-fit criteria. Biometrika, 59(3): 605–611, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334811.

Holt:1974:SAT

[HP74] J. D. Holt and R. L. Prentice. Survival analyses in twin studies and matched pair experiments. Biometrika, 61(1):17–30, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334281.

Hollander:1975:TMR

[HP75] Myles Hollander and Frank Proschan. Tests for the mean residual life. Biometrika, 62(3):585–593, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335514. See correction [HP76,

HP80].

Hollander:1976:CAT

[HP76] Myles Hollander and Frank Proschan. Corrections and amendments: “Tests for the mean residual life” (Biometrika 62 (1975), 585–593). Biometrika, 63(2):412, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335641. See [HP75].

Harvey:1979:MLE

[HP79] A. C. Harvey and G. D. A. Phillips. Maximum likelihood estimation of regression models with autoregressive- moving average disturbances. Biometrika, 66(1):49–58, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335241.

Hollander:1980:ACT

[HP80] M. Hollander and F. Proschan. Amendments and corrections: “Tests for the mean residual life” [Biometrika 62 (1975), no. 3, 585–593; MR 52 #15917]. Biometrika, 67(1):259, April 1980. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335351. See [HP75].

Hedayat:1970:ECT

[HPF70] A. Hedayat, E. T. Parker, and W. T. Federer. The existence and construction of two families of designs for two successive experiments. Biometrika, 57(2):351–355, August 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334843. See correction

[HPF71].

Hedayat:1971:CAE

[HPF71] A. Hedayat, E. T. Parker, and W. T. Federer. Correction and amendments: “The Existence and Construction of Two Families of Designs for Two Successive Experiments”. Biometrika, 58(3):687, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334412.

See [HPF70].

Hill:1977:TSB

[HR77] D. L. Hill and P. V. Rao. Tests of symmetry based on Cramérvon Mises statistics. Biometrika, 64(3):489–494, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345324.

Heyde:1972:SHP

[HS72a] C. C. Heyde and E. Seneta. Studies in the history of probability and statistics. XXXI. The simple branching process, a turning point test and a fundamental inequality: a historical note on I. J. Bienaymé. Biometrika, 59(3):680–683, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334819.

Hughes:1972:APP

[HS72b] David T. Hughes and John G. Saw. Approximating the percentage points of Hotelling’s generalized T_{0}^{2} statistic. Biometrika, 59 (1):224–226, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334640.

Hubert:1975:MLP

[HS75] Lawrence Hubert and James Schultz. Maximum likelihood pairedcomparison ranking and quadratic assignment. Biometrika, 62 (3):655–659, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335524.

Hirji:1978:THT

[HS78a] Taj Hirji and A. K. Shahani. Testing of hypotheses with trinomials generated by gauges. Biometrika, 65(1):228–231, April 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335302.

Hollander:1978:RC

[HS78b] M. Hollander and J. Sethuraman. Reply to comments. Biometrika, 65(2):411, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335223. See [HS78c].

Hollander:1978:TAB

[HS78c] Myles Hollander and Jayaram Sethuraman. Testing for agreement between two groups of judges. Biometrika, 65(2):403–410, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335221. See

comments [Sch78b] and reply [HS78b].

Harrell:1979:SIC

[HS79] Frank E. Harrell and Pranab K. Sen. Statistical inference for censored bivariate normal distributions based on induced order statistics. Biometrika, 66(2):293–298, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335661.

Hoel:1972:TSP

[HSW72] David G. Hoel, Milton Sobel, and George H. Weiss. A twostage procedure for choosing the better of two binomial populations. Biometrika, 59(2):317–322, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334576.

Hannan:1971:ECG

[HT71] E. J. Hannan and P. J. Thomson. The estimation of coherence and group delay. Biometrika, 58(3):469–481, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334382.

Hannan:1973:EGD

[HT73] E. J. Hannan and P. J. Thomson. Estimating group delay. Biometrika, 60(2):241–253, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334536.

Hutchison:1971:MCC

[Hut71] M. C. Hutchison. A Monte Carlo comparison of some ratio estimators. Biometrika, 58(2):313–321, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334519.

Herzberg:1972:BRB

[HW72] Agnes M. Herzberg and H. P. Wynn. Book review: Theory of Optimal Experiments, by V. V. Fedorov; W. J. Studden; E. M. Klimko. Biometrika, 59(3):697–698, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334826.

Hall:1973:CSD

[HW73] W. B. Hall and E. R. Williams. Cyclic superimposed designs. Biometrika, 60(1):47–53, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334904.

Hamilton:1978:IPA

[HW78] David C. Hamilton and Donald G. Watts. Interpreting partial autocorrelation functions of seasonal time series models. Biometrika, 65(1):135–140, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335288.

Hwang:1976:GTD

[Hwa76] F. K. Hwang. Group testing with a dilution effect. Biometrika, 63(3):671–673, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335750.

Halperin:1977:TIC

[HWB^{+}77] Max Halperin, James H. Ware, David P. Byar, Nathan Mantel, Charles C. Brown, James Koziol, Mitchell Gail, and Sylvan B. Green. Testing for interaction in an I × J × K contingency table. Biometrika, 64(2):271–275, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335693.

Hyde:1977:TSU

[Hyd77] John Hyde. Testing survival under right censoring and left truncation. Biometrika, 64(2):225–230, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335688.

Izenman:1976:EQG

[Ize76] Alan Julian Izenman. On the extremal quotient from a gamma sample. Biometrika, 63(1):185–190, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335101.

James:1978:CPB

[Jam78] A. T. James. Comments on paper by B. Efron and D. V. Hinkley. Biometrika, 65(3):484–485, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335895. See [EH78b].

Janacek:1975:EMM

[Jan75] G. Janacek. Estimation of the minimum mean square error of prediction. Biometrika, 62(1):175–180, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334501.

Jarrett:1977:BEF

[Jar77] Richard G. Jarrett. Bounds for the efficiency factor of block designs. Biometrika, 64(1):67–72, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335772.

Jarrett:1979:NIB

[Jar79] R. G. Jarrett. A note on the intervals between coal-mining disasters. Biometrika, 66(1):191–193, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335266.

Jensen:1979:LMM

[Jen79] D. R. Jensen. Linear models without moments. Biometrika, 66 (3):611–617, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335183.

Jarrett:1978:GCI

[JH78] R. G. Jarrett and W. B. Hall. Generalized cyclic incomplete block designs. Biometrika, 65(2):397–401, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335220.

Johnson:1971:SNT

[JK71a] N. L. Johnson and J. O. Kitchen. Some notes on tables to facilitate fitting S_{B} curves. Biometrika, 58(1):223–226, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334335. See extensions and cor-

rections [Joh74b].

Johnson:1971:TFF

[JK71b] N. L. Johnson and J. O. Kitchen. Tables to facilitate fitting S_{B} curves II: Both terminals known. Biometrika, 58(3):657–668, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334403.

Johnson:1972:PTG

[JK72] N. L. Johnson and S. Kotz. Power transformations of gamma variables. Biometrika, 59(1):226–229, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334641.

Johnson:1973:EMT

[JK73] N. L. Johnson and S. Kotz. Extended and multivariate Tukey Lambda distributions. Biometrika, 60(3):655–661, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335017.

Jones:1978:DCD

[JM78] E. R. Jones and T. J. Mitchell. Design criteria for detecting model inadequacy. Biometrika, 65(3):541–551, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335906.

Jensen:1975:ODL

[JMM75] D. R. Jensen, L. S. Mayer, and R. H. Myers. Optimal designs and large-sample tests for linear hypotheses. Biometrika, 62(1):71–78, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334488.

Joanes:1972:STC

[Joa72] D. N. Joanes. Sequential tests of composite hypotheses. Biometrika, 59(3):633–637, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2334814. See correction [Joa75].

Joanes:1975:CAS

[Joa75] D. N. Joanes. Corrections and amendments: “Sequential tests of composite hypotheses” (Biometrika 59 (1972), 633–637). Biometrika, 62(1):221, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334510. See [Joa72].

Johnson:1970:NIS

[Joh70] Mark A. Johnson. A note on the interpretation of Springer’s simulation studies of the Kiefer–Wolfowitz process. Biometrika, 57 (3):673–674, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334791.

John:1971:SOM

[Joh71] S. John. Some optimal multivariate tests. Biometrika, 58(1):123– 127, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334322.

John:1972:DSU

[Joh72] S. John. The distribution of a statistic used for testing sphericity of normal distributions. Biometrika, 59(1):169–173, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334628.

John:1973:GCD

[Joh73] J. A. John. Generalized cyclic designs in factorial experiments. Biometrika, 60(1):55–63, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334905.

John:1974:ASI

[Joh74a] S. John. Acceptability and statistical inference. Biometrika, 61 (2):285–290, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334356.

Johnson:1974:ECT

[Joh74b] N. L. Johnson. Extensions and corrections to: “Tables to facilitate fitting S_{U} frequency curves” (Biometrika 52 (1965), 547– 558). Biometrika, 61(1):203–205, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334309. See [JK71a].

Johnson:1979:NPK

[Joh79] N. S. Johnson. Nonnull properties of Kendall’s partial rank correlation coefficient. Biometrika, 66(2):333–337, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335667.

Jones:1974:JEF

[Jon74a] Howard L. Jones. Jackknife estimation of functions of stratum means. Biometrika, 61(2):343–348, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334363.

Jones:1974:NBS

[Jon74b] P. W. Jones. A note on the Bayesian sequential estimation of a binomial parameter. Biometrika, 61(3):642–644, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334755.

Jones:1975:TAB

[Jon75] P. W. Jones. The two-armed bandit. Biometrika, 62(2):523–524, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335396.

Joreskog:1970:GMA

[Jör70] K. G. Jöreskog. A general method for analysis of covariance structures. Biometrika, 57(2):239–251, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334833.

Joshi:1972:SST

[Jos72] Prakash C. Joshi. Some slippage tests of mean for a single outlier in linear regression. Biometrika, 59(1):109–120, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334621.

Joshi:1973:TII

[Jos73] Prakash C. Joshi. Two identities involving order statistics. Biometrika, 60(2):428–429, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334560.

Jackson:1970:MEE

[JOZD70] D. A. Jackson, T. M. O’Donovan, W. J. Zimmer, and J. J. Deely. G_{∈}-minimax estimators in the exponential family. Biometrika, 57 (2):439–443, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334853.

Johnson:1969:TPP

[JP69] N. L. Johnson and E. S. Pearson. Tables of percentage points of non-central χ. Biometrika, 56(2):255–272, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334419. See correction [JP70].

Johnson:1970:CTP

[JP70] N. L. Johnson and E. S. Pearson. Correction: “Tables of Percentage Points of Noncentral χ”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334861. See [JP69].

Juola:1970:PSR

[Juo70] R. C. Juola. Population symmetry and the regions X < t^{2}/X. Biometrika, 57(3):668–669, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334788.

James:1971:FRO

[JW71] A. T. James and G. N. Wilkinson. Factorization of the residual operator and canonical decomposition of nonorthogonal factors in the analysis of variance. Biometrika, 58(2):279–294, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334516.

Jones:1979:SFL

[JW79] David Jones and John Whitehead. Sequential forms of the log rank and modified Wilcoxon tests for censored data. Biometrika, 66(1):105–113, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335249. See correction [JW81].

Jones:1981:ACS

[JW81] David Jones and John Whitehead. Amendments and corrections: “Sequential forms of the log rank and modified Wilcoxon tests for censored data” [Biometrika 66 (1979), no. 1, 105–113; MR 80i:62069]. Biometrika, 68(2):576, August 1981. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335610. See [JW79].

Kalbfleisch:1974:SEC

[Kal74] John D. Kalbfleisch. Some efficiency calculations for survival distributions. Biometrika, 61(1):31–38, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334282.

Kalbfleisch:1975:RC

[Kal75a] J. D. Kalbfleisch. Reply to comments. Biometrika, 62(2):268, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335361. See

[Kal75b].

Kalbfleisch:1975:SC

[Kal75b] John D. Kalbfleisch. Sufficiency and conditionality. Biometrika, 62(2):251–259, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2335356. See [Bar75b, BN75, Bir75, McL75] and reply [Kal75a].

Kappenman:1970:CPE

[Kap70] Russell F. Kappenman. On a comparison of the proportions exceeding a fixed value in two different normal populations. Biometrika, 57(3):669–670, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334789.

Kay:1979:SFA

[Kay79] Richard Kay. Some further asymptotic efficiency calculations for survival data regression models. Biometrika, 66(1):91–96, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335247.

Kerridge:1976:YAS

[KC76] D. F. Kerridge and G. W. Cook. Yet another series for the normal integral. Biometrika, 63(2):401–403, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335636.

Kedem:1976:SSA

[Ked76] Benjamin Kedem. Sufficient statistics associated with a two-state second-order Markov chain. Biometrika, 63(1):127–132, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335092.

Kedem:1978:SNL

[Ked78] Benjamin Kedem. Sufficiency and the number of level crossings by a stationary process. Biometrika, 65(1):207–210, April 1978.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335296.

Keiding:1974:CAE

[Kei74a] Niels Keiding. Corrections and amendments: “Estimation in the Birth Process”. Biometrika, 61(3):647, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334759. See [Kei74b].

Keiding:1974:EBP

[Kei74b] Niels Keiding. Estimation in the birth process. Biometrika, 61(1): 71–80, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2334286. See correction [Kei74a].

Kempton:1975:GFF

[Kem75] R. A. Kempton. A generalized form of Fisher’s logarithmic series. Biometrika, 62(1):29–38, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334484.

Kempthorne:1979:CPP

[Kem79] Oscar Kempthorne. Comments on paper by P. D. Finch. Biometrika, 66(2):206–207, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335649. See [Fin79a].

Kendall:1971:SHP

[Ken71] M. G. Kendall. Studies in the history of probability and statistics. XXVI. The work of Ernst Abbe. Biometrika, 58(2):369–373, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334525. See

corrections [Ken72].

Kendall:1972:CAS

[Ken72] M. G. Kendall. Corrections and amendments: “Studies in the history of probability and statistics. XXVI. The work of Ernst Abbe” (Biometrika 58 (1971), 369–373). Biometrika, 59(2):498, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334604.

See [Ken71].

Kettenring:1971:CAS

[Ket71] J. R. Kettenring. Canonical analysis of several sets of variables. Biometrika, 58(3):433–451, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334380.

Koziol:1976:CMS

[KG76] James A. Koziol and Sylvan B. Green. A Cramér-von Mises statistic for randomly censored data. Biometrika, 63(3):465–474, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335723.

Khatri:1978:RNS

[Kha78] C. G. Khatri. A remark on the necessary and sufficient conditions for a quadratic form to be distributed as chi-squared. Biometrika, 65(1):239–240, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335305. See [Goo69].

Kastenbaum:1970:SSRa

[KHB70a] Marvin A. Kastenbaum, David G. Hoel, and K. O. Bowman. Sample size requirements: One-way analysis of variance. Biometrika, 57(2):421–430, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334851.

Kastenbaum:1970:SSRb

[KHB70b] Marvin A. Kastenbaum, David G. Hoel, and K. O. Bowman. Sample size requirements: Randomized block designs. Biometrika, 57(3):573–577, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334775.

Kiefer:1975:ODV

[Kie75] J. Kiefer. Optimal design: Variation in structure and performance under change of criterion. Biometrika, 62(2):277–288, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335363.

Kingman:1972:RSS

[Kin72] J. F. C. Kingman. On random sequences with spherical symmetry. Biometrika, 59(2):492–494, August 1972. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334601.

Krishnaiah:1976:DLR

[KLC76] P. R. Krishnaiah, Jack C. Lee, and T. C. Chang. The distributions of the likelihood ratio statistics for tests of certain covariance structures of complex multivariate normal populations. Biometrika, 63(3):543–549, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335733.

Kupper:1973:NAM

[KM73] Lawrence L. Kupper and Edward F. Meydrech. A new approach to mean squared error estimation of response surfaces. Biometrika, 60(3):573–579, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335007.

Kalbfleisch:1977:ESD

[KM77] J. D. Kalbfleisch and A. A. McIntosh. Efficiency in survival distributions with time-dependent covariables. Biometrika, 64(1):47–50, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335769.

Kalbfleisch:1979:CSM

[KM79] J. D. Kalbfleisch and R. J. MacKay. On constant-sum models for censored survival data. Biometrika, 66(1):87–90, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335246.

Kaminsky:1975:BSL

[KMN75] Kenneth S. Kaminsky, Nancy R. Mann, and Paul I. Nelson. Best and simplified linear invariant prediction of order statistics in location and scale families. Biometrika, 62(2):525–527, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335397.

Knoke:1975:TRA

[Kno75a] James D. Knoke. Testing for randomness against autocorrelated alternatives: The parametric case. Biometrika, 62(3):571–575, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335512.

Knott:1975:MEL

[Kno75b] M. Knott. On the minimum efficiency of least squares. Biometrika, 62(1):129–132, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334494.

Knoke:1977:TRA

[Kno77] James D. Knoke. Testing for randomness against autocorrelation: Alternative tests. Biometrika, 64(3):523–529, December 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345328.

Kochar:1979:DFC

[Koc79] Subhash C. Kochar. Distribution-free comparison of two probability distributions with reference to their hazard rates. Biometrika, 66(3):437–441, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335162.

Koehler:1979:GFM

[Koe79] Kenneth Koehler. A general formula for moments of the Pearson goodness-of-fit statistic for alternatives. Biometrika, 66(2):397– 399, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335678.

Kohn:1977:NCA

[Koh77] R. Kohn. Note concerning the Akaike and Hannan estimation procedures for an autoregressive-moving average process. Biometrika, 64(3):622–625, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345343.

Konishi:1978:ADS

[Kon78] Sadanori Konishi. An approximation to the distribution of the sample correlation coefficient. Biometrika, 65(3):654–656, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335923.

Korin:1972:SCH

[Kor72] Basil P. Korin. Some comments on the homoscedasticity criterion M and the multivariate analysis of variance tests T^{2}, W and R. Biometrika, 59(1):215–216, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334636.

Kalbfleisch:1973:MLB

[KP73] J. D. Kalbfleisch and R. L. Prentice. Marginal likelihoods based on Cox’s regression and life model. Biometrika, 60(2):267–278, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334538.

Kok:1976:ART

[KP76] Kim-Lian Kok and H. D. Patterson. Algebraic results in the theory of serial factorial design. Biometrika, 63(3):559–565, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335736.

Koziol:1978:STE

[KP78] James A. Koziol and A. John Petkau. Sequential testing of the equality of two survival distributions using the modified Savage statistic. Biometrika, 65(3):615–623, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335914.

Kelly:1976:NSM

[KR76] F. P. Kelly and B. D. Ripley. A note on Strauss’s model for clustering. Biometrika, 63(2):357–360, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335630.

Kryscio:1972:EIR

[Kry72] Richard J. Kryscio. On estimating the infection rate of the simple stochastic epidemic. Biometrika, 59(1):213–214, April 1972. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334635.

Kryscio:1974:ESS

[Kry74] Richard J. Kryscio. On the extended simple stochastic epidemic model. Biometrika, 61(1):200–202, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334308.

Krzanowski:1979:SLT

[Krz79] W. J. Krzanowski. Some linear transformations for mixtures of binary and continuous variables, with particular reference to linear discriminant analysis. Biometrika, 66(1):33–39, April 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335239.

Kanofsky:1972:ACP

[KS72] Paul Kanofsky and R. Srinivasan. An approach to the construction of parametric confidence bands on cumulative distribution functions. Biometrika, 59(3):623–631, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334813.

Lachenbruch:1975:ZMD

[Lac75] Peter A. Lachenbruch. Zero-mean difference discrimination and the absolute linear discriminant function. Biometrika, 62(2):397–401, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335379.

Laird:1978:EBM

[Lai78] N. M. Laird. Empirical Bayes methods for two-way contingency tables. Biometrika, 65(3):581–590, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335910.

Latta:1977:GWS

[Lat77] Robert B. Latta. Generalized Wilcoxon statistics for the twosample problem with censored data. Biometrika, 64(3):633–635, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345346.

Lawless:1972:CCI

[Law72] J. F. Lawless. Conditional confidence interval procedures for the location and scale parameters of the Cauchy and logistic distributions. Biometrika, 59(2):377–386, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334582.

Lawless:1974:ACI

[Law74] J. F. Lawless. Approximations to confidence intervals for parameters in the extreme value and Weibull distributions. Biometrika, 61(1):123–129, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334293.

Layard:1974:MCC

[Lay74] M. W. J. Layard. A Monte Carlo comparison of tests for equality of covariance matrices. Biometrika, 61(3):461–465, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334727.

Laycock:1975:ODR

[Lay75] P. J. Laycock. Optimal design: Regression models for directions. Biometrika, 62(2):305–311, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335365.

Littell:1977:DFP

[LB77] Ramon C. Littell and James M. Boyett. Designs for R × C factorial paired comparison experiments. Biometrika, 64(1):73–77, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335773.

Ljung:1978:MLF

[LB78] G. M. Ljung and G. E. P. Box. On a measure of lack of fit in time series models. Biometrika, 65(2):297–303, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335207.

Ljung:1979:LFS

[LB79] Greta M. Ljung and George E. P. Box. The likelihood function of stationary autoregressive-moving average models. Biometrika, 66(2):265–270, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335657.

Lam:1976:PCS

[LC76] K. Lam and W. K. Chiu. On the probability of correctly selecting the best of several normal populations. Biometrika, 63(2):410–411, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335640.

Lee:1975:MCS

[LDG75] Elisa T. Lee, M. M. Desu, and E. A. Gehan. A Monte Carlo study of the power of some two-sample tests. Biometrika, 62(2):425–432, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335383.

Lee:1971:AFD

[Lee71] Yoong-Sin Lee. Asymptotic formulae for the distribution of a multivariate test statistic: Power comparisons of certain multivariate tests. Biometrika, 58(3):647–651, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334401.

Lee:1972:SRD

[Lee72a] Yoong Sin Lee. Some results on the distribution of Wilks’s likelihood-ratio criterion. Biometrika, 59(3):649–664, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334816.

Lee:1972:TUP

[Lee72b] Yoong Sin Lee. Tables of upper percentage points of the multiple correlation coefficient. Biometrika, 59(1):175–189, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334629.

Lee:1974:NRR

[Lee74] Yeong Hwi K. Lee. A note on Rao’s reduction of Potthoff & Roy’s generalized linear model. Biometrika, 61(2):349–351, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334364.

Lee:1979:CEC

[Lee79] Sik Yum Lee. Constrained estimation in covariance structure analysis. Biometrika, 66(3):539–545, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335173.

Leonard:1972:BMB

[Leo72] T. Leonard. Bayesian methods for binomial data. Biometrika, 59(3):581–589, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334809.

Leonard:1973:BMH

[Leo73] T. Leonard. A Bayesian method for histograms. Biometrika, 60 (2):297–308, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334541.

Leonard:1974:MBE

[Leo74] Tom Leonard. A modification to the Bayes estimate for the mean of a normal distribution. Biometrika, 61(3):627–628, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334749.

Leonard:1976:SAA

[Leo76] Tom Leonard. Some alternative approaches to multiparameter estimation. Biometrika, 63(1):69–75, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335085.

Lepage:1971:CWA

[Lep71] Yves Lepage. A combination of Wilcoxon’s and Ansari-Bradley’s statistics. Biometrika, 58(1):213–217, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334333.

Lepage:1973:TCW

[Lep73] Yves Lepage. A table for a combined Wilcoxon Ansari-Bradley statistic. Biometrika, 60(1):113–116, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334912.

Levenbach:1972:EAP

[Lev72] Hans Levenbach. Estimation of autoregressive parameters from a marginal likelihood function. Biometrika, 59(1):61–71, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334615.

Lewis:1972:STB

[Lew72] J. L. Lewis. A k-sample test based on range intervals. Biometrika, 59(1):155–160, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334626.

Lewis:1975:RED

[Lew75] S. M. Lewis. Robust estimation of density for a two-dimensional point process. Biometrika, 62(2):519–521, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335394.

Lininger:1979:CFT

[LGGB79] Lloyd Lininger, Mitchell H. Gail, Sylvan B. Green, and David P. Byar. Comparison of four tests for equality of survival curves in the presence of stratification and censoring. Biometrika, 66(3):419–428, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335160.

Lindgren:1974:SME

[Lin74] Georg Lindgren. Spectral moment estimation by means of level crossings. Biometrika, 61(3):401–418, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334721.

Lindley:1977:PFS

[Lin77] D. V. Lindley. A problem in forensic science. Biometrika, 64 (2):207–213, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335686.

Little:1976:IAM

[Lit76] R. J. A. Little. Inference about means from incomplete multivariate data. Biometrika, 63(3):593–604, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335740.

Lewis:1976:TME

[LJ76] S. M. Lewis and J. A. John. Testing main effects in fractions of asymmetrical factorial experiments. Biometrika, 63(3):678–680, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335753.

Lawley:1973:RFA

[LM73] D. N. Lawley and A. E. Maxwell. Regression and factor analysis. Biometrika, 60(2):331–338, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334544.

Lieberman:1967:USD

[LMH67] G. J. Lieberman, R. G. Miller, Jr., and M. A. Hamilton. Unlimited simultaneous discrimination intervals in regression. Biometrika, 54 (1/2):133–145, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333857. See correction [LMH71].

Lieberman:1971:CAU

[LMH71] G. J. Lieberman, R. G. Miller, Jr., and M. A. Hamilton. Corrections and amendments: “Unlimited Simultaneous Discrimination Intervals in Regression”. Biometrika, 58(3):687, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334413. See [LMH67].

Liddell:1978:LCC

[LO78] I. G. Liddell and J. K. Ord. Linear-circular correlation coefficients: Some further results. Biometrika, 65(2):448–450, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335230.

Louis:1975:OAS

[Lou75] Thomas A. Louis. Optimal allocation in sequential tests comparing the means of two Gaussian populations. Biometrika, 62 (2):359–369, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335373. See correction [Lou76].

Louis:1976:CAO

[Lou76] T. A. Louis. Corrections and amendments: “Optimal allocation in sequential tests comparing the means of two Gaussian populations” (Biometrika 62 (1975), no. 2, 359–369). Biometrika, 63(1):218, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335114. See

[Lou75].

Loynes:1979:NPU

[Loy79] R. M. Loynes. A note on Prescott’s upper bound for normed residuals. Biometrika, 66(2):387–389, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335675.

Lin:1974:DMI

[LS74] Pi Erh Lin and Lawrence E. Stivers. On difference of means with incomplete data. Biometrika, 61(2):325–334, August 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334361.

Li:1975:SPT

[LS75a] Loretta Li and William R. Schucany. Some properties of a test for concordance of two groups of rankings. Biometrika, 62(2):417–423, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335382.

Liao:1975:SSC

[LS75b] Henry Liao and J. Sedransk. Sequential sampling for the comparison of domain means. Biometrika, 62(3):690–693, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335533.

Lewis:1976:SNP

[LS76a] P. A. W. Lewis and G. S. Shedler. Simulation of nonhomogeneous Poisson processes with log linear rate function. Biometrika, 63(3):501–505, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335727.

Locke:1976:UST

[LS76b] Charles Locke and John D. Spurrier. The use of U-statistics for testing normality against nonsymmetric alternatives. Biometrika, 63(1):143–147, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335095. See correction [LS77a].

Locke:1977:CAU

[LS77a] Charles Locke and John D. Spurrier. Corrections and amendments: “The use of U-statistics for testing normality against nonsymmetric alternatives” (Biometrika 63 (1976), no. 1, 143–147). Biometrika, 64(1):175, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335797. See [LS76b].

Locke:1977:UST

[LS77b] Charles Locke and John D. Spurrier. The use of U-statistics for testing normality against alternatives with both tails heavy or both tails light. Biometrika, 64(3):638–640, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345348.

Lucas:1977:ECP

[LS77c] H. A. Lucas and G. A. F. Seber. Estimating coverage and particle density using the line intercept method. Biometrika, 64(3): 618–622, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345342.

Langenberg:1979:NDH

[LS79] Patricia Langenberg and R. Srinivasan. Null distribution of the Hollander–Proschan statistic for decreasing mean residual life. Biometrika, 66(3):679–680, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335194.

Li:1979:SSP

[LSG79] Shou-Hua Li, Richard M. Simon, and John J. Gart. Small sample properties of the Mantel–Haenszel test. Biometrika, 66(1):181–183, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335263. See

correction [LSG83].

Li:1983:ACS

[LSG83] Shou-Hua Li, Richard M. Simon, and John J. Gart. Amendments and corrections: ‘Small Sample Properties of the Mantel–Haenszel Test’. Biometrika, 70(1):304, April 1983. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335977. See [LSG79].

Lagakos:1978:SMM

[LSZ78] S. W. Lagakos, C. J. Sommer, and M. Zelen. Semi-Markov models for partially censored data. Biometrika, 65(2):311–317, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335209.

Lucas:1977:DEV

[Luc77] James M. Lucas. Design efficiencies for varying numbers of centre points. Biometrika, 64(1):145–147, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335786.

Lagakos:1978:MCS

[LW78] S. W. Lagakos and J. S. Williams. Models for censored survival analysis: A cone class of variable-sum models. Biometrika, 65 (1):181–189, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335293.

Lwin:1975:EFD

[Lwi75] Thaung Lwin. Exponential family distribution with a truncation parameter. Biometrika, 62(1):218–220, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334509.

Macdonald:1970:SIF

[Mac70] P. D. M. Macdonald. Statistical inference from the fraction labelled mitoses curve. Biometrika, 57(3):489–503, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334767.

MacNeill:1972:COP

[Mac72] Ian B. MacNeill. Censored overlapping Poisson counts. Biometrika, 59(2):427–434, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334587.

MacLean:1974:ETE

[Mac74a] Charles J. MacLean. Estimation and testing of an exponential polynomial rate function within the nonstationary Poisson process. Biometrika, 61(1):81–85, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334287.

MacNeill:1974:TPC

[Mac74b] Ian B. MacNeill. Tests for periodic components in multiple time series. Biometrika, 61(1):57–70, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334285.

MacKenzie:1977:SHP

[Mac77a] Donald A. MacKenzie. Studies in the history of probability and statistics. XXXVI. Arthur Black: a forgotten pioneer of mathematical statistics. Biometrika, 64(3):613–616, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345340.

MacNeill:1977:TWS

[Mac77b] Ian B. MacNeill. A test of whether several time series share common periodicities. Biometrika, 64(3):495–508, December 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345325.

Manly:1970:CWT

[Man70a] B. F. J. Manly. The choice of a Wald test on the mean of a normal population. Biometrika, 57(1):91–95, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334939.

Manly:1970:DDS

[Man70b] B. F. J. Manly. On the distribution of the decisive sample number of certain sequential tests. Biometrika, 57(2):367–376, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334845.

Mardia:1970:MMS

[Mar70a] K. V. Mardia. Measures of multivariate skewness and kurtosis with applications. Biometrika, 57(3):519–530, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334770.

Maruyama:1970:RDG

[Mar70b] Takeo Maruyama. Rate of decrease of genetic variability in a subdivided population. Biometrika, 57(2):299–311, August 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334837.

Mardia:1971:ENS

[Mar71] K. V. Mardia. The effect of nonnormality on some multivariate tests and robustness to nonnormality in the linear model. Biometrika, 58(1):105–121, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334321.

Marcus:1976:PST

[Mar76a] Ruth Marcus. The powers of some tests of the equality of normal means against an ordered alternative. Biometrika, 63(1):177–183, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335100.

Mardia:1976:LCC

[Mar76b] K. V. Mardia. Linear-circular correlation coefficients and rhythmometry. Biometrika, 63(2):403–405, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335637.

Margolin:1977:DIS

[Mar77] Barry H. Margolin. The distribution of internally Studentized statistics via Laplace transform inversion. Biometrika, 64(3):573– 582, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2345335.

Maritz:1979:NER

[Mar79a] J. S. Maritz. A note on exact robust confidence intervals for location. Biometrika, 66(1):163–166, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335258.

Martin:1979:SLP

[Mar79b] R. J. Martin. A subclass of lattice processes applied to a problem in planar sampling. Biometrika, 66(2):209–217, August 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335651.

Matthews:1970:CLT

[Mat70] Jane P. Matthews. A central limit theorem for absorbing Markov chains. Biometrika, 57(1):129–139, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334943.

Morrison:1973:PLR

[MB73] Donald F. Morrison and Dinesh S. Bhoj. Power of the likelihood ratio test on the mean vector of the multivariate normal distribution with missing observations. Biometrika, 60(2):365–368, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334549.

McClave:1973:BAA

[McC73] James T. McClave. On the bias of autoregressive approximations to moving averages. Biometrika, 60(3):599–605, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335010.

McCullagh:1977:LMP

[McC77] P. McCullagh. A logistic model for paired comparisons with ordered categorical data. Biometrika, 64(3):449–453, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345320.

McCullagh:1978:CPM

[McC78] P. McCullagh. A class of parametric models for the analysis of square contingency tables with ordered categories. Biometrika, 65(2):413–418, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335224.

McKeon:1974:ADH

[McK74] James J. McKeon. F approximations to the distribution of Hotelling’s T_{0}^{2}. Biometrika, 61(2):381–383, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334369.

McKay:1977:SPV

[McK77] R. J. McKay. Simultaneous procedures for variable selection in multiple discriminant analysis. Biometrika, 64(2):283–290, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335695.

McKinlay:1978:ENS

[McK78] Sonja M. McKinlay. The effect of nonzero second-order interaction on combined estimators of the odds ratio. Biometrika, 65(1):191– 202, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335294.

McLachlan:1974:ADC

[McL74] G. J. McLachlan. The asymptotic distributions of the conditional error rate and risk in discriminant analysis. Biometrika, 61(1):131– 135, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334294.

McLaren:1975:CPJ

[McL75] A. D. McLaren. Comments on paper by J. D. Kalbfleisch: Some personal comments on sufficiency and conditionality. Biometrika, 62(2):264–268, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335360. See [Kal75b].

McLachlan:1976:BAE

[McL76] G. J. McLachlan. The bias of the apparent error rate in discriminant analysis. Biometrika, 63(2):239–244, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335615.

McLeod:1977:IBJ

[McL77] A. I. McLeod. Improved Box-Jenkins estimators. Biometrika, 64(3):531–534, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345329.

McNeil:1972:SSE

[McN72] Donald R. McNeil. On the simple stochastic epidemic. Biometrika, 59(2):494–497, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334602.

Maag:1971:KST

[MD71] Urs R. Maag and Ghislaine Dicaire. On Kolmogorov–Smirnov type one-sample statistics. Biometrika, 58(3):653–656, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334402.

Mudholkar:1974:ELH

[MDS74] G. S. Mudholkar, M. L. Davidson, and P. Subbaiah. Extended linear hypotheses and simultaneous tests in multivariate analysis of variance. Biometrika, 61(3):467–477, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334728.

Mardia:1976:BIM

[MEA76] K. V. Mardia and S. A. M. El-Atoum. Bayesian inference for the von Mises–Fisher distribution. Biometrika, 63(1):203–206, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335106.

Meeker:1978:STI

[Mee78] William Q. Meeker. Sequential tests of independence for 2 × 2 contingency tables. Biometrika, 65(1):85–90, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335281.

Meeks:1979:CEF

[Mee79] S. L. Meeks. Conditional estimation following tests. Biometrika, 66(3):668–671, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335191.

Meshalkin:1973:MNS

[Mes73] L. D. Meshalkin. Maximum number of statistically significant associations between risk factors and the frequency of an event in observational studies. Biometrika, 60(1):175–178, April 1973. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334919.

Mann:1972:AOC

[MG72] Nancy R. Mann and Frank E. Grubbs. Approximately optimum confidence bounds on series system reliability for exponential time to failure data. Biometrika, 59(1):191–204, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334630.

Mehta:1973:TEM

[MG73] J. S. Mehta and John Gurland. A test for equality of means in the presence of correlation and missing values. Biometrika, 60(1):211– 213, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334927.

Mudholkar:1978:RSL

[MG78] Govind S. Mudholkar and E. Olusegun George. A remark on the shape of the logistic distribution. Biometrika, 65(3):667–668, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335929.

Myers:1971:AAT

[MH71] Raymond H. Myers and Richard B. Howe. On alternative approximate F tests for hypotheses involving variance components. Biometrika, 58(2):393–396, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334527.

McKean:1978:RAG

[MH78] Joseph W. McKean and Thomas P. Hettmansperger. A robust analysis of the general linear model based on one step R-estimates. Biometrika, 65(3):571–579, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335909.

Mendoza:1977:TPS

[MI77] Gaston Mendoza and Boris Iglewicz. A three-phase sequential model for clinical trials. Biometrika, 64(2):201–205, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335685.

Mickey:1975:ATP

[Mic75] M. R. Mickey. Approximate tail probabilities for Student’s t distribution. Biometrika, 62(1):216–217, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334508.

Miller:1973:NEM

[Mil73] Rupert G. Miller. Nonparametric estimators of the mean tolerance in bioassay. Biometrika, 60(3):535–542, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335002.

Miller:1974:JR

[Mil74] Rupert G. Miller. The jackknife — a review. Biometrika, 61(1):1– 15, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334280.

See [Que49, Que56, Efr79].

Miller:1976:LSR

[Mil76] Rupert G. Miller. Least squares regression with censored data. Biometrika, 63(3):449–464, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335722.

Mathai:1979:EPP

[MK79] A. M. Mathai and R. S. Katiyar. Exact percentage points for testing independence. Biometrika, 66(2):353–356, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335670.

Martz:1974:EBE

[ML74a] H. F. Martz and M. G. Lian. Empirical Bayes estimation of the binomial parameter. Biometrika, 61(3):517–523, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334734.

Morris:1974:DAD

[ML74b] J. E. Morris and P. J. Laycock. Discriminant analysis of directional data. Biometrika, 61(2):335–341, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334362.

Margolin:1976:TKS

[MM76] Barry H. Margolin and Willi Maurer. Tests of the Kolmogorov– Smirnov type for exponential data with unknown scale, and related problems. Biometrika, 63(1):149–160, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335096.

Mehrotra:1974:UEP

[MN74] K. G. Mehrotra and P. Nanda. Unbiased estimation of parameters by order statistics in the case of censored samples. Biometrika, 61(3):601–606, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334745.

Mohn:1979:CEM

[Moh79] Erik Mohn. Confidence estimation of measures of location in the log normal distribution. Biometrika, 66(3):567–575, December 1979.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335177.

Molinari:1977:DCS

[Mol77] Luciano Molinari. Distribution of the chi-squared test in nonstandard situations. Biometrika, 64(1):115–121, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335780.

Moore:1973:NSG

[Moo73] David S. Moore. A note on Srinivasan’s goodness-of-fit test. Biometrika, 60(1):209–211, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334926.

Moran:1970:NSC

[Mor70a] P. A. P. Moran. A note on serial correlation coefficients. Biometrika, 57(3):670–673, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334790.

Moran:1970:AOT

[Mor70b] P. A. P. Moran. On asymptotically optimal tests of composite hypotheses. Biometrika, 57(1):47–55, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334934.

Moran:1973:APH

[Mor73a] P. A. P. Moran. Asymptotic properties of homogeneity tests. Biometrika, 60(1):79–85, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334908.

Morrison:1973:TEM

[Mor73b] Donald F. Morrison. A test for equality of means of correlated variates with missing data on one response. Biometrika, 60(1):101– 105, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334910.

Moran:1975:EEA

[Mor75a] M. A. Moran. On the expectation of errors of allocation associated with a linear discriminant function. Biometrika, 62(1):141–148, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334496.

Moran:1975:ESE

[Mor75b] P. A. P. Moran. The estimation of standard errors in Monte Carlo simulation experiments. Biometrika, 62(1):1–4, April 1975. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334481.

Morton:1978:RAL

[Mor78] R. Morton. Regression analysis of life tables and related nonparametric tests. Biometrika, 65(2):329–333, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335211.

Moran:1979:CPR

[Mor79] P. A. P. Moran. The closest pair of N random points on the surface of a sphere. Biometrika, 66(1):158–162, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335257.

Mardia:1978:SCC

[MP78] Kanti V. Mardia and Madan L. Puri. A spherical correlation coefficient robust against scale. Biometrika, 65(2):391–395, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335219.

Marcus:1976:CTP

[MPG76] Ruth Marcus, Eric Peritz, and K. R. Gabriel. On closed testing procedures with special reference to ordered analysis of variance. Biometrika, 63(3):655–660, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335748.

Miller:1970:ESM

[MR70] K. S. Miller and M. M. Rochwarger. Estimation of spectral moments of time series. Biometrika, 57(3):513–517, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334769.

Metropolis:1953:ESC

[MRR^{+}53] Nicholas Metropolis, Arianna W. Rosenbluth, Marshall N. Rosenbluth, Augusta H. Teller, and Edward Teller. Equation of state calculations by fast computing machines. Journal of Chemical Physics, 21(6):1087–1092, June 1953. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.

org/link/doi/10.1063/1.1699114;   http://pubs.acs.org/
cgi-bin/chemport/version=1.0&coi=1:CAS:528:DyaG3sXltlKhsw%
253D%253D&pissn=0095-2338&pyear=2005&&md5=2eba6ee7d50b361b924c3ff8efeda4b1.

This article introduces the Metropolis algorithm, which the journal Computing in Science and Engineering cited in the top 10 algorithms having the “greatest influence on the development and practice of science and engineering in the 20th Century.” See [BBLS00, BS00], and the Hasting–Metropolis generalization in [Has70]. See also [Bar65, Pes73, Hit03]. According to [Har10, page 263], this paper has been cited more than 10,000 times.

Mehta:1970:BFP

[MS70] J. S. Mehta and R. Srinivasan. On the Behrens–Fisher problem. Biometrika, 57(3):649–655, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334783. See correction [Pfa74a].

Mardia:1975:MMT

[MS75] K. V. Mardia and T. W. Sutton. On the modes of a mixture of two von Mises distributions. Biometrika, 62(3):699–701, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335536.

Mikulski:1976:VBU

[MS76] Piotr W. Mikulski and Paul J. Smith. A variance bound for unbiased estimation in inverse sampling. Biometrika, 63(1):216–217, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335111.

McDonald:1967:RSM

[MT67] B. J. McDonald and W. A. Thompson, Jr. Rank sum multiple comparisons in one- and two-way classifications. Biometrika, 54 (3/4):487–497, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2335040. See correction [MT72].

McDonald:1972:CAR

[MT72] B. J. McDonald and W. A. Thompson, Jr. Corrections and amendments: “Rank sum multiple comparisons in one- and two-way classifications” (Biometrika 54 (1967), 487–497). Biometrika, 59 (3):699, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334827. See [MT67].

Morgan:1977:CIM

[MT77] B. J. T. Morgan and D. M. Titterington. A comparison of iterative methods for obtaining maximum likelihood estimates in contingency tables with a missing diagonal. Biometrika, 64(2):265–269, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335692.

Muirhead:1974:BDF

[Mui74] R. J. Muirhead. Bounds for the distribution functions of the extreme latent roots of a sample covariance matrix. Biometrika, 61(3):641–642, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334754.

Mulholland:1970:SSD

[Mul70] H. P. Mulholland. On singularities of sampling distributions, in particular for ratios of quadratic forms. Biometrika, 57(1):155–174, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334945.

Mulholland:1977:NDS

[Mul77] H. P. Mulholland. On the null distribution of \sqrt{b_{1}} for samples of size at most 25, with tables. Biometrika, 64(2):401–409, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335708.

Murray:1977:NEP

[Mur77] Gordon D. Murray. A note on the estimation of probability density functions. Biometrika, 64(1):150–152, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335788.

Murray:1979:EMN

[Mur79] Gordon D. Murray. The estimation of multivariate normal density functions using incomplete data. Biometrika, 66(2):375–380, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335673.

Muller:1976:PSF

[MV76] Peter-H. Müller and Heiko Vahl. Pearson’s system of frequency curves whose left boundary and first three moments are known. Biometrika, 63(1):191–194, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335102.

McNeil:1977:LPA

[MW77] Donald R. McNeil and George H. Weiss. A large population approach to estimation of parameters in Markov population models. Biometrika, 64(3):553–558, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345332.

Naik:1974:GBF

[Nai74] Umesh D. Naik. On a generalization of the Behrens-Fisher problem. Biometrika, 61(3):636–640, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334753.

Naik:1975:TEM

[Nai75] Umesh D. Naik. On testing equality of means of correlated variables with incomplete data. Biometrika, 62(3):615–622, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335518.

Neave:1970:SAS

[Nea70] Henry R. Neave. Spectral analysis of a stationary time series using initially scarce data. Biometrika, 57(1):111–122, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334941.

Nelson:1977:TTO

[Nel77] Lloyd S. Nelson. Tables for testing ordered alternatives in an analysis of variance. Biometrika, 64(2):335–338, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335701.

Newbold:1971:OAS

[New71] Paul Newbold. Optimum allocation in stratified two-phase sampling for proportions. Biometrika, 58(3):681–683, December 1971.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334408.

Newbold:1974:ELF

[New74] Paul Newbold. The exact likelihood function for a mixed autoregressive-moving average process. Biometrika, 61(3):423–426, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334723.

Nicholls:1979:ELF

[NH79] D. F. Nicholls and A. D. Hall. The exact likelihood function of multivariate autoregressive-moving average models. Biometrika, 66(2):259–264, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335656.

Nicholson:1970:ELP

[Nic70] W. L. Nicholson. Estimation of linear properties of particle size distributions. Biometrika, 57(2):273–297, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334836.

Nicholls:1973:FDE

[Nic73] D. F. Nicholls. Frequency domain estimation procedures for linear models. Biometrika, 60(1):202–205, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334924.

Nicholls:1976:EEV

[Nic76] D. F. Nicholls. The efficient estimation of vector linear time series models. Biometrika, 63(2):381–390, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335633.

Nicholls:1977:CEM

[Nic77] D. F. Nicholls. A comparison of estimation methods for vector linear time series models. Biometrika, 64(1):85–90, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335775.

Nagarsenker:1973:DLR

[NP73] B. N. Nagarsenker and K. C. S. Pillai. Distribution of the likelihood ratio criterion for testing a hypothesis specifying a covariance matrix. Biometrika, 60(2):359–364, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334548.

Nebenzahl:1972:PWS

[NS72] Elliott Nebenzahl and Milton Sobel. Play-the-winner sampling for a fixed sample size binomial selection problem. Biometrika, 59(1): 1–8, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334609.

Nimmo-Smith:1979:LRS

[NS79] Ian Nimmo-Smith. Linear regressions and sphericity. Biometrika, 66(2):390–392, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335676.

Oakes:1977:AIC

[Oak77] David Oakes. The asymptotic information in censored survival data. Biometrika, 64(3):441–448, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345319.

Odeh:1972:PJS

[Ode72] Robert E. Odeh. On the power of Jonckheere’s k-sample test against ordered alternatives. Biometrika, 59(2):467–471, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334593.

Oden:1973:SCI

[Odé73] Anders Odén. Simultaneous confidence intervals in inverse linear regression. Biometrika, 60(2):339–343, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334545.

Ofosu:1973:TSP

[Ofo73] J. B. Ofosu. A two-sample procedure for selecting the population with the largest mean from several normal populations with unknown variances. Biometrika, 60(1):117–124, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334913. See correction [Ofo75].

Ofosu:1975:CAT

[Ofo75] J. B. Ofosu. Corrections and amendments: “A Two-Sample Procedure for Selecting the Population With the Largest Mean from Several Normal Populations with Unknown Variances”. Biometrika, 62(1):221, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2334511. See [Ofo73].

OHagan:1973:BEC

[O’H73] A. O’Hagan. Bayes estimation of a convex quadratic. Biometrika, 60(3):565–571, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335006.

OHagan:1976:PJM

[O’H76] A. O’Hagan. On posterior joint and marginal modes. Biometrika, 63(2):329–333, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335627.

OHagan:1976:BES

[OL76] A. O’Hagan and Tom Leonard. Bayes estimation subject to uncertainty about parameter constraints. Biometrika, 63(1):201–203, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335105.

Olkin:1972:MPD

[Olk72] Ingram Olkin. Monotonicity properties of Dirichlet integrals with applications to the multinomial distribution and the analysis of variance. Biometrika, 59(2):303–307, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334574.

ONeill:1979:DFE

[O’N79] Terence J. O’Neill. Distribution-free estimation of cure time. Biometrika, 66(1):184–187, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335264.

Orban:1978:OCS

[OW78] John Orban and Douglas A. Wolfe. Optimality criteria for the selection of a partially sequential indicator set. Biometrika, 65 (2):357–362, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335215.

Paloheimo:1971:TS

[Pal71] J. E. Paloheimo. On a theory of search. Biometrika, 58(1):61–75, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334317.

Paloheimo:1972:SBP

[Pal72] J. E. Paloheimo. A spatial bivariate Poisson distribution. Biometrika, 59(2):489–492, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334600.

Patterson:1970:NCD

[Pat70] H. D. Patterson. Nonadditivity in change-over designs for a quantitative factor at four levels. Biometrika, 57(3):537–549, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334772.

Patterson:1973:QCD

[Pat73] H. D. Patterson. Quenouille’s changeover designs. Biometrika, 60(1):33–45, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334903.

Patefield:1978:UUR

[Pat78] W. M. Patefield. The unreplicated ultrastructural relation: Large sample properties. Biometrika, 65(3):535–540, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335905.

Prentice:1978:RSF

[PB78] R. L. Prentice and N. E. Breslow. Retrospective studies and failure time models. Biometrika, 65(1):153–158, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335290.

Pearce:1974:BCE

[PCdCM74] S. C. Pearce, T. Caliński, and T. F. de C. Marshall. The basic contrasts of an experimental design with special reference to the analysis of data. Biometrika, 61(3):449–460, December 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334726.

Pearson:1977:TDN

[PDB77] E. S. Pearson, R. B. D’Agostino, and K. O. Bowman. Tests for departure from normality: Comparison of powers. Biometrika, 64(2):231–246, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335689.

Pham-Dinh-Tuan:1977:EPC

[PDT77] Pham-Dinh-Tuan. Estimation of parameters of a continuous time Gaussian stationary process with rational spectral density. Biometrika, 64(2):385–399, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335707.

Pham-Dinh-Tuan:1978:FMP

[PDT78] Pham-Dinh-Tuan. On the fitting of multivariate processes of the autoregressive-moving average type. Biometrika, 65(1):99–107, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335283.

Pearson:1945:EN

[Pea45] E. S. Pearson. Editorial note. Biometrika, 33(3):252–253, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332304. See

corrigenda [Ano48b].

Pearce:1970:EBD

[Pea70] S. C. Pearce. The efficiency of block designs in general. Biometrika, 57(2):339–346, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334841. See correction [Pea72].

Pearce:1971:PBE

[Pea71] S. C. Pearce. Precision in block experiments. Biometrika, 58 (1):161–167, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334327.

Pearce:1972:CAE

[Pea72] S. C. Pearce. Corrections and amendments: “The Efficiency of Block Designs in General”. Biometrika, 59(2):498, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334605. See [Pea70].

Pedersen:1979:ACD

[Ped79] Bo V. Pedersen. Approximating conditional distributions by the mixed Edgeworth-saddlepoint expansion. Biometrika, 66(3):597– 604, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335181.

Peers:1971:LRA

[Pee71] H. W. Peers. Likelihood ratio and associated test criteria. Biometrika, 58(3):577–587, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334392.

Peers:1978:SOS

[Pee78] H. W. Peers. Second-order sufficiency and statistical invariants. Biometrika, 65(3):489–496, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335899.

Peskun:1973:OMC

[Pes73] P. H. Peskun. Optimum Monte–Carlo sampling using Markov chains. Biometrika, 60(3):607–612, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335011.

Pesotchinsky:1975:OQO

[Pes75] L. L. Pesotchinsky. D-optimum and quasi-D-optimum secondorder designs on a cube. Biometrika, 62(2):335–340, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335369. See correction

[Pes76].

Pesotchinsky:1976:CAO

[Pes76] L. L. Pesotchinsky. Corrections and amendments: “D-optimum and quasi-D-optimum second-order designs on a cube” (Biometrika 62 (1975), 335–340). Biometrika, 63(2):412, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335642. See [Pes75].

Peto:1972:RTM

[Pet72] Richard Peto. Rank tests of maximal power against Lehmann-type alternatives. Biometrika, 59(2):472–475, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334594.

Pettitt:1976:CMS

[Pet76a] A. N. Pettitt. Cramér–von Mises statistics for testing normality with censored samples. Biometrika, 63(3):475–481, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335724.

Pettitt:1976:TSA

[Pet76b] A. N. Pettitt. A two-sample Anderson-Darling rank statistic. Biometrika, 63(1):161–168, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335097.

Pettitt:1977:TED

[Pet77] A. N. Pettitt. Tests for the exponential distribution with censored data using Cramér-von Mises statistics. Biometrika, 64(3):629–632, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345345.

Pettitt:1978:GCM

[Pet78] A. N. Pettitt. Generalized Cramér-von Mises statistics for the gamma distribution. Biometrika, 65(1):232–235, April 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335303.

Pfanzagl:1974:CAB

[Pfa74a] J. Pfanzagl. Corrections and amendments: “On the Behrens– Fisher Problem”. Biometrika, 61(3):647, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334760. See [MS70].

Pfanzagl:1974:BFP

[Pfa74b] J. Pfanzagl. On the Behrens-Fisher problem. Biometrika, 61(1):39– 47, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334283.

Pillai:1969:EDW

[PG69] K. C. Sreedharan Pillai and Arjun K. Gupta. On the exact distribution of Wilks’s criterion. Biometrika, 56(1):109–118, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334695. See cor-

rection [PG70].

Pillai:1970:CED

[PG70] K. C. Sreedharan Pillai and Arjun K. Gupta. Corrections: “On the Exact Distribution of Wilks’s Criterion”. Biometrika, 57(1): 225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334963.

See [PG69].

Phipps:1971:IHP

[Phi71] Mary C. Phipps. An identity in Hermite polynomials. Biometrika, 58(1):219–222, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334334.

Phillips:1978:ESA

[Phi78] P. C. B. Phillips. Edgeworth and saddlepoint approximations in the first-order noncircular autoregression. Biometrika, 65(1):91–98, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335282.

Paulson:1975:EPS

[PHL75] A. S. Paulson, E. W. Holcomb, and R. A. Leitch. The estimation of the parameters of the stable laws. Biometrika, 62(1):163–170, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334499.

Pierce:1971:LSE

[Pie71] David A. Pierce. Least squares estimation in the regression model with autoregressive-moving average errors. Biometrika, 58(2):299– 312, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2334518.

Pierce:1972:LSE

[Pie72] David A. Pierce. Least squares estimation in dynamic-disturbance time series models. Biometrika, 59(1):73–78, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334616.

Phadke:1978:CEL

[PK78] M. S. Phadke and G. Kedem. Computation of the exact likelihood function of multivariate moving average models. Biometrika, 65(3):511–519, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335902.

Pierce:1979:TGF

[PK79] Donald A. Pierce and Kenneth J. Kopecky. Testing goodness of fit for the distribution of errors in regression models. Biometrika, 66 (1):1–5, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335235.

Plackett:1972:SHP

[Pla72] R. L. Plackett. Studies in the history of probability and statistics. XXIX. The discovery of the method of least squares. Biometrika, 59(2):239–251, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334569.

Plackett:1977:MTT

[Pla77] R. L. Plackett. The marginal totals of a 2 × 2 table. Biometrika, 64(1):37–42, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335767.

Pocock:1977:GSM

[Poc77] Stuart J. Pocock. Group sequential methods in the design and analysis of clinical trials. Biometrika, 64(2):191–199, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335684.

Pollock:1975:STR

[Pol75] K. H. Pollock. A K-sample tag-recapture model allowing for unequal survival and catchability. Biometrika, 62(3):577–583, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335513.

Pearson:1975:RBS

[PP75] E. S. Pearson and N. W. Please. Relation between the shape of population distribution and the robustness of four simple test statistics. Biometrika, 62(2):223–241, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335354.

Paul:1978:ISP

[PP78] S. R. Paul and R. L. Plackett. Inference sensitivity for Poisson mixtures. Biometrika, 65(3):591–602, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335911.

Prentice:1979:LDI

[PP79] R. L. Prentice and R. Pyke. Logistic disease incidence models and case-control studies. Biometrika, 66(3):403–411, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335158.

Preece:1973:ODT

[PPK73] D. A. Preece, S. C. Pearce, and J. R. Kerr. Orthogonal designs for three-dimensional experiments. Biometrika, 60(2):349–358, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334547.

Persson:1977:SHE

[PR77] Tore Persson and Holger Rootzen. Simple and highly efficient estimators for a Type I censored normal sample. Biometrika, 64 (1):123–128, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335781.

Prescott:1970:ESD

[Pre70] P. Prescott. Estimation of the standard deviation of a normal population from doubly censored samples using normal scores. Biometrika, 57(2):409–419, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334850.

Prescott:1971:DNS

[Pre71a] P. Prescott. Distribution of the normal scores estimator of the standard deviation of a normal population. Biometrika, 58(3):631–636, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334398.

Prescott:1971:USR

[Pre71b] P. Prescott. Use of a simple range-type estimator of σ in tests of hypotheses. Biometrika, 58(2):333–340, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334521.

Preston:1971:EBP

[Pre71c] P. F. Preston. An empirical Bayes problem with a Markovian parameter. Biometrika, 58(3):535–543, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334388.

Prentice:1973:ESC

[Pre73] R. L. Prentice. Exponential survivals with censoring and explanatory variables. Biometrika, 60(2):279–288, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334539.

Prentice:1974:DFM

[Pre74a] R. L. Prentice. Degrees-of-freedom modifications for F tests based on nonnormal errors. Biometrika, 61(3):559–563, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334739.

Prentice:1974:LGM

[Pre74b] R. L. Prentice. A log gamma model and its maximum likelihood estimation. Biometrika, 61(3):539–544, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334737.

Prescott:1974:NTS

[Pre74c] P. Prescott. Normalizing transformations of Student’s t distribution. Biometrika, 61(1):177–180, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334301.

Prescott:1974:VCO

[Pre74d] P. Prescott. Variances and covariances of order statistics from the gamma distribution. Biometrika, 61(3):607–613, December 1974.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334746.

Prentice:1975:DAS

[Pre75] R. L. Prentice. Discrimination among some parametric models. Biometrika, 62(3):607–614, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335517.

Prescott:1976:CTN

[Pre76] P. Prescott. Comparison of tests for normality using stylized sensitivity surfaces. Biometrika, 63(2):285–289, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335621.

Prentice:1978:LRT

[Pre78] R. L. Prentice. Linear rank tests with right censored data. Biometrika, 65(1):167–179, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335292. See correction [Pre83].

Prentice:1979:PIR

[Pre79a] M. J. Prentice. On the problem of m incomplete rankings. Biometrika, 66(1):167–170, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335259.

Prescott:1979:MDE

[Pre79b] P. Prescott. A mean difference estimator of standard deviation in asymmetrically censored normal samples. Biometrika, 66(3): 684–686, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335196.

Prentice:1983:ACL

[Pre83] R. L. Prentice. Amendments and corrections: “Linear rank tests with right censored data” [Biometrika 65 (1978), no. 1, 167–179; MR0497517 (80a:62060)]. Biometrika, 70(1):304, April 1983. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335978. See [Pre78].

Pruscha:1976:MLE

[Pru76] Helmut Pruscha. Maximum likelihood estimation in linear learning models. Biometrika, 63(3):537–542, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335732.

Pradhan:1973:PWS

[PS73] Meena Pradhan and Y. S. Sathe. Play-the-winner sampling for a fixed sample size with curtailment. Biometrika, 60(2):424–427, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334559.

Pettitt:1976:MCM

[PS76] A. N. Pettitt and M. A. Stephens. Modified Cramér-von Mises statistics for censored data. Biometrika, 63(2):291–298, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335622.

Pearson:1970:SNR

[PT70] E. S. Pearson and M. L. Tiku. Some notes on the relationship between the distributions of central and non-central F. Biometrika, 57(1):175–179, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334946.

Patterson:1971:RIB

[PT71] H. D. Patterson and R. Thompson. Recovery of inter-block information when block sizes are unequal. Biometrika, 58(3):545–554, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334389.

Pusey:1970:ROO

[Pus70] Mary C. Pusey. Representation of one orthonormal set in terms of another. Biometrika, 57(2):445–450, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334854.

Pandit:1975:UEP

[PW75] S. M. Pandit and S. M. Wu. Unique estimates of the parameters of a continuous stationary stochastic process. Biometrika, 62(2):497–501, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335392.

Patterson:1976:NCR

[PW76] H. D. Patterson and E. R. Williams. A new class of resolvable incomplete block designs. Biometrika, 63(1):83–92, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335087.

Quenouille:1949:ATC

[Que49] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246.

URL http://www.jstor.org/stable/2983696. See extension in

[Que56, Tuk58], and reviews [Mil74, Efr79].

Quenouille:1956:NBE

[Que56] M. H. Quenouille. Notes on bias in estimation. Biometrika, 43 (3–4):353–360, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2332914. This extends earlier work [Que49, Tuk58] on what is now known as the Quenouille–Tukey jacknife technique. See reviews [Mil74, Efr79].

Quine:1970:NMS

[Qui70] M. P. Quine. A note on the moment structure of the multitype Galton–Watson process. Biometrika, 57(1):219–222, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334958.

Ramey:1979:SPS

[RA79] James T. Ramey and Khursheed Alam. A sequential procedure for selecting the most probable multinomial event. Biometrika, 66(1):171–173, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335260.

Rabinovitch:1970:SHP

[Rab70] Nachum L. Rabinovitch. Studies in the history of probability and statistics. XXIV. Combinations and probability in Rabbinic literature. Biometrika, 57(1):203–205, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334951.

Ramasubban:1972:ADN

[Ram72] T. A. Ramasubban. An approximate distribution of a noncircular serial correlation coefficient. Biometrika, 59(1):79–84, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334617.

Ramsey:1979:PML

[Ram79] Fred L. Ramsey. Parametric models for line transect surveys. Biometrika, 66(3):505–512, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335169.

Rankin:1974:HMM

[Ran74] Neil O. Rankin. The harmonic mean method for one-way and twoway analyses of variance. Biometrika, 61(1):117–122, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334292.

Rao:1973:CDS

[Rao73a] J. N. K. Rao. Corrections: “On double sampling for stratification and analytical surveys” (Biometrika 60 (1973), 125–133). Biometrika, 60(3):669, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335022. See [Rao73b].

Rao:1973:DSS

[Rao73b] J. N. K. Rao. On double sampling for stratification and analytical surveys. Biometrika, 60(1):125–133, April 1973. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334914. See correction

[Rao73a].

Rayment:1971:PEM

[Ray71] P. R. Rayment. The performance of established and modified compound decision rules. Biometrika, 58(1):183–194, April 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334330.

Read:1973:MRC

[Rea73] Campbell B. Read. On minimizing the risk in certain sequential tests. Biometrika, 60(3):509–515, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334999.

Reeves:1972:DML

[Ree72] J. E. Reeves. The distribution of the maximum likelihood estimator of the parameter in the first-order autoregressive series. Biometrika, 59(2):387–394, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334583.

Renshaw:1972:BDM

[Ren72] E. Renshaw. Birth, death and migration processes. Biometrika, 59(1):49–60, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334614.

Robertson:1970:BAM

[RF70] C. A. Robertson and J. G. Fryer. The bias and accuracy of moment estimators. Biometrika, 57(1):57–65, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334935.

Rencher:1975:SEB

[RK75] Alvin C. Rencher and Richard G. Krutchkoff. Some empirical Bayes estimators allowing for varying error variances. Biometrika, 62(3):643–650, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335522.

Robinson:1970:BIS

[Rob70] J. Robinson. Blocking in incomplete split plot designs. Biometrika, 57(2):347–350, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334842.

Robertson:1974:LST

[Rob74] C. A. Robertson. Large-sample theory for the linear structural relation. Biometrika, 61(2):353–359, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334365.

Robinson:1975:SCT

[Rob75] G. K. Robinson. Some counterexamples to the theory of confidence intervals. Biometrika, 62(1):155–161, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334498. See correction [Rob77].

Robinson:1977:CAS

[Rob77] G. K. Robinson. Corrections and amendments: “Some Counterexamples to the Theory of Confidence Intervals”. Biometrika, 64 (3):655, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345358. See [Rob75].

Robinson:1978:CPB

[Rob78a] G. K. Robinson. Comments on paper by B. Efron and D. V. Hinkley. Biometrika, 65(3):485, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335896. See [EH78b].

Robinson:1978:NBI

[Rob78b] G. K. Robinson. On the necessity of Bayesian inference and the construction of measures of nearness to Bayesian form. Biometrika, 65(1):49–52, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335275.

Robinson:1978:SCO

[Rob78c] J. A. Robinson. Sequential choice of an optimal dose: A prediction intervals approach. Biometrika, 65(1):75–78, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335279.

Rodriguez:1977:GBT

[Rod77] Robert N. Rodriguez. A guide to the Burr Type XII distributions. Biometrika, 64(1):129–134, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335782.

Roger:1977:STC

[Rog77] J. H. Roger. A significance test for cyclic trends in incidence data. Biometrika, 64(1):152–155, April 1977. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335789. See correction [Rog78].

Roger:1978:ACS

[Rog78] J. H. Roger. Amendments and corrections: “A Significance Test for Cyclic Trends in Incidence Data”. Biometrika, 65(1):241, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335306. See

[Rog77].

Rosenberg:1973:LRR

[Ros73] Barr Rosenberg. Linear regression with randomly dispersed parameters. Biometrika, 60(1):65–72, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334906.

Rosner:1979:MLE

[Ros79] Bernard Rosner. Maximum likelihood estimation of interclass correlations. Biometrika, 66(3):533–538, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335172.

Rothery:1979:NMI

[Rot79] P. Rothery. A nonparametric measure of intraclass correlation. Biometrika, 66(3):629–639, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335185.

Royall:1970:FPS

[Roy70] Richard M. Royall. On finite population sampling theory under certain linear regression models. Biometrika, 57(2):377–387, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334846.

Royall:1976:LFF

[Roy76] Richard M. Royall. Likelihood functions in finite population sampling theory. Biometrika, 63(3):605–614, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335741.

Roy:1977:ABS

[Roy77] Roch Roy. On the asymptotic behaviour of the sample autocovariance function for an integrated moving average process.

Biometrika, 64(2):419–421, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335711.

Rao:1971:SSR

[RR71] Poduri S. R. S. Rao and J. N. K. Rao. Small sample results for ratio estimators. Biometrika, 58(3):625–630, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334397.

Raghunandanan:1970:SEP

[RS70] K. Raghunandanan and R. Srinivasan. Simplified estimation of parameters in a logistic distribution. Biometrika, 57(3):677–679, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334793.

Rao:1978:PIE

[RS78a] C. Radhakrishna Rao and Nobuo Shinozaki. Precision of individual estimators in simultaneous estimation of parameters. Biometrika, 65(1):23–30, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335272.

Ripley:1978:QTS

[RS78b] B. D. Ripley and B. W. Silverman. Quick tests for spatial interaction. Biometrika, 65(3):641–642, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335917.

Rousseau:1978:SAB

[RS78c] Pascale Rousseau and David Sankoff. Singularities in the analysis of binomial data. Biometrika, 65(3):603–608, December 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335912.

Rao:1973:STT

[RT73] T. Subba Rao and H. Tong. On some tests for the time dependence of a transfer function. Biometrika, 60(3):589–597, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335009.

Rubin:1976:IMD

[Rub76] Donald B. Rubin. Inference and missing data. Biometrika, 63 (3):581–592, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2335739. With comments by R. J. A. Little and a reply by the author.

Ruda:1974:PEF

[Rud74] M. Ruda. Parameter estimation in the first-order autoregressive process. Biometrika, 61(3):632–633, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334751.

Rubin:1975:VLC

[RW75] Donald B. Rubin and Sanford Weisberg. The variance of a linear combination of independent estimators using estimated weights. Biometrika, 62(3):708–709, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335540.

Samiuddin:1976:TEV

[SA76] M. Samiuddin and M. Atiqullah. A test for equality of variances. Biometrika, 63(1):206–208, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335107.

Samiuddin:1970:TAV

[Sam70] M. Samiuddin. On a test for an assigned value of correlation in a bivariate normal distribution. Biometrika, 57(2):461–464, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334858.

Sankoff:1971:DDS

[San71] David Sankoff. Duration of detectible synchrony in a binary branching process. Biometrika, 58(1):77–81, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334318.

Sarndal:1971:SHP

[Sär71] Carl-Erik Särndal. Studies in the history of probability and statistics. XXVII. The hypothesis of elementary errors and the Scandinavian school in statistical theory. Biometrika, 58(2):375–391, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334526.

Sarkadi:1975:CSF

[Sar75] K. Sarkadi. The consistency of the Shapiro-Francia test. Biometrika, 62(2):445–450, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335386.

Sathe:1977:SVB

[Sat77] Y. S. Sathe. Sharper variance bounds for unbiased estimation in inverse sampling. Biometrika, 64(2):425–426, August 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335713.

Saw:1958:CMS

[Saw58a] J. G. Saw. Corrigenda: “Moments of Sample Moments of Censored Samples from a Normal Population”. Biometrika, 45(3/4):587–588, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333213.

See [Saw58b].

Saw:1958:MSM

[Saw58b] J. G. Saw. Moments of sample moments of censored samples from a normal population. Biometrika, 45(1/2):211–221, June 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333057. See corrigenda

[Saw58a, Saw79].

Saw:1970:MLH

[Saw70] J. G. Saw. The multivariate linear hypothesis with nonnormal errors and a classical setting for the structure of inference in a special case. Biometrika, 57(3):531–535, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334771.

Saw:1978:FDS

[Saw78] John G. Saw. A family of distributions on the m-sphere and some hypothesis tests. Biometrika, 65(1):69–73, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335278.

Saw:1979:ACM

[Saw79] J. G. Saw. Amendments and corrections: “Moments of Sample Moments of Censored Samples from a Normal Population”. Biometrika, 66(1):194, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335267. See [Saw58b].

Shenton:1979:CPR

[SBL79] L. R. Shenton, K. O. Bowman, and H. K. Lam. Comments on a paper by R. C. Geary on standardized mean deviation. Biometrika, 66(2):400–401, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335679. See [Gea36].

Schuster:1975:EDF

[Sch75] Eugene F. Schuster. Estimating the distribution function of a symmetric distribution. Biometrika, 62(3):631–635, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335520. See correction

[Sch76].

Schuster:1976:CAE

[Sch76] E. F. Schuster. Corrections and amendments: “Estimating the distribution function of a symmetric distribution” (Biometrika 62 (1975), 631–635). Biometrika, 63(2):412, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335643. See [Sch75].

Schou:1978:ECP

[Sch78a] Geert Schou. Estimation of the concentration parameter in von Mises–Fisher distributions. Biometrika, 65(2):369–377, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335217.

Schucany:1978:CPM

[Sch78b] William R. Schucany. Comments on paper by M. Hollander and J. Sethuraman. Biometrika, 65(2):410–411, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335222. See [HS78c].

Scott:1979:ODB

[Sco79] David W. Scott. On optimal and data-based histograms. Biometrika, 66(3):605–610, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335182.

Stone:1972:BII

[SD72] M. Stone and A. P. Dawid. Un-Bayesian implications of improper Bayes inference in routine statistical problems. Biometrika, 59 (2):369–375, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334581.

Shuster:1976:TWC

[SD76] J. J. Shuster and D. J. Downing. Two-way contingency tables for complex sampling schemes. Biometrika, 63(2):271–276, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335619.

Singh:1979:BDN

[SD79] M. Singh and A. Dey. Block designs with nested rows and columns. Biometrika, 66(2):321–326, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335665.

Sobel:1975:GTN

[SE75] Milton Sobel and R. M. Elashoff. Group testing with a new goal, estimation. Biometrika, 62(1):181–193, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334502.

Seal:1977:SHP

[Sea77] Hilary L. Seal. Studies in the history of probability and statistics. XXXV. Multiple decrements or competing risks. Biometrika, 64(3):429–439, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345318.

Seber:1970:ETS

[Seb70] G. A. F. Seber. Estimating time-specific survival and reporting rates for adult birds from band returns. Biometrika, 57(2):313–318, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334838.

Seber:1971:EAS

[Seb71] G. A. F. Seber. Estimating age-specific survival rates from birdband returns when the reporting rate is constant. Biometrika, 58(3):491–497, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334384.

Seely:1975:EIA

[See75] Justus Seely. An example of an inquadmissible analysis of variance estimator for a variance component. Biometrika, 62(3):689–690, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335532.

Seheult:1978:PFS

[Seh78] Allan Seheult. On a problem in forensic science. Biometrika, 65(3):646–648, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335920.

Shafiq:1979:GAB

[SF79] M. Shafiq and W. T. Federer. Generalized N-ary balanced block designs. Biometrika, 66(1):115–123, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335250.

Schafer:1972:GFT

[SFC72] R. E. Schafer, J. M. Finkelstein, and John Collins. On a goodnessof-fit test for the exponential distribution with mean unknown. Biometrika, 59(1):222–224, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334639.

Srinivasan:1974:CMT

[SG74] R. Srinivasan and L. B. Godio. A Cramér-von Mises type statistic for testing symmetry. Biometrika, 61(1):196–198, April 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334306.

Sugg:1976:GKQ

[SH76] Merritt N. Sugg and Kermit Hutcheson. Gauss–Krawtchouk quadrature approximations for moments of multinomial type distributions. Biometrika, 63(2):395–400, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335635.

Shah:1970:DHM

[Sha70a] B. K. Shah. On the distribution of half the mean square successive difference. Biometrika, 57(1):193–198, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334949.

Shanbhag:1970:DQF

[Sha70b] D. N. Shanbhag. On the distribution of a quadratic form. Biometrika, 57(1):222–223, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334959.

Sharpe:1970:RNT

[Sha70c] K. Sharpe. Robustness of normal tolerance intervals. Biometrika, 57(1):71–78, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334937.

Shaman:1973:ICM

[Sha73] Paul Shaman. On the inverse of the covariance matrix for an autoregressive-moving average process. Biometrika, 60(1):193–196, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334921.

Sharp:1975:CLR

[Sha75] Stephen A. Sharp. Corrections to likelihood ratio tests of hypotheses concerning the parameters of Markov chains. Biometrika, 62(3):595–598, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335515.

Shaikh:1976:SIM

[Sha76a] A. Shaikh. Statistical inference for Markov chains with lumped states. Biometrika, 63(1):211–213, April 1976. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335109. See correction [Sha78b].

Sharot:1976:SJ

[Sha76b] T. Sharot. Sharpening the jackknife. Biometrika, 63(2):315–321, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335625.

Shaman:1977:SBC

[Sha77] Paul Shaman. Some Bayesian considerations in spectral estimation. Biometrika, 64(1):79–84, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335774.

Samiuddin:1978:SCB

[SHA78a] M. Samiuddin, M. Hanif, and H. Asad. Some comparisons of the Bartlett and cube root tests of homogeneity of variance. Biometrika, 65(1):218–221, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335299.

Shaikh:1978:ACS

[Sha78b] A. Shaikh. Amendments and corrections: “Statistical inference for Markov chains with lumped states” (Biometrika 63 (1976), no. 1, 211–213). Biometrika, 65(2):455, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335232. See [Sha76a].

Sheynin:1970:SHP

[She70] O. B. Sheynin. Studies in the history of probability and statistics. XXIII. Daniel Bernoulli on the Normal Law. Biometrika, 57(1):199–202, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334950.

Sherman:1971:ENA

[She71a] R. Sherman. Error of the normal approximation to the sum of N random variables. Biometrika, 58(2):396–398, August 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334528.

Sheynin:1971:SHP

[She71b] O. B. Sheynin. Studies in the history of probability and statistics. XXV. On the history of some statistical laws of distribution. Biometrika, 58(1):234–236, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334339.

Shibata:1976:SOA

[Shi76] Ritei Shibata. Selection of the order of an autoregressive model by Akaike’s information criterion. Biometrika, 63(1):117–126, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335091.

Shirahata:1977:TPC

[Shi77] S. Shirahata. Tests of partial correlation in a linear model. Biometrika, 64(1):162–164, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335792.

Shirahata:1978:AOS

[Shi78] S. Shirahata. An approach to a one-sided test in the bivariate normal distribution. Biometrika, 65(1):61–67, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335277.

Siegmund:1977:RST

[Sie77] D. Siegmund. Repeated significance tests for a normal mean. Biometrika, 64(2):177–189, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335683.

Siegmund:1978:EFS

[Sie78] D. Siegmund. Estimation following sequential tests. Biometrika, 65(2):341–349, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335213.

Siegel:1979:NCS

[Sie79] Andrew F. Siegel. The noncentral chi-squared distribution with zero degrees of freedom and testing for uniformity. Biometrika, 66(2):381–386, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335674.

Silverman:1978:CWW

[Sil78a] B. W. Silverman. Choosing the window width when estimating a density. Biometrika, 65(1):1–11, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335270.

Silvey:1978:ODM

[Sil78b] S. D. Silvey. Optimal design measures with singular information matrices. Biometrika, 65(3):553–559, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335907.

Simon:1977:NTT

[Sim77] Gary Simon. A nonparametric test of total independence based on Kendall’s Tau. Biometrika, 64(2):277–282, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335694.

Singh:1970:DRS

[Sin70] C. Singh. On the distribution of range of samples from nonnormal populations. Biometrika, 57(2):451–456, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334855.

Singh:1972:OSN

[Sin72] C. Singh. Order statistics from nonnormal populations. Biometrika, 59(1):229–233, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334642.

Singh:1978:SST

[Sin78a] Padam Singh. The selection of samples of two units with inclusion probabilities proportional to size. Biometrika, 65(2):450–454, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335231.

Sinha:1978:RTP

[Sin78b] Kishore Sinha. A resolvable triangular partially balanced incomplete block design. Biometrika, 65(3):665, December 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335927.

Singh:1979:MSE

[Sin79] R. S. Singh. Mean squared errors of estimates of a density and its derivatives. Biometrika, 66(1):177–180, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335262.

Siskind:1972:SMI

[Sis72] Victor Siskind. Second moments of inverse Wishart-matrix elements. Biometrika, 59(3):690–691, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334823.

Siskind:1976:API

[Sis76] V. Siskind. Approximate probability integrals and critical values for Bartholomew’s test for ordered means. Biometrika, 63 (3):647–654, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335747.

Slakter:1973:LVN

[Sla73] Malcolm J. Slakter. Large values for the number of groups with the Pearson chi-squared goodness-of-fit tes. Biometrika, 60(2):420–421, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334557.

Slivka:1970:OSN

[Sli70] John Slivka. A one-sided nonparametric multiple comparison control percentile test: Treatments versus control. Biometrika, 57 (2):431–438, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334852.

Shuster:1972:TWA

[SM72] J. J. Shuster and C. Miura. Two-way analysis of reciprocals. Biometrika, 59(2):478–481, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334596.

Small:1978:PSR

[Sma78] N. J. H. Small. Plotting squared radii. Biometrika, 65(3):657–658, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335924.

Smith:1973:BEO

[Smi73] A. F. M. Smith. Bayes estimates in one-way and two-way models. Biometrika, 60(2):319–329, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334543.

Smith:1975:BAI

[Smi75] A. F. M. Smith. A Bayesian approach to inference about a changepoint in a sequence of random variables. Biometrika, 62(2):407– 416, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/

2335381. See correction [Smi76].

Smith:1976:CAB

[Smi76] A. F. M. Smith. Corrections and amendments: “A Bayesian approach to inference about a change-point in a sequence of random variables” (Biometrika 62 (1975), 407–416). Biometrika, 63(1):218, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335115. See

[Smi75].

Somerville:1971:GFT

[Som71a] Paul N. Somerville. A generalization of a fundamental theorem of ranking and selection. Biometrika, 58(1):227–228, April 1971. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334336. See correction

[Som76].

Somerville:1971:TOP

[Som71b] Paul N. Somerville. A technique for obtaining probabilities of correct selection in a two-stage selection problem. Biometrika, 58 (3):615–623, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334396.

Somerville:1976:CAG

[Som76] P. N. Somerville. Corrections and amendments: “A Generalization of a Fundamental Theorem of Ranking and Selection”. Biometrika, 63(2):412, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335644. See [Som71a].

Sowden:1971:BAP

[Sow71] R. R. Sowden. Bias and accuracy of parameter estimates in a quantal response model. Biometrika, 58(3):595–603, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334394.

Sowden:1972:FOB

[Sow72] B. R. Sowden. On the first-order bias of parameter estimates in a quantal response model under alternative estimation procedures. Biometrika, 59(3):573–579, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334808.

Schucany:1978:CSF

[SPB78] William R. Schucany, William C. Parr, and John E. Boyer. Correlation structure in Farlie–Gumbel–Morgenstern distributions. Biometrika, 65(3):650–653, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335922.

Spiegelhalter:1977:TNA

[Spi77] D. J. Spiegelhalter. A test for normality against symmetric alternatives. Biometrika, 64(2):415–418, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335710.

Spjotvoll:1972:JCI

[Spj72] Emil Spjøtvoll. Joint confidence intervals for all linear functions of means in the one-way layout with unknown group variances. Biometrika, 59(3):683–685, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334820.

Spjotvoll:1977:APM

[Spj77a] Emil Spjotvoll. Alternatives to plotting C_{P} in multiple regression. Biometrika, 64(1):1–8, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2335762. See amendments and corrections [Spj78].

Spjotvoll:1977:OOP

[Spj77b] Emil Spjøtvoll. Ordering ordered parameters. Biometrika, 64 (2):327–334, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335700.

Spjotvoll:1978:ACA

[Spj78] Emil Spjotvoll. Amendments and corrections: Alternatives to Plotting C_{p} in Multiple Regression. Biometrika, 65(1):241, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335307. See [Spj77a].

Sprott:1973:NLT

[Spr73] D. A. Sprott. Normal likelihoods and their relation to large sample theory of estimation. Biometrika, 60(3):457–465, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334994.

Sprott:1975:MCS

[Spr75] D. A. Sprott. Marginal and conditional sufficiency. Biometrika, 62(3):599–605, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335516.

Sprott:1978:CPB

[Spr78] D. A. Sprott. Comments on paper by B. Efron and D. V. Hinkley. Biometrika, 65(3):485–486, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335897. See [EH78b].

Srinivasan:1970:ATG

[Sri70] R. Srinivasan. An approach to testing the goodness of fit of incompletely specified distributions. Biometrika, 57(3):605–611, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334778.

Scott:1970:NMA

[SS70] Alastair Scott and T. M. F. Smith. A note on Moran’s approximation to Student’s t. Biometrika, 57(3):681–682, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334795.

Subrahmaniam:1973:MBF

[SS73] Kocherlakota Subrahmaniam and Kathleen Subrahmaniam. On the multivariate Behrens-Fisher problem. Biometrika, 60(1):107– 111, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334911.

Scott:1975:MDS

[SS75] A. J. Scott and T. M. F. Smith. Minimax designs for sample surveys. Biometrika, 62(2):353–357, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335372.

Singh:1977:PDI

[SS77] Bahadur Singh and J. Sedransk. The preposterior distribution of the inverse of a sum of two cross-product matrices. Biometrika, 64(3):649–651, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2345353.

Sobel:1971:OAO

[ST71] Milton Sobel and Yung Liang Tong. Optimal allocation of observations for partitioning a set of normal populations in comparison with a control. Biometrika, 58(1):177–181, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334329.

Silvey:1973:GAO

[ST73] S. D. Silvey and D. M. Titterington. A geometric approach to optimal design theory. Biometrika, 60(1):21–32, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334902.

Silvey:1974:LAO

[ST74] S. D. Silvey and D. M. Titterington. A Lagrangian approach to optimal design. Biometrika, 61(2):299–302, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334358.

Stephens:1975:APC

[Ste75a] M. A. Stephens. Asymptotic properties for covariance matrices of order statistics. Biometrika, 62(1):23–28, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334483.

Stephens:1975:NTM

[Ste75b] M. A. Stephens. A new test for the modal vector of the Fisher distribution. Biometrika, 62(1):171–174, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334500.

Stephens:1977:GFE

[Ste77] M. A. Stephens. Goodness of fit for the extreme value distribution. Biometrika, 64(3):583–588, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2345336.

Stephens:1979:TFL

[Ste79a] M. A. Stephens. Tests of fit for the logistic distribution based on the empirical distribution function. Biometrika, 66(3):591–595, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335180.

Stephens:1979:VC

[Ste79b] M. A. Stephens. Vector correlation. Biometrika, 66(1):41–48, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335240.

Stigler:1970:EAG

[Sti70] Stephen M. Stigler. Estimating the age of a Galton–Watson branching process. Biometrika, 57(3):505–512, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334768.

Stigler:1971:EPE

[Sti71] Stephen M. Stigler. The estimation of the probability of extinction and other parameters associated with branching processes. Biometrika, 58(3):499–508, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334385.

Stigler:1973:SHP

[Sti73] Stephen M. Stigler. Studies in the history of probability and statistics. XXXII. Laplace, Fisher, and the discovery of the concept of sufficiency. Biometrika, 60(3):439–445, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334992.

Stigler:1974:SHP

[Sti74] Stephen M. Stigler. Studies in the history of probability and statistics. XXXIII. Cauchy and the witch of Agnesi: an historical note on the Cauchy distribution. Biometrika, 61(2):375–380, August 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334368.

Stigler:1975:SHP

[Sti75] Stephen M. Stigler. Studies in the history of probability and statistics. XXXIV. Napoleonic statistics: The work of Laplace. Biometrika, 62(2):503–517, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335393.

Stone:1972:BRB

[Sto72] M. Stone. Book review: Foundations of Statistical Inference, by V. P. Godambe; D. A. Sprott. Biometrika, 59(1):236–237, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334645.

Stone:1974:CVM

[Sto74] M. Stone. Cross-validation and multinomial prediction. Biometrika, 61(3):509–515, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334733.

Stone:1977:AAC

[Sto77]  M. Stone.  Asymptotics for and against cross-validation.

Biometrika, 64(1):29–35, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335766.

Strauss:1975:MC

[Str75] David J. Strauss. A model for clustering. Biometrika, 62(2):467– 475, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2335389.

Strauss:1977:ROC

[Str77] David Strauss. Runs of occupied cells. Biometrika, 64(1):170–171, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335795.

Styan:1970:NDQ

[Sty70] George P. H. Styan. Notes on the distribution of quadratic forms in singular normal variables. Biometrika, 57(3):567–572, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334774.

Subrahmaniam:1969:OSC

[Sub69] K. Subrahmaniam. Order statistics from a class of non-normal distributions. Biometrika, 56(2):415–428, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334435. See correction [Sub70].

Subrahmaniam:1970:COS

[Sub70] K. Subrahmaniam. Corrections: “Order Statistics from a Class of Non-Normal Distributions”. Biometrika, 57(1):226, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334965. See [Sub69].

Sobel:1970:PWS

[SW70] Milton Sobel and George H. Weiss. Play-the-winner sampling for selecting the better of two binomial populations. Biometrika, 57 (2):357–365, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334844.

Srinivasan:1973:LDR

[SW73] R. Srinivasan and R. M. Wharton. The limit distribution of a random variable used in the construction of confidence bands. Biometrika, 60(2):431–432, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334562.

Sweeting:1978:ETB

[Swe78] Trevor J. Sweeting. On efficient tests for branching processes. Biometrika, 65(1):123–127, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335286.

Simon:1975:SAB

[SWH75] Richard Simon, George H. Weiss, and David G. Hoel. Sequential analysis of binomial clinical trials. Biometrika, 62(1):195–200, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334503.

Switzer:1976:CPT

[Swi76] Paul Switzer. Confidence procedures for two-sample problems. Biometrika, 63(1):13–25, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335079.

Shuster:1974:OAO

[SY74] Jonathan J. Shuster and Mark C. K. Yang. Optimal allocation of observations for partitioning a set of new items into two control groups. Biometrika, 61(1):191–193, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334304.

Tiao:1971:ACR

[TA71] G. C. Tiao and M. M. Ali. Analysis of correlated random effects: Linear model with two random components. Biometrika, 58(1):37– 51, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334315.

Tamhane:1978:IBR

[Tam78] Ajit C. Tamhane. Inference based on regression estimator in double sampling. Biometrika, 65(2):419–427, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335225.

Tarone:1975:TTL

[Tar75] Robert E. Tarone. Tests for trend in life table analysis. Biometrika, 62(3):679–682, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335528.

Targhetta:1979:CIO

[Tar79a] M. Luisa Targhetta. Confidence intervals for a one-parameter family of multivariate distributions. Biometrika, 66(3):687–688, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335197.

Tarone:1979:TGF

[Tar79b] R. E. Tarone. Testing the goodness of fit of the binomial distribution. Biometrika, 66(3):585–590, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335179. See correction [Tar90].

Tarone:1990:ACT

[Tar90] R. E. Tarone. Amendments and corrections: “Testing the Goodness of fit of the Binomial Distribution”. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2337011.

See [Wis68, Tar79b].

Tiao:1978:SCD

[TH78] G. C. Tiao and S. C. Hillmer. Some consideration of decomposition of a time series. Biometrika, 65(3):497–502, December 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335900.

Theobald:1975:IAM

[The75] C. M. Theobald. An inequality with application to multivariate analysis. Biometrika, 62(2):461–466, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335388. See correction [The76].

Theobald:1976:CAI

[The76] C. M. Theobald. Corrections and amendments: “An inequality with application to multivariate analysis” (Biometrika 62 (1975), no. 2, 461–466). Biometrika, 63(3):685, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335757. See [The75].

Thiebaux:1974:MVE

[Thi74] H. Jean Thiebaux. Minimum variances estimation of coefficient matrices in a dependent system. Biometrika, 61(1):87–90, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334288.

Thorburn:1976:SAP

[Tho76] Daniel Thorburn. Some asymptotic properties of jackknife statistics. Biometrika, 63(2):305–313, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335624.

Tiao:1972:ABT

[Tia72] G. C. Tiao. Asymptotic behaviour of temporal aggregates of time series. Biometrika, 59(3):525–531, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334803.

Tiku:1970:MCS

[Tik70] M. L. Tiku. Monte Carlo study of some simple estimators in censored normal samples. Biometrika, 57(1):207–211, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334953.

Tiku:1971:EMS

[Tik71a] M. L. Tiku. Estimating the means and standard deviation from two censored normal samples. Biometrika, 58(1):241–243, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334341.

Tiku:1971:NDH

[Tik71b] M. L. Tiku. A note on the distribution of Hotelling’s generalized T_{0}^{2}. Biometrika, 58(1):237–241, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334340.

Titterington:1975:ODS

[Tit75] D. M. Titterington. Optimal design: Some geometrical aspects of D-optimality. Biometrika, 62(2):313–320, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335366.

Titterington:1977:AIM

[Tit77] D. M. Titterington. Analysis of incomplete multivariate binary data by the kernel method. Biometrika, 64(3):455–460, December 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345321.

Tjostheim:1978:MAS

[Tjø78] Dag Tjøstheim. A measure of association for spatial variables. Biometrika, 65(1):109–114, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335284.

Tolley:1979:TTE

[TN79] H. Dennis Tolley and James E. Norman. Time on trial estimates with bivariate risks. Biometrika, 66(2):285–291, August 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335660.

Tong:1971:IEO

[Ton71] Yung Liang Tong. Interval estimation of the ordered means of two normal populations based on hybrid estimators. Biometrika, 58(3):605–613, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334395.

Thigpen:1974:MRT

[TP74] Charles C. Thigpen and A. S. Paulson. A multiple range test for analysis of covariance. Biometrika, 61(3):479–484, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334729.

Troutman:1979:SRP

[Tro79] Brent M. Troutman. Some results in periodic autoregression. Biometrika, 66(2):219–228, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335652.

Tuan:1979:EPA

[Tua79] Pham Dinh Tuan. The estimation of parameters for autoregressivemoving average models from sample autocovariances. Biometrika, 66(3):555–560, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335175.

Tukey:1958:BCQ

[Tuk58] John W. Tukey. Bias and confidence in not-quite large samples (abstract). Annals of Mathematical Statistics, 29(2):614, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/

1177706647. This is the other half of the Quenouille–Tukey jackknife technique; see [Que49, Que56, Mil74, Efr79].

Tiao:1976:ETA

[TW76] G. C. Tiao and W. S. Wei. Effect of temporal aggregation on the dynamic relationship of two time series variables. Biometrika, 63(3):513–523, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335729.

Tarone:1977:DFT

[TW77] Robert E. Tarone and James Ware. On distribution-free tests for equality of survival distributions. Biometrika, 64(1):156–160, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335790.

Tong:1979:BPF

[TW79] Y. L. Tong and David E. Wetzell. On the behaviour of the probability function for selecting the best normal population. Biometrika, 66(1):174–176, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335261.

Taheri:1974:CSS

[TY74] H. Taheri and D. H. Young. A comparison of sequential sampling procedures for selecting the better of two binomial populations. Biometrika, 61(3):585–592, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334743.

Upton:1976:IPH

[UL76] G. J. G. Upton and R. D. Lee. The importance of the patient horizon in the sequential analysis of binomial clinical trials. Biometrika, 63(2):335–342, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335628.

Upton:1973:SST

[Upt73] G. J. G. Upton. Single-sample tests for the von Mises distribution. Biometrika, 60(1):87–99, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334909.

Upton:1974:NAD

[Upt74] G. J. G. Upton. New approximations to the distribution of certain angular statistics. Biometrika, 61(2):369–373, August 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334367.

VanderParren:1970:TDF

[Van70] J. L. Van der Parren. Tables for distribution-free confidence limits for the median. Biometrika, 57(3):613–617, December 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334779.

VanderParren:1973:TDF

[Van73a] J. L. Van der Parren. Tables of distribution-free one-sided and twosided confidence limits for quartiles. Biometrika, 60(2):433–434, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334564.

VanNess:1973:ACP

[Van73b] John W. Van Ness. Admissible clustering procedures. Biometrika, 60(2):422–424, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334558.

Venable:1978:GTI

[VB78] T. C. Venable and V. P. Bhapkar. Gart’s test of interaction in a 2 × 2 × 2 contingency table for small samples. Biometrika, 65(3):669–672, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335930.

Vit:1973:IPP

[Vit73] Pedro Vit. Interval prediction for a Poisson process. Biometrika, 60(3):667–668, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335020.

Vit:1974:THG

[Vit74] Pedro Vit. Testing for homogeneity: The geometric distribution. Biometrika, 61(3):565–568, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334740.

vanKampen:1973:BDP

[vK73] N. G. van Kampen. Birth and death processes in large populations. Biometrika, 60(2):419–420, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334556.

Vuchkov:1977:RTP

[Vuc77] I. N. Vuchkov. A ridge-type procedure for design of experiments. Biometrika, 64(1):147–150, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335787.

Wood:1978:LSR

[WA78] Constance L. Wood and Michele M. Altavela. Large-sample results for Kolmogorov-Smirnov statistics for discrete distributions. Biometrika, 65(1):235–239, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335304.

Wahrenberger:1977:BRT

[WAK77] David L. Wahrenberger, Charles E. Antle, and Lawrence A. Klimko. Bayesian rules for the two-armed bandit problem. Biometrika, 64(1):172–174, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335796.

Walker:1967:STS

[Wal67] A. M. Walker. Some tests of separate families of hypotheses in time series analysis. Biometrika, 54(1/2):39–68, June 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333851. See correction

[Wal70a].

Walker:1970:CST

[Wal70a] A. M. Walker. Corrections: “Some Tests of Separate Families of Hypotheses in Time Series Analysis”. Biometrika, 57(1):226, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334966. See

[Wal67].

Walton:1970:NNN

[Wal70b] Gerald S. Walton. A note on nonrandomized Neyman-shortest unbiased confidence intervals for the binomial and Poisson parameters. Biometrika, 57(1):223–224, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334960.

Walker:1971:EHC

[Wal71] A. M. Walker. On the estimation of a harmonic component in a time series with stationary independent residuals. Biometrika, 58(1):21–36, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334314.

Walter:1975:DLM

[Wal75] S. D. Walter. The distribution of Levin’s measure of attributable risk. Biometrika, 62(2):371–374, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335374.

Watson:1971:EFP

[Wat71] G. S. Watson. Estimating functionals of particle size distributions. Biometrika, 58(3):483–490, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334383.

Watson:1972:PEL

[Wat72] G. S. Watson. Prediction and the efficiency of least squares. Biometrika, 59(1):91–98, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334619.

Watson:1974:BII

[Wat74] S. R. Watson. On Bayesian inference with incompletely specified prior distributions. Biometrika, 61(1):193–196, April 1974. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334305.

Waternaux:1976:ADS

[Wat76] Christine M. Waternaux. Asymptotic distribution of the sample roots for a nonnormal population. Biometrika, 63(3):639–645, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335746.

Webster:1970:AME

[Web70] J. T. Webster. On the application of the method of Das in evaluating a multivariate normal integral. Biometrika, 57(3):657–660, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334784.

Wedderburn:1974:QLF

[Wed74] R. W. M. Wedderburn. Quasi-likelihood functions, generalized linear models, and the Gauss-Newton method. Biometrika, 61 (3):439–447, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334725.

Wedderburn:1976:EUM

[Wed76] R. W. M. Wedderburn. On the existence and uniqueness of the maximum likelihood estimates for certain generalized linear models. Biometrika, 63(1):27–32, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335080.

Weisberg:1974:ECP

[Wei74] Sanford Weisberg. An empirical comparison of the percentage points of W and W′. Biometrika, 61(3):644–646, December 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334756.

Wei:1978:RAD

[Wei78] L. J. Wei. On the random allocation design for the control of selection bias in sequential experiments. Biometrika, 65(1):79–84, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335280.

White:1973:EGE

[Whi73a] Lynda V. White. An extension of the general equivalence theorem to nonlinear models. Biometrika, 60(2):345–348, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334546.

Whittaker:1973:BMM

[Whi73b] J. Whittaker. The Bhattacharyya matrix for the mixture of two distributions. Biometrika, 60(1):201–202, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334923.

Whitehead:1978:LSS

[Whi78] John Whitehead. Large sample sequential methods with application to the analysis of 2 × 2 contingency tables. Biometrika, 65(2):351–356, August 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335214.

Wichern:1973:BSA

[Wic73] Dean W. Wichern. The behaviour of the sample autocorrelation function for an integrated moving average process. Biometrika, 60(2):235–239, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334535.

Wilkinson:1970:GRP

[Wil70a] G. N. Wilkinson. A general recursive procedure for analysis of variance. Biometrika, 57(1):19–46, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334933.

Williams:1970:CMC

[Wil70b] E. J. Williams. Comparing means of correlated variates. Biometrika, 57(2):459–461, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2334857.

Williams:1975:MSW

[Wil75a] C. B. Williams. Mendenhall’s studies of word-length distribution in the works of Shakespeare and Bacon. Biometrika, 62(1):207–212, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334505.

Williams:1975:EBD

[Wil75b] E. R. Williams. Efficiency-balanced designs. Biometrika, 62(3): 686–689, December 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335531.

Williams:1976:ILR

[Wil76] D. A. Williams. Improved likelihood ratio tests for complete contingency tables. Biometrika, 63(1):33–37, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335081. See correction [Wil78].

Williams:1977:SIP

[Wil77a] David A. Williams. Some inference procedures for monotonically ordered normal means. Biometrika, 64(1):9–14, April 1977. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335763.

Williams:1977:SRS

[Wil77b] E. J. Williams. Some representations of stable random variables as products. Biometrika, 64(1):167–169, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335794.

Williams:1978:CIL

[Wil78] D. A. Williams. Correction to: “Improved likelihood ratio tests for complete contingency tables” (Biometrika 63 (1976), no. 1, 33– 37). Biometrika, 65(3):677, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335935. See [Wil76].

Wisniewski:1968:THB

[Wis68] T. K. M. Wisniewski. Testing for homogeneity of a binomial series. Biometrika, 55(2):426–428, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2334888. See corrections [Tar90, Wis90].

Wisniewski:1990:ACT

[Wis90] T. K. M. Wisniewski. Amendments and corrections: Testing for Homogeneity of a Binomial Series. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337012. See

[Wis68].

Whitehead:1979:ASC

[WJ79] John Whitehead and David Jones. The analysis of sequential clinical trials. Biometrika, 66(3):443–452, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335163.

Williams:1977:MCS

[WL77] J. S. Williams and S. W. Lagakos. Models for censored survival analysis: Constant-sum and variable-sum models. Biometrika, 64 (2):215–224, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335687.

Wolfe:1976:TER

[Wol76] Douglas A. Wolfe. On testing equality of related correlation coefficients. Biometrika, 63(1):214–215, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335110.

Wood:1970:VST

[Woo70] J. T. Wood. A variance stabilizing transformation for coefficients of concordance and for Spearman’s rho and Kendall’s Tau. Biometrika, 57(3):619–627, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334780.

Woodroofe:1976:FPB

[Woo76] Michael Woodroofe. Frequentist properties of Bayesian sequential tests. Biometrika, 63(1):101–110, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335089.

Wood:1978:LSK

[Woo78] Constance L. Wood. A large-sample Kolmogorov-Smirnov test for normality of experimental error in a randomized block design. Biometrika, 65(3):673–676, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335931.

Woodroofe:1979:RLR

[Woo79] Michael Woodroofe. Repeated likelihood ratio tests. Biometrika, 66(3):453–463, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335164.

Worthington:1975:GIM

[Wor75] B. A. Worthington. General iterative method for analysis of variance when block structure is orthogonal. Biometrika, 62(1):113– 120, April 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334492.

Wynn:1975:IOS

[Wyn75] H. P. Wynn. Integrals for one-sided confidence bounds: A general result. Biometrika, 62(2):393–396, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335378.

Wynn:1976:CFP

[Wyn76] H. P. Wynn. Connected finite population sampling plans. Biometrika, 63(1):208–210, April 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335108.

Wynn:1977:ICB

[Wyn77] H. P. Wynn. An inequality for certain bivariate probability integrals. Biometrika, 64(2):411–414, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335709.

Yang:1976:DPD

[Yan76] Mark C. K. Yang. A design problem for determining the population direction of movement. Biometrika, 63(1):77–82, April 1976. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335086.

Yang:1978:DSS

[Yan78] Mark C. K. Yang. Determination of sample size for testing the relation between an incident and a set of random variables in a sample survey. Biometrika, 65(3):635–640, December 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335916. See correction [Yan80].

Yang:1980:ACD

[Yan80] Mark C. K. Yang. Amendments and corrections: “Determination of Sample Size for Testing the Relation Between an Incident and a Set of Random Variables in a Sample Survey”. Biometrika, 67(3):728, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335154. See [Yan78].

Yates:1972:MCT

[Yat72] F. Yates. A Monte-Carlo trial on the behaviour of the nonadditivity test with nonnormal data. Biometrika, 59(2):253–261, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334570.

Yuen:1973:ABP

[YD73] Karen K. Yuen and W. J. Dixon. The approximate behaviour and performance of the two-sample trimmed t. Biometrika, 60 (2):369–374, August 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334550.

Young:1970:CPS

[You70a] D. H. Young. Consideration of power for some two sample tests with censoring based on a given order statistic. Biometrika, 57 (3):595–604, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334777. See correction [You72a].

Young:1970:OSN

[You70b] D. H. Young. The order statistics of the negative binomial distribution. Biometrika, 57(1):181–186, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334947.

Young:1971:MRO

[You71] D. H. Young. Moment relations for order statistics of the standardized gamma distribution and the inverse multinomial distribution. Biometrika, 58(3):637–640, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334399.

Young:1972:CAC

[You72a] D. H. Young. Corrections and amendments: “Consideration of Power for Some Two-Sample with Censoring Based on a Given Order Statistic”. Biometrika, 59(1):238, April 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334647. See [You70a].

Young:1972:IPM

[You72b] J. C. Young. An investigation of procedures for multiple-stage selection for a variate subject to errors of measurement. Biometrika, 59(2):323–334, August 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334577.

Young:1973:DSC

[You73a] D. H. Young. Distributions of some censored rank statistics under Lehmann alternatives for the two-sample case. Biometrika, 60(3):543–549, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335003.

Young:1973:SRO

[You73b] D. H. Young. Some results for the order statistics of the negative binomial distribution under slippage configuration. Biometrika, 60(1):206–209, April 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334925.

Young:1974:MSS

[You74] J. C. Young. Multiple-stage screening from nonnormal populations. Biometrika, 61(1):155–164, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334298.

Young:1977:BAP

[You77] A. S. Young. A Bayesian approach to prediction using polynomials. Biometrika, 64(2):309–317, August 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335698.

Yuen:1974:TST

[Yue74] Karen K. Yuen. The two-sample trimmed t for unequal population variances. Biometrika, 61(1):165–170, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334299.

Zelen:1971:ASC

[Zel71] M. Zelen. The analysis of several 2 × 2 contingency tables. Biometrika, 58(1):129–137, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334323.

Zentgraf:1975:NLD

[Zen75] R. Zentgraf. A note on Lancaster’s definition of higher-order interactions. Biometrika, 62(2):375–378, August 1975. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335375.

Zajta:1977:TSP

[ZP77] A. J. Zajta and W. Pandikow. A table of selected percentiles for the Cramér-von Mises-Lehmann test: equal sample sizes. Biometrika, 64(1):165–167, April 1977. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335793.